A Complete Bibliography of Publications in Technometrics for the decade 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

03 June 2024 Version 1.02

Title word cross-reference

2 [DT68]. 2^{17−9} [Add65, Pre66]. 2^{k−p} [BH61a, BH61b]. 2^{n} [Add61, Coo68, Mar69b]. 2^{n−1} [Joh66a]. 2^{n}3^{m} [Mar69a, Mar69b]. 2^{p−q} [DW66a]. 2Î [Ku63b]. 2 × 2 [Owe65c, Owe66a]. 2 × 3 [BN63, BN64]. 3(2^{n−k}) [Joh64a]. 32 [Add65]. 3^{4n+1} [Rag65]. 50 [Har66]. ^{3} [CHDF67, Kao69]. ^{m} [Mar67, Mar68b]. ^{n} [Mar67, Mar68b, Rec67]. χ^{2} [JB61]. D_{mn}(n ≤ m ≤ 20) [Zel62a]. E(y) = α + δx + βρ^{x} [SK65]. F [Fox63, Web68c]. k [Bar67b]. λ(r, n) [Wil65]. m [HM65b]. N [Ano65i, WW64]. n = 26 [Wil65]. s [Har66]. s = 1 [Har66]. t [DT68, GB65, Hal63, Kra66b, Owe68]. T^{2} [JB61]. U [Sho67b]. V [Add65, Pre66]. V_{1}^{(s)} [Har66]. W [SW68].

-Distribution [Kra66b, Owe68]. -Statistic [Web68c, Sho67b]. -Tests [JB61]. -Way [Ano65i, WW64].

1 [KLTW65b, McK69d, Sam69, Zys64]. 123rd [Ano64a]. 124th [Ano65a].

125th [Ano66a]. 126th [Ano67a]. 127th [Ano68a]. 128th [Ano69a]. 1620 [Chu66a, Rob64]. 1949 [For66]. 1958 [Dod60, Zel63a]. 1963 [Fis67]. 1966 [Str69].

2 [KLTW65c, Pfa63a].

3 [BCK78, Hay68, Luk64a].

4 [Luk64b].

5 [Ter65].

Abadie [Ran68b]. above [Bar67b]. Abramowitz [Owe65a]. Absolute [Kru66]. Abstracts [Ano64a, Ano65a, Ano66a, Ano67a, Ano68a, Ano69a]. Abuse [Box66c]. Accelerated [BCM62, Che62]. Acceptable [FPS65]. Acceptance [Cox60, Cra68, Eps60d, HB60, JB61, Owe69, Pat63c, PZB62, SD69, ZT68]. Acceptance-Rejection [HB60]. Accuracy [Joh62b]. Acknowledgement [Pat63c]. Ackoff [Thr64, Olm62]. Acton [Dyk67a]. Adaptive [Pel69]. Addelman [Ano62f, Pre66]. Addendum [Ano62a]. Additional [JA67]. Additive [Pot62b, Pot62a]. Adequacy [PL65]. Adjust [Hil60]. Adjusted [TG67]. Advanced [Hay68, Pfa63a, Str69, Str69]. Advances [Hes68, Ter65]. after [Hin67]. against [DW66a, JS68b, Kus69, Man62]. Age [Gla67]. Ahmed [Zel63b]. Aid [CJ64]. Aided [KS69]. Aids [GD60]. Albert [Mee69]. Alder [Cra65, Dra63a]. Algebra [Pet66]. Algebraic [Koo64]. Algorithm [Coo68, Hun66]. Aliasing [Mar68b]. Alison [For66, Zel63a]. Allocation [LJ69b, Mor61, Wei61a]. Alt [Ter65]. Alternatives [MH68, Sho67b]. American [Ano64a, Ano65a, Ano68a, Ano69a]. among [BH67, Gov63]. Analogue [Bar62a]. Analyser [PG65]. Analyses [Wei61a]. Analysing [Kru63]. Analysis [And60, AT63, Aro68, Boc63, Bra66, Bra68, Chu66b, Coc68a, Cot69, CL67, CHDF67, Cro68, De 64, Dem60, Dra62, Dra63b, Enr60, Goo61, HL67, Jac66, JL67, Jen61, JA67, Jon65, Kap66, Kra66a, Kra67, Kra63b, Lac66, LM68, LE64, Man64a, Man64b, McK68b, McL69, Nob63, NH69, Olm63, Pas62, Pot67, PO61, Pra67, Rei62, See68, Str69, TG67, Tuk61, Wat69a, Wes62, Won61, dB64, Bai67, Bla66, Hea61, Mai66, Mee66, Roo63, Sha66, Dyk67a, Low65]. Analyzing [Mar67]. Anderson [Gor66]. Animal [Rob63]. Annual [Ano64a, Ano65a, Ano66a, Ano67a, Ano68a, Ano69a]. ANOVA [Cox92, Fow69, HB69, JL69, LJ69a]. Anscombe [De 64, Olm63, KLTW65a]. Anthanasios [Mil66]. Anthony [Mai66]. Any [Coo68]. Appearance [TS66]. Application [Add64a, Bec67, Bur67, DeC65, Dun60, Eva67, GM66, JB61, JS68a, Joh61, LE64, Mal69, Pos65, PJ67, Sch69, SHH62, Sum65, VT68, Neu68]. Applications [Goo69, Gru67, HB61, HH66a, LJ69a, Mar67, Sam69, SG61, Hal63, Owe68, Cow67, Eri66, Joh66b, Tra65, Bak64, Ros65a]. Applied [All68, Eva63, LJ69b, Wat69b, Wei62a, Bai67, Gil65, Ran69, Sto61]. Appraisal [Bro61]. Approach [Dun65, Kao69, Koc67a, Koc68, Wol68, Gra69, Rog68a]. Approximate [DT68, Mad65, MS68, Rob65, Tal67, Web68c]. Approximating [Bar66]. Approximation [Bur67, GH66, Kra66b, LSL68, Mee69, ML60, GB65]. Approximations [Bin68, Hal63, KLTW65c, SW68, Tip63]. Approximator [Aig68]. Arbitrary [Bra68]. Area [Eck69]. Arguments [Goo62]. Arising [Bar63b, Wei61a]. Arithmetic [Ham65b, Ham65a, JS68b]. Army [Str69]. Arrangements [Bra68]. Arrow [Wei62a]. Arthur [All62, Owe63, Owe65b]. ASA [Ano66a, Ano67a]. Assay [PL65]. Assessment [Dyk66a, WS68]. Assignments [You64]. Associated [Eck69, Won61]. Association [Ano69a, Ano64a, Ano65a, Ano68a]. Assuming [Kem67, Kus69]. Assumption [Eps60e, Eps60f]. Assurance [HB60, Wal69]. Astronomy [Bar63c]. Asymmetric [MD68]. Asymmetrical [Add62a, Add62b]. Asymptotic [Dub65, HB61]. Asymptotically [JL66]. Atkinson [Fer66]. Attack [Dun62]. Attribute [Hal67, Hal68]. Attributes [Cra68, Dod60, Ell63, PB67, Pha68a]. Augmented [Box66a]. Augmenting [GM68, Joh66a]. Author [KLTW65a]. Autocorrelated [Ler65]. Automatic [Lie65, MBM64, Moh69, Wil66]. Automobiles [BM60]. Availabilities [Nel68]. Availability [GL67]. Average [Ano61e, Cra68, GW61, MT65]. Averages [Lin64]. Avila [Neu68]. Avila-Beloso [Neu68].

B [Cad69, Coo66, Cra65, Dra63a, Fle63, Gei62, Gei63, Ham65b, Ham65a, Har62, Joh68, Lau68, Owe65c, Owe66a, Rei63b, Sys63]. Back [Ano60a, Ano60b, Ano60c, Ano61a, Ano61b, Ano61c, Ano62b, Ano62c, Ano62d, Ano63a, Ano63b, Ano63c, Ano63d, Ano64b, Ano64c, Ano64d, Ano65b, Ano65c, Ano65d, Ano66b, Ano66c, Ano66d, Ano66e, Ano67b, Ano67c, Ano67d, Ano67e, Ano68b, Ano68c, Ano68d, Ano68e, Ano69b, Ano69c, Ano69d, Ano69e]. Balanced [Cla67, Doe65, Joh61, KLTW65b, KLTW65c, Pre69, Rob61, LN66]. Banbury [Cha62b]. Bancroft [McK69d]. Barlow [Ams65a]. Bartko [Ano67g]. Bartlett [Jen62, Pfa64, Rou68, Won61]. Barton [Gil63]. Based [Coh65, Dod62, Dun60, Eps64a, Hal60, Har64, HS66, Hew68, Jae65, Owe67, Pfa63b, PZB62, HM65b]. Bases [BF65, Wol68]. Basic [Pri62, Cow67, Dav64, Gaf65, McK69a, Put65]. Bass [Tho67]. Batch [DGR61]. Batching [HB60]. Bauer [Dan61]. Bayes [Cox60, SZ68]. Bayesian [Gei66, Bra66, Dun65, Hal68, Pos63, ST68, Wet60, Ran68a]. Be [HO63, VH69]. Beale [GM65]. Beckenbach [Gil65]. behavior [DT68]. Behavioral [Van64]. Beloso [Neu68]. below [Bar67b]. Bendat [Cro68]. Bennett [Owe65c, Owe66a]. Berger [Rei63a]. Bergman [Nob63]. Bergström [Tuc65]. Berkeley [All63, Bar63c]. Berman [Cor69]. Bernard [Rei63a, Str69, Zel63b]. Bernoulli [Bra66]. Bert [Van64]. Best [Dol60, GQS67, HL67, Man67, McC65, SD68]. Beta [Bra66, Fox63, GPH67, McK68a, Wei65]. Betteley [Chu68a]. Better [DHT69]. between [Ano64g, EP63b, HR65, PW65, Qur64, QNA65, SZ68, Tuk61]. Bevan [Tuc68]. Bias [CW69, KMH69]. Bibliography [Fis67, For66, Zel63a]. Billingsley [Pyk63]. Bimodal [Eis64]. Binary [Wou66]. Binomial [Ano66g, Ano67f, Bar62c, Bar66, Bar67a, Coh60, Gue69, Mac69, ML60, Pat63c, SG66, Ste65, Was64, Wei65, Pat60]. Biological [Pet66, And66]. Biologists [Kra66a]. Biology [All63, Lea67]. Biomedical [Gaf65]. Biostatistics [McK69b]. Biplot [BG78]. Bivariate [CD67, CB62, LE64, Nad67, Smi68b]. Black [Fra68]. Blackwell [McK69a]. Bliss [Lea67]. Block [AB64, Doe65, Joh61, Kap66, Pre69, Rob61]. Blocking [Joh64a]. Blocks [Add65, Pre69]. Bloemena [Har67a]. Book [All62, All68, All63, Ams65b, Ams65a, And66, Bai67, Bak64, Ban69, Bar63a, Bar63c, Bar68, Bas68, Ber61, Ber66, Ber67, Bla66, Blo66, Blu66, BP68, Boo62, Bor66, Bro68, Buc67, Bur68, Bur69, Cad69, Cap66, Cha62a, Cha62b, Chu68a, Chu66a, Coh66d, Coh66c, Con63, Coo66, Cor69, Cot69, Cow67, Cra65, Cro68, Cry67, Dan61, Dan63, Dav64, Dav66, Dem60, Dod60, Dod62, Doo69, Dra63a, Dre63, Dun60, Dun61, Dyk65a, Dyk65b, Dyk66a, Dyk67a, Dyk67b, Eps64a, Eri66, Fei66, Fer66, Fis67, Fle63, For66, Gaf65, Gan69, Gei62, Gei63, Gei64, Gei66, Gho67, Gil63, Gil65, Gil66, Gol63, Gov66, Gra69, Gre68, Gru67, Hag66, Ham62, Ham65b, Ham65a, Har67a, Har62, Har66, Hay68, Hea61, Hes68, Hic68, Hog66, Hol68, Hon66, Hus69, Ire65, Jac66]. Book [Jef63, Jen62, Joh66b, Joh68, Kan61, KR68, Kaw65, Kei63, Kip65, Koe65, Koo64, Koo68, Kor68, Kra66a, Kru63, Ku63a, Lau68, Lea67, Lei68, Lev66, Li68, Lig69, Llo61, Low65, Luk63, Luk64a, Luk64b, Lut68, Mad68, Mai66, Man61, Man64c, McK68b, McK68a, McK69d, McK69a, McK69b, McK69c, McL69, Mee66, Mee69, Mil63, Mil66, Min65b, Min65a, Mul65, Nad69b, Nau68, Neu68, Ney65, Nob63, Noe63, Ode66a, Ode66b, Ode68, Olm62, Ost64, Owe63, Owe65a, Owe65b, Owe65c, Owe66a, Par64, Pet66, Pet64, Pfa63a, Pfa64, Pos66, Pot67, Put65, Pyk63, Ran68b, Ran68a, Ran69, Rei62, Rei63b, Rei63a, Rob64, Rob63, Rog68a, Rog68b, Rog69b, Rog69a, Roo63, Ros64, Ros65b, Ros65a, Rou68, Roy67, S.65a, S.65b, Sam69, Sat65, Sch66, Sew66, Sha66, Ste65, Sto61, Str69]. Book [Sys63, Tan68, Ter65, Tho63, Tho67, Thr64, TC68, Tra65, Tuc65, Tuc68, Utt66, Van64, Ver66, Wal68, Wat69a, Wat69b, Wei62a, Wol68, Wou63, Wou64, Wou65, Zac68, Zel62b, Zel62a, Zel63b, Zel63a, Zim68, Zys64]. Books [Ano63e, Ano63f, Ano63g, Ano64e, Ano64f, Ano65e, Ano65f, Ano65g, Ano65h]. Borel [Hag66]. Bose [Ano78]. Both [CB67, Mal69, Owe64]. Bound [Gut62, JL66]. Bounded [Eps61]. Bounds [Har64]. Bower [Fer66]. Box [Fra68, Ano61f, Ano61g, Ano63i]. Brabb [Hus69]. Bracken [Owe65b]. Bradu [Ano78]. Branching [Ney65]. Bretherton [Ste65]. Brooks [Chu68a]. Brown [Cow67, Fle63, Roo63, Lut68]. Brownlee [Dyk65a, Llo61]. Brunk [Coh66c, Ku63a]. Bruyn [Bur69]. Bryant [Sha66]. Buckland [Dyk66a]. Building [BH62, CJ64]. Bulk [Bec67, Dun62]. Bulmer [Roy67]. Bunt [Gho67]. Burford [Cry67]. Burn [BMPS68, Law66]. Burn-In [BMPS68, Law66]. Bursk [Cow67]. Business [Cow67, Gre68, McL69].

C [All68, Ams65a, And66, Blu66, Bur69, Cha62a, Chu68a, Cow67, Dyk67b, Fer66, GM66, Ham65b, Ham65a, Joh66b, Kip65, Lea67, Low65, McK68b, Mee66, Neu68, Owe63, Owe66a, Sha66, Tra65, Wat69a, Wat69b]. Cahen [Bak64]. Calculating [Wel62]. Calculation [Bur65, Cho62, STF68]. Calculations [Fow69]. Calibration [Ber69, Kru67, Kru69, Wil69]. Cani [Cow67]. Capsules [Rob69]. Carlo [Mul65, BMW65, NG68]. Cartesian [Kan61]. Case [Aig68, GKKK61, GM62, HH69, JL66, Ler65]. Cases [Gue69]. Cell [CG62]. Censored [Ano68f, Ano73, Coh61b, Coh63, Coh65, Eps60c, GK69, Hal66, HM65a, Har67b, HM68, Pha68a, Sal67]. Censoring [Boh68]. Center [Str69]. Centerless [Moh69]. Centerless-Grinding [Moh69]. central [Hal63]. Certain [Cho62, Dub67a, Hew68, Wei62b, Wou66, Lev66]. Chain [HB69, Har62]. Chain-Pooling [HB69]. Chains [BPM63, Joh66d, KKK62, Ran68a]. Chakravarti [All68]. Chance [Gil63]. Changes [DS68a]. Changing [AR66]. Characteristic [Ros65a, Dyk66a]. Characteristics [MT65]. Charles [Ire65]. Chart [Ano61e, GW61, JS68b, Rob66]. Charts [Ewa63, Joh63, Joh66c, Pag61, Tay68]. Chemical [DGR61, Hsi66, Man69a, Wei61a]. Chemists [Chu68a, Dan61]. Chi [Che67, Nel66]. Chi-Square [Che67]. Chiang [McK69b]. Chin [McK69b]. Choice [Cla65]. Choosing [Dol63]. Chover [Doo69]. Clark [Sch66]. Class [Aig68, Cra62]. Classical [Kru67, Kru69, Nau68]. Classification [Ano65i, Bra68, Goo62, Pel69, Que64, WW64]. Classifications [KLTW65b, KLTW65c]. Classified [Bec68a, Bec68b, Sau68]. Clelland [Cow67]. Clipping [Hin67]. Clustering [Nau66]. Coalescence [Mar68a]. Codes [Man61]. Coefficient [Sum65, Wei66]. Coefficients [Ler65, Wol77, Wil65]. Cohen [Rei63a, Ano66j]. Coherent [EP63a]. Collaborative [Bur68]. collected [Cox92]. Combination [Gun65]. Combinations [GH69]. Combinatorial [Gil65, Gan69, Gil63, Man64c, Zac68]. Combining [Ben64, Kap66, Nel66, Won61]. Comer [Smi69b]. Comment [Eps64b, Smi68a]. Commentary [Coc68a]. Comments [Ans65, Goo61, GM66, Jac65, JP61, KLTW65a]. Communication [GM66, Pos65, Rob63, TC68]. Communications [DW66b]. Comparative [Ril62]. Comparing [Kao64, NG68, Rob60, Ste60]. Comparison [Ano67f, Bar67a, BA63, Cox65, JL69, MS68, Nau66, Rob66]. Comparisons [AC67, Bec68a, Dun65, Dun64, KLTW65b, KLTW65c, Ren69, RS65, Rob61, She65a, TSRR67]. Compartmental [BC69]. Compleat [Gil66]. Complete [Ano68f, Ano73, Coh65, Coo68, HM65a, Har67b, Wil65]. Completed [Bar67b]. Complex [And60, Kra63a]. Component [Ano65j, BES61, Dia67, EP63b, Mye64, Sch64]. Components [BA63, EP63a, JL69, Koc67a, Koc68, KA67, Lar66, LNJE68, Mor61, TM63, LN66, Wei61b]. Composite [KA67, PO61]. Composition [Mye64]. Compound [Hal60].

Comprehensive [Fra68]. Computation [Coo66]. Computer [BCK60, BCK61, Cot69, KS69, MH69, Ode68, S.65b, Wol68, BCK78, Fis67, Hon66, Kor68]. Computers [Ano61d, Jos64, Leo61a, Leo61b, Leo61c, Leo61d, Leo62a, Leo62b, Leo62c, Rob64, Ter65, Van64]. Computing [Ham65b, Ham65a]. Comrie [Owe63]. Concepts [Ire65, Put65]. concerning [Koc68]. Conclusions [Tuk60]. Conditional [HM68]. Conducted [Str69]. Conference [Cha62b]. Confidence [Goo65, GL67, Hal66, Har64, HW66, JL66, Kra63a, Kru66, Mad65, Mos65, MS68, Nád69a, QH64, Ros66, San68, ST68, SZ68]. Confounded [CHDF67, JA67]. Confounding [Add64b, You61]. Connectedness [Ano65i, WW64]. Connection [Tuk61]. Connor [Dan63]. Consideration [Ber69]. Considerations [Jen61, Par61, Pri62]. Constants [You62]. Constrained [Atk69, Cho62, Wes62]. Constructing [AB64, BCK60, BCK61, BCK78]. Construction [Add65, McC65]. Consultant [Cam69]. Consulting [Dan69]. Consumer [Sin66, Hal67]. Contagious [Nau68]. Contingency [BN63, BN64, CL67, Hea69, Ku63b, KKK62, Owe65c, Owe66a, BK68]. Contingent [Dur66]. Continued [LMW64]. Continuous [Dod60, Hal68, HO63, Hil64, Mil60]. Contributions [All63, Bar63c, Zel63b]. Control [Ano61e, Bec68a, Dod60, Dod62, Dun60, Eps64b, GW61, Hon66, Hsi66, Joh63, Joh66c, Lie65, Owe64, RS65, Rob66, Rob69, Rob63, Rob61, SG61, Tay67, Tay68, TS64, Wig67, Wil66, Zim68, Kip65]. Controlling [Moh69, Pag63, Que64]. Convergence [Tor63]. convex [Man69b, Blo66]. Convolution [Ano65j, Sch64]. Convolutions [Tuc65]. Corporation [Utt66]. Correct [Kru66]. Corrected [Wel62]. Correcting [Man61]. Correction [Ano61d, Con62, De 64, Olm63]. Correlated [Ben64, Coc65, Llo63]. Correlation [Han69, Wol77]. Corrigenda [Ano80]. Cost [Ril62, Wig67]. Costs [Hal60, Pfa63b]. Course [Mai66, Rog69a]. Coverage [Eck69]. Cox [Gov66, Bar63a, BLH^{+}60, Man62]. Cramér [Ber67]. Criteria [Hil64, Tin63]. Criterion [Eps64a, HHW68, ZT68]. Critical [Bin68, LSL68]. Cross [Ano65i, WW64]. Crothers [Fer66]. Cu [Ewa63]. Cu-Sum [Ewa63]. Cumulants [Fin63, FC60]. Cumulative [Ano61e, GW61, HO63, Joh63, Joh66c, Kra66b, Man64b, Pag61, Pag62, Tay68, Wei61b, Bur69]. Curtailed [PB67, Pha68a, Pha68b]. Curve [Chu66b, SK65]. Curves [LB69]. Cusum [Kem67]. Cusums [Pag63]. Cut [KLTW65b, KLTW65c, OF69]. Cybernetics [Rob63].

D [Bar63a, Coh66c, Fle63, Gei66, Gil63, Gil66, Gov66, Ku63a, Owe65c, Rob64, Tra65]. damage [Wei61b]. Daniel [KFT^{+}60]. Data [Ano66g, Chu66b, Coh60, Eps60b, GM68, GT66, Kra63b, Man64b, Man69a, OF69, Pat66, Pha68b, PO61, Que64, Tal67, TC68, Cro68, Dyk67a, Jac66, Kru63]. David [Gil63, McK69a, Sew66, Ban69]. Davidon [VT68]. Dealing [GS69]. Decimal [Con62]. Decision [Gup65, Hol68, Ran68a, Ran69, Rog68a, Sto61, Lut68]. decisions [Tuk60, McL69]. Decomposition [Goo69]. Decreasing [Pro63]. Defective [NWF63, PB67, Que64]. Defectives [SG66]. DeGray [Ano69f]. Degree [Hoe68b, Kus69]. Degrees [Che67, GH69]. Deming [Dyk67a]. Density [PW65]. Department [Str69]. Dependent [Bas68, SD69]. Derivations [KLTW65c]. Derived [Cox60, ZD61]. Descriptive [Ver66]. Design [AH68, Atk69, Box63, Coh66d, Dav63, DeG68a, Gor66, HHW68, Ire65, Joh61, Kao69, KS69, McK68b, MA66, Rob65, She65b, Tay68, KMH69, Gol63]. Designs [Add64a, AB64, AC67, Ano63h, BC62, BH61a, BH61b, Box66a, Che62, Cla67, Dan63, Dia67, Doe65, Dra61, Dra62, DL67, DS68a, DS68b, Dyk60, Fry61, GH62, Hil60, HH69, HR65, Joh64a, Joh66a, Kap66, KA67, LNJE68, Mar68b, Mar69a, Mar69b, MD68, OM68, Pat63b, Pot62b, Pre69, Rec67, Rob61, SHH62, TS64, TM68, Web68a, Web68b, WM61, BB60, LN66, Pot62a]. Destructive [NWF63]. Detecting [Gru69]. Detection [TC68]. Determination [Ano65i, FW68, Hal67, Sal67, Sau68, WW64]. Development [BS66, HS66, Hin68]. Developments [Wil66]. Deviation [Eis68, Gut62, Kra67, Kru66, Pag63, Ros66, Tip63]. Device [Moh69]. Devices [Wil66]. Diagnostic [BG78]. Diamond [Moh67, Hog66]. Diamond-Pin [Moh67]. Dictionary [Rog68b, Ros65b]. Difference [Bin68, Ham62]. Differences [NG68]. Different [BA63, MH68]. Differential [Ham62]. Difficulties [Pat66]. Digital [Coo66, Cot69, Rob65, TC68, Ode68, Van64]. Dilution [Wei61a]. Dimension [Moh69]. Dimensional [BPM63, Dia67, FVM69, BK68]. Direct [Aro68, AR69]. Discrete [Nau68, SB67, Wei62b, Cry67]. Discriminant [Coc64, Coc68a, Lac66, LM68]. Discriminating [HR65, PW65, QNA65]. Discrimination [Ano64g, BH67, HHW68, HH69, Qur64]. Discussion [BLH^{+}60, Coh66a, Gor66, KFT^{+}60, Tuk61]. discussions [JP61]. Dispersion [LSL68]. Disputes [Tin63]. Distribution [Aig68, AR66, AR69, BA67, BG66, Bar62b, Bar63b, Bar62c, Bra66, Bur65, Bur67, CD67, CB62, CW69, Coh65, CR66, Dod62, Dow66, Eis68, ES69, Ela61, Eps60e, Eps60f, Eps61, Eps64a, Fin63, GF69, GPH67, GD60, Gun65, Gup60, GQS67, GS69, Hal60, HB61, HO63, Har69, Jae64a, Jae65, JL66, Joh62b, Joh63, Joh64b, Joh66c, Kra66b, LSL68, LNN61, Lev66, Lin64, LE64, LMW64, Mac69, Man67, Mar64, Men63, MH69, Ode67, Owe64, Owe65b, Owe67, Sal67, Sär64, SV69, Sat64, SW68, SL63, SM65, TL68, TBA69, TB69, Wei64b, WGH62, Wol77, ZD61, DT68, HM67, Owe68, Pat60, Sho67b, KR68, Ste65]. Distribution-Free [BG66, HO63, LSL68, Wol77]. Distributions [Bas64, BPM63, CB67, Coh66a, Coh67, Dub67a, Eis64, Eps60c, FC60, Fox63, GK69, Gup62, Hal60, Hal68, Hal64, HM68, Has66, LNJE68, Man68, NM69, Pfa63b, Sau68, SB67, SG61, Wei65, WB60, LN66, Rif69, Nau68, Noe63, Pos66]. Divisible [Kap66]. Dobben [Bur69]. Dodge [Dun61, Hal62]. Doig [For66, Zel63a]. Donald [Cow67, Cry67, Gol63, Rei63b]. Doornbos [Nad69b]. Double [Cra68, Dun61, KLTW65b, KLTW65c, ZT68]. Downs [Eis63]. Draper [Bai67, Ano69f]. Dual [HHW68]. Dubes [Wat69b]. Dubins [Ode66b]. Duckworth [Zel62b]. Duncan [Bak64]. Dunn [Gaf65]. Duplicated [Pat63b]. Duplication [Dyk60]. during [BS66]. Dynamic [PJ67, Wou65, Ber61]. Dynkin [Fle63].

Early [Coh66b]. Eckmann [Gol63]. Econometricians [Bas68]. Economic [Tay68, Tra65]. Economically [Owe65b]. Economy [Pet64]. Edgerton [Rog68b]. Editor [Bai63, BC63, BGCR67, Cha67, HTJ^{+}66, JRK^{+}68, JCP64, NF68, Nel60, Owe66b, PO63, Sha65, SH67, Tic62]. Editorial [Ano66f, Smi69a]. Edward [Cra65, GM66]. Edwards [Dyk67a]. Edwin [Gil65]. Effect [Add62a, HB60, Mar68b, Mar69a, Sta64]. Effects [Box63, Eis63, Koc67b, Wei61b]. Efficiency [BH60, EBT62]. Efficient [JL66, STF68]. Eight [Add65, DS68a]. Eight-Run [DS68a]. Electric [Utt66]. Electronic [Rob64]. Element [Jae64b]. Elementary [Bak64, Gei62, Cad69, Hic68, Mad68]. Elements [Cor69, Eps60a, Hag66, Tho67, Tra65, Wal68]. Eliminating [WW61]. Elimination [HO63]. Elliptical [CB62]. Elphyma [S.65a]. Elzey [Li68]. Emanuel [Mil63]. Émile [Hag66]. Emphasis [Owe69]. Emphasizing [Tuk61]. Empirical [Box66b, Man69a, LN66, LNJE68]. Empty [CG62]. Encoding [Pos63]. Enders [Cot69, TC68]. Engineering [Blu66, Cap66, Coh66d, Gra69, Bro68, Dyk65a, Llo61]. Engineers [Bor66, Min65b]. English [Ros65b]. Enrick [Dun60, Zim68]. Equality [Nel68]. Equations [Bas68, GT66, Smi66, Ham62]. Equivalence [Pat63c]. Eric [Chu66a, Ste65, Zel62b]. Erich [Ost64, Ros64]. Erik [S.65a]. Errata [Ano60d, Ano60e, Ano61e, Ano61f, Ano61g, Ano62f, Ano62e, Ano63i, Ano63h, Ano64g, Ano64h, Ano65j, Ano65i, Ano65k, Ano66i, Ano66g, Ano66j, Ano66h, Ano67g, Ano67h, Ano67f, Ano69f, Ano78, BC67, BCK78]. Erratum [Ano68f, Ano73]. Erroneously [Kem67]. Error [BF65, BPM63, CB62, Chu68b, CB67, Coc68a, Con62, DG67, Eps61, GM66, Jae64b, LM68, Mal69, Pos65, Coo66, Man61]. Error-Distributions [BPM63]. Error-Free [Pos65]. Errors [Box63, Coc68b, Kem67, Ker67, MBM64, You62]. Especial [Aig68]. Essays [Nob63, Pfa64]. Essentials [Bar68]. Estimate [Bar63b, Han69, Jae64b, Koc67b]. Estimates [And65, Bas64, Ben64, Coh61b, Dow66, Eps61, Man67, McC65, Tal67, TM63]. Estimating [AC67, BA63, CB67, Coh61a, Ela61, FPS65, GH69, MBM64, Mye64, OM68, Pra67, Sal67, TL68]. Estimation [AC67, Ano65j, Ano65k, Ano68f, Ano73, AH68, BA67, BMPS68, BC66, BC69, Ber69, BM60, Che62, CW69, Coc68a, Coh65, Coh66a, Coh67, Eps60b, Eps60c, GPH67, GS69, HM65a, Har67b, Has66, HHW68, Hin67, Hol67, Hun64, Jae64a, Jae65, Jen63, JS68a, Joh62a, Joh62b, Kip65, Koc67a, Koc68, KA67, LM68, Lar66, Mal69, Man64a, Man68, Mar69a, Men63, NWF63, Par61, Pas62, Pat66, PB67, Pri62, PG65, Sär64, SV69, Sch64, Sch66, SD68, SM65, Tic61, VH69, Was64, BC67, HM67, KMH69]. Estimators [Dub65, Dub67b, Dub67a, Gla62, GQS67, HM68, LJE67, Ler65, SB67, HM65b].

Ettinger [Fei66]. Evaluating [GF69, SD69]. Evaluation [Jae64b, MH69, She65b, Tin63, Pat60]. Even [Web68a]. Evolutionary [Box66b, HK66, SHH62]. EVOP [BD68b]. Exact [AR69, BN64, Har64, Hea69, JL66, LSL68, Rob60, Sal67]. Examination [AT63, De 64, Olm63]. Example [Con62, Jen63, Kem67]. Examples [DH69, Pat60]. exhibiting [Wei61b]. Existence [EM64]. Existing [GM68, Mie67]. Expected [Wil66]. Experiment [Coo68, CHDF67, JA67]. Experimental [Ano63h, Box63, BH65, Bra68, Coh66d, Fry61, Hil60, Jac66, OM68, KMH69, Ham65b, Ham65a, Koe65]. Experimentation [TS64]. Experiments [Add61, Add62a, Ano60d, Ano66h, AH68, Atk69, Bro61, Dan60a, Dan63, Dyk66b, GF69, Gor66, Hew68, HB69, KS69, Mar67, MA66, Sta64, TM68, Zel59, Zys65, Ams65b, Ire65, McK68b]. Explanation [Pro63]. Exponential [Ano65j, AR66, Dod62, ES69, Eps60c, Eps60e, Eps60f, Eps61, GM62, Har64, Hol67, PO61, Sal67, SV69, Sch64, TL68, ZD61]. Expressions [Ros65b]. Extension [Coo68]. Extrapolation [Kru69, LJ69b]. Extreme [Dia67, Dow66, Eps60a, Gor66, GM66, HM68, Man68, MA66, Pos65, HM67]. Extreme-Value [HM68, Man68, Pos65, HM67].

F [Ban69, De 64, Dun61, Gho67, Gil63, Gil65, KR68, Lei68, Li68, McK69c, Olm63, Pet64, Rob64, Sat65, Van64]. Factor [Box63, DS68a, Mar68b, Mar69a, Pot62b, Pot62a]. Factorial [Add61, Add62a, Add62b, Add64b, Add69, Ano60d, Ano63h, Ano65k, BH61a, BH61b, Coo68, CHDF67, Dan60a, Dan63, DW66a, DS68a, Fry61, HB69, Hun64, JA67, Kao69, Mar67, Mar69a, Mar69b, Pat63b, Rec67, Sta64, Web68b, Zel59, Zys65, Cox92]. Factors [Add64a, Dan63, Dra62, DS68b, Man64a, WB60]. Failure [Coh66b, Eis63, EP63b, NWF63, PW65, PO61, Pro63, Sau68, ST68]. Failures [Jae64a]. Fallacious [Goo62]. Families [Cla67]. Fatigue [Ams65b, LJ69b]. Federal [Utt66]. Feedback [Box66b]. Fehr [Gho67]. Feller [Gru67, Sam69]. Ferdinard [Pet64]. Ferguson [Bla66, Rog68a]. Fill [Rob69]. Filters [Jae67, Rob65]. Finding [Ano63h, Fry61]. Finite [Bec67, Cry67, Wal69, Ham62]. Finite-Difference [Ham62]. Finney [Owe65c, Ano63i]. First [Mai66, ST68]. Fisher [Ano60e, Rob60]. Fisk [Bas68]. Fisz [Gei64]. Fit [Wei66]. Fitting [Dol60, GT66, GD60, Man69a, SK65, Har61b]. Five [WM61]. Fletcher [Owe63]. Flow [DGR61, Enr60]. Fluctuations [NH69]. Folded [Ela61, Joh62b, Joh63, LNN61]. Formalization [Rei63a]. Forman [Dyk67a]. Formula [GH69]. Formulae [Dyk67b]. Formulas [Owe65a]. Forsythe [Ham62]. Fortran [Wol68]. Foundations [Jef63]. Four [Ano68f, DeC65, Dia67, DS68b, Har67b, Pot62a]. Four-factor [Pot62a]. Four-Moment [DeC65]. Four-Parameter [Ano68f, Har67b]. Fourier [NH69]. Fourth [All63, Bar63c]. Fraction [FPS65, PB67]. Fractional [Add62b, Add69, BH61a, BH61b, Dan60b, Pat63b, TL68, You61, Dan63]. Fractions [Add61, Ano63h, Fry61, Rag65, Rec67]. Francis [Cow67]. Frank [Ams65a, Dun66]. Franklin [Zys64]. Franz [Ter65]. Fred [Coh66d]. Free [BG66, GM62, GM66, HO63, HB69, LSL68, Pos65, Wol77]. Freedom [Che67, GH69]. Freeman [Min65a]. Frequencies [Ku63b, Ren69]. Frequency [Aig68, GKKK61]. Freund [Bor66, Hag66, Tho63]. Friedlander [S.65b]. Fröberg [Mee66]. Front [Ano60f, Ano60g, Ano60h, Ano60i, Ano61h, Ano61i, Ano61j, Ano61k, Ano62g, Ano62h, Ano62i, Ano62j, Ano63j, Ano63k, Ano63l, Ano63m, Ano64i, Ano64j, Ano64k, Ano64l, Ano65l, Ano65m, Ano65n, Ano65o, Ano66k, Ano66l, Ano66m, Ano66n, Ano67i, Ano67j, Ano67k, Ano67l, Ano68g, Ano68h, Ano68i, Ano68j, Ano69g, Ano69h, Ano69i, Ano69j]. Fry [Blu66]. Fuel [Jae64b]. Function [BN64, Ber69, Bur67, Coc64, Jen63, Nád69a, SL63, McK68a]. Functions [CB67, DL67, HO63, Man62, MH69, Owe65a, PW65, Tic61, Wei65, Har61b, Ros65a]. Fundamental [Ire65]. Further [Koc68]. Future [Wil66].

G [Ams65b, BP68, Bro68, Buc67, Cad69, Chu68a, Cro68, Dun61, Dyk65b, Dyk67b, Fer66, For66, Gho67, Gre68, Hay68, Hol68, Joh68, Kan61, Koe65, Lau68, Lei68, Lev66, Lut68, McL69, Nau68, Ode68, Pfa63a, Rei63a, Roo63, Ros65a, Roy67, Ver66, Wal68, Wou63, Wou65, Zel62a, Zel63b, Zel63a]. G. [Gur65]. Gain [LJ69b]. Gamble [Ode66b]. Gamma [Ano68f, Ano73, CW69, GD60, Gup60, HM65a, Har67b, KR67, Pas62, Sär64, SM65, WGH62]. Gangolli [Cry67]. Ganguli [Cha62a]. Gardner [Mee69]. Gasoline [BM60]. Gauging [Pag62]. Gauss [Har61b]. Gaussian [Pos66, Hin67, Tay67, TS66, WR69]. Geary [Tra65]. Geiringer [Ber66]. General [Dan69, DGR61, Jen61, Koc67a, Koc68, Par64, Pot62b, Pot62a]. Generalized [Ano68f, Har67b, Mal69, SM65]. Generating [Mar64]. Generation [Fox63, SS62]. Genesis [Eis64]. George [Bla66, Ham62, Hus69, Nob63]. Gilbarg [Nob63]. Given [Hal67, JS68b]. Gnedenko [Gei62, Sys63]. Goldstein [Ver66]. Good [McC65]. Goodman [JP61]. Goodness [Wei66]. Gotkin [Ver66]. Graph [Har67a]. Graphical [CL67, Dol60, Ren69, San68]. Graphs [Owe65a]. Graybill [Sat65, Zys64]. Greco [Pot62a]. Green [Van64]. Greenberg [Zel63b]. Greenwood [All62, Owe63]. Grenander [Koo64]. Griffiths [Wat69a]. Grinding [Moh69]. Grossman [Gho67]. Group [Ano62a, Kap66, Pat62, Rei63a, She65b, SG66, Wat61, Gur65]. Group-Screening [Pat62]. Group-Testing [SG66]. Grouped [Aig68, Tal67]. Groups [Jae65]. Growth [BS66]. Gruenberger [Rob64]. Guenther [Low65]. Guide [Thr64, Zel62b, Owe63, All62]. Gupta [Ano60e]. Guterman [Ano62f, Ano63i]. Guttman [Cap66].

H [All62, Bai67, Ber67, Cha62a, Coh66c, Dod60, Dod62, Dra63a, Fer66, Gho67, Gov66, Ham65b, Ham65a, Har62, Kra66a, Ku63a, Man64c, Owe63, Rog68b, S.65b, Ste65, Sto61, Wei62a]. H108 [Eps64b]. Hadley [Gre68, Hol68, Wou63, Wou65]. Hahn [Bro68]. Haight [KR68]. Half [Rif69]. Half-rectified [Rif69]. Halperin [Ano61f, Ano64h]. Hamilton [Sch66].

Handbook [Dod60, Dod62, Eps64b, Buc67, Ode68, All68, Dav66, KR68, Noe63, Owe65a, Rei63b]. Hannan [Hea61]. Hansen [Bak64]. Harald [Tuc65]. Hard [Hin67]. Harmonic [JL69]. Harold [Dun61, Min65a]. Harris [Ney65, Str69]. Harry [Dun61]. Harter [Ano64h, Ano67h]. Hartley [All62, Owe63]. Hasselblad [Coh66a]. Having [FC60]. Hazard [Eps64a, Gav63, Nád69a]. Height [Pas62]. Henry [Cra65, Dem60]. Herbert [Rei62]. Hicks [Ire65]. High [Ano61d, Ham65b, Ham65a, Jos64, Leo61a, Leo61b, Leo61c, Leo61d, Leo62a, Leo62b, Leo62c]. Highway [Pet64]. Hilda [Ber66]. Hillier [Lei68]. Hines [Neu68]. Histories [GKKK61]. Hodges [Put65]. Honor [Nob63]. Hooke [Con63]. Hope [Hic68]. Horst [Owe66a]. Howard [Ran69, Ber61]. Hsu [Owe65c]. Huitson [Pot67]. Hunter [Ams65a]. Hypergeometric [Gue69, Hal60, SZ68]. Hypotheses [BK68, Har66]. Hypothesis [Gho67, OF69, Sch66, Rif69].

IBM [Chu66a, Rob64]. Identical [EP63a, Pre69]. Identification [Kip65]. Identities [Gan69]. If [Ode66b]. II [BCK78, BCK61, BH61b, BMW66, Eps60f, Hal66, HM68, LJ69a, LNJE68, Smi68b, Coo66, Dav66, Gru67, Kor68]. III [Bar63c, Har69]. Impact [Lie65]. Improving [WW61]. Inadmissible [You64]. Inclusion [DS66]. Incomplete [AB64, Doe65, Joh61, McK68a, Pre69, Rob61, Wei65]. Incurred [Kem67]. Independent [Bar62b, BT62, OM68, WS68]. Index [Ano66i, Owe63, Ram66]. Indices [Kra63a]. Industrial [Hon66, Cha62a]. Inequalities [DeC65, Ode66b]. Inequality [Bar62a, Sum65]. Inference [Joh66b, Pyk63, Gei66, BP68, Con63, Jef63, Min65a]. Inferences [TBA69]. Inflows [Llo63]. Information [Ano60j, Ano61l, Ano62l, Ano63n, Ano64m, Ano65p, Ano66o, Ano67m, Ano68k, Ano69k, ES69, Kor68]. Ingelstam [S.65a]. Initial [Lac66]. Inner [Won61]. Inspection [Dun61, Ell63, Hal60, Hal62, PJ67, QNA65, Wet60, Dod60, Dod60]. Instrumental [Mal69]. Integration [Eva67]. Inter [Kap66, Mil60]. Inter-Plant [Mil60]. Interaction [Mar68b, BK68]. Interactions [DeG68a, Mar69a]. Interference [Jae67]. Interim [Dod60, Dod62]. Interlaboratory [Kra67]. Intermediate [McK69d]. International [Cha62b, S.65b]. Interpretations [Goo62]. Interval [Aig68, Bow68, GL67, Man68, Ros66, San68, HM65b]. Intervals [Che68, Goo65, Hal66, Hew68, Kra63a, Kru66, MS68, QH64, SZ68]. Intra [Kap66]. Intraclass [Han69]. Introduction [Coh66c, Fer66, Gei62, Gho67, Gru67, Ku63a, Li68, Neu68, Rou68, Sam69, Utt66, Van64, Zys64, Cry67, Gei66, And66, Kan61, Con63, Cra65, Dra63a, Gei66, Gre68, Hol68, Hus69, Kei63, Lau68, Lei68, McK69b, Mee66, Min65a, Rob64, Roo63, Sat65, Tan68, Tuc68]. Introductory [Bas68, Cap66, Eri66]. Invariant [Man67]. Inventory [Lut68, SG61, Wig67]. Inverse [Ano67f, Bar67a, Hun66, Kru67, Kru69, MH69, Pat63c, WR69]. Inversion [Ano62e, Dwy64, SGR62]. Investigation [FVM69, GS69, Law66, Noe63]. Involved [Pat66]. Involving [BPM63, Ku63b]. Irene [Owe65a]. Ireson [Buc67]. Irregular [Add61]. Irwin [Cap66]. Isn’t [BD68b]. Isotope [Wei61a]. Item [Kra63a, Rei62]. Items [WW61]. IV [All63, Mar69b].

J [All62, All68, Bar63c, Bor66, Bro68, Bur68, Cha62b, Chu68a, Cow67, Cro68, De 64, Fei66, Fer66, Gan69, Gil66, Har62, Hea61, Hus69, Jef63, Kor68, Lei68, Lev66, Man64c, Ode66b, Olm63, Owe63, Owe65c, Put65, Ran68b, Ran68a, Rei63a, Rob64, Roo63, Tho67, Tuc68, Wei62a]. James [McK69c]. Jean [Gaf65]. Jerome [Owe65b]. John [Cow67, Dav66, De 64, Hag66, Ham65b, Ham65a, Jac66, Noe63, Olm63, Rei63a, Tho63, Wat69a, Cox92, Smi69b]. Johnson [Coh66d, Ams65b, Dyk65b, McK68a]. Joint [HW66, HHW68, WS68]. Joseph [Rei63a]. Joshua [Doo69]. Journal [S.65b]. Jump [Mod62].

Kamat [Lev66]. Karl [McK68a]. Karlin [Wei62a]. Kemeny [Rei63a]. Kemperman [Har62]. Kendall [For66, Hay68, Pfa63a, Zel63a]. Kenneth [Pos66]. Khinchin [Gei62]. Klerer [Ode68]. Known [FC60, HM65b]. Korn [Ode68]. Kotz [Ros65b]. Kovary [Fle63]. Kurty [Rog68b]. Kurtz [Ans65, Dav64, Jac65].

L [Coo66, Cra65, Cry67, Dan61, Dra63a, Dun60, Eri66, Gho67, Gor66, Gra69, Jef63, Kra66a, McK68a, Mee69, Ode66b, Olm62, Owe63, Par64, Put65, Ter65, Thr64, Zac68, Zim68]. Laboratory [Cam69, Hsi66]. Laha [All68, Ros65a]. Lajos [Jen62, Kei63]. Lambe [Dyk67b]. Large [QH64, Rog69b, SB67]. Larry [Ams65a]. Lassar [Ver66]. Latin [BCK78, BCK60, BCK61, Cla67, Pot62a, Pot62b]. Latscha [Owe65c]. Lattice [GH62]. Laurie [Rei63a]. Lavi [Hes68]. Laws [Rog69b]. Leadbetter [Ber67]. leading [Wil65]. Learning [Fer66, SV68]. Least [Lar66, Sch66, Har61b, Man69b]. Lee [Kip65]. Lehmann [Put65, Sho67b]. Leimkuhler [Pet64]. Length [Bar67b, Blu67]. Lengths [Ano61e, GW61]. Leo [Ver66]. Leonard [Ams65b, Bak64, Dyk65b]. Leone [Ano80, Coh66d]. Letter [Bai63, BC63, Owe66b, Sha65, Tic62]. Letters [BGCR67, Cha67, HTJ^{+}66, JRK^{+}68, JCP64, NF68, Nel60, PO63, SH67]. Level [Add64b, Add69, Dod60, DS68a, HB69, BB60]. Levels [Box63, Cla65, DS68b, Dan63]. Lieberman [Lei68]. Life [Ano60d, AR66, BG66, Bar63b, Bas64, BCM62, Blu67, Che62, Coh63, Coh66b, Dod62, Dyk66a, EN65, Eps60b, Eps60d, Eps60e, Eps60f, Eps61, Eps64a, EM64, Gup62, Hew68, ML60, Nel63, Ril62, Sau68, Sho67a, WW61, Zel59, ZD61]. Likelihood [Ano68f, Ano73, CW69, CB67, Coh61b, Coh65, GPH67, GD60, Han69, HM65a, Har67b, HM68, Joh62b, Ler65, Tal67]. Limit [Hal62, Tin63, WB60, Tuc65]. Limits [Boh68, FW68, GH66, GM62, HO63, HW66, Mac69, Mad65, Mos65, ST68]. Lindgren [Lau68, Gei63]. Lindley [Gei66, Gei66]. Line [Ber69, Dol60, Dyk67a, Joh66d, Lie61, Mal69, PJ67, Neu68]. Linear [Aig68, Ber69, Coc64, DW66a, Dow66, DS66, DS68a, GH69, GQS67, Jen63, JS68a, Man64a, Man67, McC65, Par64, Pas62, Roo63, SD68, Tra65, Zys64, Har61b, Joh66b, Wou63]. Linearity [GM65]. Lines [Dun62, Pag63]. Link [Ans65, Jac65]. Liquid [Mye64]. Literature [Fis67, For66, Zel63a, HH66b]. Lives [WW61]. Load [HS66]. Locating [Dan60a, Mod62]. Location [Dub67a, LB69, Moh67]. Loève [Kaw65]. Loewner [Nob63]. Log [GK69, Nád69a, Whi69]. Log-Normal [GK69, Nád69a]. Log-Weibull [Whi69]. Logarithmic [Lev66]. Logistic [GK69, GQS67]. Lognormal [Gup62, JS68b]. Long [McK69b]. Loop [Jen63]. Lot [Jae64a]. Lots [Hil62, Wal69]. Lowell [Min65b]. Loxston [Chu68a]. LTPD [Hal62]. Lukacs [Ros65a].

M [All68, Ber67, Chu68a, Cor69, Gei64, Gra69, Hay68, Jen62, Kaw65, Ode68, Owe65c, Owe66a, Pfa63a, Pfa64, Rou68, Roy67, Sat65, Tan68, Tra65, Tuc68]. Machine [Rob63]. Machining [MBM64]. Madison [Str69]. Main [Add62a, Mar68b, Mar69a, Sta64]. Main-Effect [Add62a, Mar68b, Mar69a, Sta64]. Maitland [Cha62b]. Making [Ren69, TSRR67]. Management [Wei62a, Hus69, Lut68]. Manager [Thr64]. Mandel [Jac66]. Mann [Sho67b]. Manual [Dan61, Dun60]. Manufacture [Jae67]. Manufacturing [Dun60, Mil60]. Markov [Ber61, Fle63, Har62, Pyk63, Ran68a, HS66, KKK62, Doo69]. Marks [Aig68]. Marliss [Zel62a]. Martin [Ran68a]. Marvin [Dre63]. Mary [Koe65]. masse [See68]. Material [DGR61]. Materials [McK69d, Pet64]. Mathai [Tan68]. Mathematical [All62, All63, Bar63c, Coh66c, Fer66, Gei64, Ku63a, Luk63, Luk64a, Luk64b, Nob63, Owe63, Owe65a, Par61, Wou64, Ams65a, Ber66, Boo62, Rei63a, Rog68a, Tho63]. Mathematics [Str69, Gil65, Ham65b, Ham65a, Man64c, S.65b, Tan68, Chu68a]. Matrix [Ano62e, Dwy64, Goo69, Lar66, SGR62, Pet66]. Matter [Ano60a, Ano60b, Ano60c, Ano60f, Ano60g, Ano60h, Ano60i, Ano61a, Ano61b, Ano61c, Ano61h, Ano61i, Ano61j, Ano61k, Ano62b, Ano62c, Ano62d, Ano62g, Ano62h, Ano62i, Ano62j, Ano63a, Ano63b, Ano63c, Ano63d, Ano63j, Ano63k, Ano63l, Ano63m, Ano64b, Ano64c, Ano64d, Ano64i, Ano64j, Ano64k, Ano64l, Ano65b, Ano65c, Ano65d, Ano65l, Ano65m, Ano65n, Ano65o, Ano66b, Ano66c, Ano66d, Ano66e, Ano66k, Ano66l, Ano66m, Ano66n, Ano67b, Ano67c, Ano67d, Ano67e, Ano67i, Ano67j, Ano67k, Ano67l, Ano68b, Ano68c, Ano68d, Ano68e, Ano68g, Ano68h, Ano68i, Ano68j, Ano69b, Ano69c, Ano69d, Ano69e, Ano69g, Ano69h, Ano69i, Ano69j]. Maurice [For66, Zel63a]. Maxima [Cho62]. Maximisation [Atk69]. Maximum [Ano68f, Ano73, CW69, Coh61b, Coh65, GPH67, GD60, Han69, HM65a, Har67b, HM68, Joh62b, Ler65, Ode67, Tal67]. Maximum-Likelihood [Ano68f, Ano73, HM65a, Har67b, HM68]. Maxwell [Kru63]. May [VH69]. McCarthy [Tra65]. McCracken [Rob64]. McElrath [Lau68]. McKay [Tho67]. McLean [Gor66]. Mean [AR69, Bin68, Eis68, Eps61, GH69, GS69, Koc67b, Kra67, Kru66, NG68, Pat66, Sat64, Tip63, VH69, WW61]. Means [Bec67, Ben64, Cra62, JS68b, MH68]. Measurement [Coc68b, Jae64b, Tin63, Cro68]. Measures [GM65, Lig69]. Measuring [Wei66]. Mechanisms [BH65]. Mechanistic [BH67]. Median [Bar67b]. Medicine [All63]. Meeting [Ano64a, Ano65a, Ano66a, Ano67a, Ano68a, Ano69a]. Memoriam [Smi69b]. Messrs [JP61]. Messrs. [BLH^{+}60, KFT^{+}60]. Meteorology [Bar63c]. Method [And60, Ano62a, Ano62e, Aro68, BH62, Dol63, Dwy64, Man69a, PW65, Ren69, Rob60, SGR62, See68, SK65, Smi66, VT68, Wat61, Wat69a, Wel62, Gur65, Har61b, Mul65]. Methodology [Dyk65a, Llo61, HH66b]. Methods [All68, AR69, BCK60, BCK61, Cox65, Dyk67b, Ela61, Ham62, Joh68, Kru67, Kru69, Mar67, Mye64, MS68, NH69, Wil69, Zac68, BCK78, Cad69, Mul65]. Metric [Lig69]. Metropolis [Ham65b, Ham65a]. Meyer [Eri66]. Michael [Ste65]. Mijares [Har66]. Miller [Bor66, Owe63, BP68, Gov66, Pos66]. Milton [Owe65a]. Minimum [Gla62, KMH69, SV69, Wig67]. Misclassified [Ano66g, Coh60, Lac66, Pha68b]. Mises [Ber66]. Missing [Dra61, Nad67, Smi68b, Zys65]. Mixed [SD69, TL68]. Mixture [Coh66a, Gor66, Has66, MA66, Mye64]. Mixtures [Coh67, Dia67, Kno67, Sch69, TM68]. Model [BC69, BH62, BM60, CJ64, Fra68, Han69, HHW68, HH69, Hol67, PL65, SV68, Wig67, LJ69a]. Model-Building [BH62]. Modeling [HH66a]. Models [BH67, BG78, Bro68, Dia67, HR65, Koc67b, Neu68, Rei63a, Zys64, Jen62]. Modified [Ano62e, Gue69, SGR62, Har61b, BMW66]. Mohan [Ano67g]. Moment [DeC65]. Moments [Ela61, Gov63, KR67, ML60, Noe63, TL68, Whi69]. Monte [Mul65, BMW65, NG68]. Mood [Sat65, LSL68]. Morethrough [Fei66]. Morris [Rei63a, Ter65]. Motivation [Cha62a]. Multi [BES61, Dod60, Sch69, BK68]. Multi-Component [BES61]. multi-dimensional [BK68]. Multi-Level [Dod60]. Multi-Sized [Sch69]. Multichannel [Cot69]. Multicomponent [GH62]. Multidimensional [CL67, Pos66]. Multinomial [Goo65, QH64]. Multiple [Bec68a, Bra68, Cox65, Dun65, Dun64, GM68, Gup65, Hin68, Ker67, KLTW65b, KLTW65c, Ren69, RS65, Rob61, She65a, TSRR67, Wei61a]. Multiplex [Eva63]. Multiply [Bec68a, Bec68b]. Multiply-Classified [Bec68a, Bec68b]. Multisample [BD68a]. Multivariate [Ben64, Boc63, Har66, SD68, Kra66a]. Murray [Cow67]. Must [Ode66b]. Mutually [BCK60, BCK61, BCK78]. My [BD68b].

N [Bai67, Bak64, Ban69, Gho67, Gil63, Ham65b, Ham65a, McK68a, Rog69a, Zim68, LJ69b]. Narula [Ano80]. Natrella [Koe65]. Negative [And65, Bar62c, Bar66, ML60, Ste65, TM63, Pat60]. Nested [AC67, KA67, LJ69a, LNJE68, LN66]. Newton [Har61b]. Neyman [Nob63, Bar63c, SZ68]. Nilsson [Roo63]. No [Jae64a, BK68]. Noether [Wal68]. Noise [GKKK61, Lin64, Par64, Par64]. Nolinear [Ran68b]. Non [EP63a, GM65, JS68a, Kao69, Nau66, Owe69, Par64, Sin66, TM63, Web68a, Hal63, Har61b]. non-central [Hal63]. Non-Identical [EP63a]. Non-Linear [JS68a, Par64, Har61b]. Non-Linearity [GM65]. Non-Negative [TM63]. Non-Normal [Sin66]. Non-Normality [Owe69]. Non-Orthogonal [Web68a]. Non-Parametric [Kao69]. Non-Random [Nau66]. noncentral [Owe68]. Nondestructive [Jae64b]. Nonlinear [BC66, FVM69, Mee69, Tor63, VT68, BC67, Wou65]. Nonparametric [BD68a, Dav66, GM62, Noe63, Pel69, Wal68, Bla66]. Norbert [Dun60, Rob63]. Normal [AR69, Bec68a, Bec68b, BPM63, Bur67, CB62, Coh66a, Coh67, Eis68, Ela61, GK69, GH66, Gun65, Gup62, GS69, Has66, Joh62b, Joh63, Kru66, LSL68, LNN61, LE64, LMW64, Mar64, MPB68, MH69, Nád69a, Owe64, Owe65b, Owe67, SS62, Sin66, Wei64b, WB60]. Normality [Kem67, Owe69, WS68]. Norman [Coh66d]. Note [Ano65i, Bar62c, Ben64, Boh68, Box66a, Cho62, Chu68b, CR66, CL67, DeC65, Hew68, HH69, KR67, Ku63b, LJ69b, Mie67, Pat63a, Rag65, See68, She65a, Sta64, VT68, WW64, Wel62, Wil69, Gur65, HM67]. Nuclepore [Jae67]. Null [SW68]. Number [BM60, CR66, Cra68]. Numbers [MPB68, Rog69b]. Numerical [Eva67, FVM69, Mai66, Mee66, Wes62].

O [All62, Owe63]. Observation [Joh64b, Ros66, VH69]. Observations [Gru69, HW66, Hew68, Kru60, Tip63]. Observed [Dur66, Jae65, JS68b, Nel66, Pro63]. Obtaining [Man67, MS68]. Occupancy [HB61]. Occur [Jae64a]. Octane [BM60]. October [Dod60, Str69]. Off [Chu68b]. Olive [Gaf65]. One [Cla65, FVM69, Har64, Kra63a, OF69, VH69]. One-Dimensional [FVM69]. One-Sided [Kra63a, OF69]. Open [Jen63]. Operate [Wei62b]. Operation [Box66b, SHH62, HK66]. Operational [Cha62b, Zel62b]. Operations [Lei68, Olm62, Thr64, Sew66]. Operators [Fow69]. Optimal [BCM62, Che62, LJ69b, OM68, Owe65b, Kip65]. Optimisation [SHH62]. Optimization [Hes68, Hsi66]. Optimum [Mor61, Sal67, Was64, Wei61a]. Order [Dra62, GPH67, Gov63, Gup60, GQS67, Har64, KR67, Sal67, Whi69, Zel63b, HM65b]. Orthogonal [Add62a, BCK60, BCK61, Mar68b, Mar69a, MD68, Sta64, Web68a, BCK78, Wil65]. Orthogonalization [Ano66h, Dyk66b]. Other [Ano65j, Dub67a, Sch64]. Outliers [AG60, Bro61, Dan60a, GS69, TG67]. Outlying [Dix64, Gru69]. Overdetermined [Smi66]. Owen [Ano66j, Rei63b].

P [Bas68, Gol63, Hes68, Lev66, Nau68, Owe63, Owe65c, Rei63a, Smi69b]. Paired [TSRR67]. Pairs [Ste60]. Pál [Rog69b]. Pala [Neu68]. Paper [Ans65, Jac65, Pat63a, Gur65]. Papers [Ano64a, Ano65a, Ano68a, Ano69a, Ban69, KFT^{+}60, BLH^{+}60]. Papoulis [Mil66]. Parallel [Dan60b, Mad65]. Parameter [Ano68f, AR66, AH68, Coh61a, Dub67a, Eps60c, GM62, Har64, Har67b, HHW68, Jae64a, Man68, SM65, SZ68, Was64, HM65b]. Parameter-Free [GM62]. Parameters [AC67, Ano68f, Ano73, BA67, CW69, Coh66a, Dow66, Dub65, Dub67b, Ela61, Eps60c, GPH67, GQS67, HM65a, Har67b, HM68, Has66, Kao64, Man67, Men63, Sal67, Sär64, TL68, TBA69]. Parametric [Kao69]. Parker [Cow67]. Part [Gei66, BH61a, BH61b, Eps60e, Eps60f, KLTW65b, KLTW65c]. Parthasarathy [Lig69]. Partial [Dyk60, Ham62, Hew68, You61]. Partially [Cla67, CHDF67, JA67, Pat63b]. Particle [Jae65, NM69]. Particles [Jae65, Kno67, Sch69]. Parts [Wig67]. Parzen [Mil63]. Passage [Har62]. Paths [Joh66d]. Patil [Lev66, Nau68]. Patrick [Pyk63, Thr64]. Patterned [Bro61]. Patterns [PL65]. Paul [Eri66]. Peak [MT65]. Pearce [And66]. Pearson [McK68a, Har69]. Peng [McK68b]. Percentage [Bra66, Har61a, Har69, WR69, Har66]. Percentages [Owe64, Rob60]. Percentile [Dub67b, FC60]. Percentiles [Noe63]. Performance [Coc64]. Periodic [Wou66]. Periodogram [Jon65]. Permissible [Dub67a]. Permitting [Mar69a]. PERT [Utt66]. Peter [S.65b]. Peters [McL69]. Peterson [Man61]. Physical [BH65, Coh66d, Man69a, Sch66, You62]. Physics [Bar63c]. Piersol [Cro68]. Pieruschka [Ost64, Ros64]. Pin [Moh67]. Plan [Add65, Pre66, She65b]. Planning [Fle65]. Plans [Add62a, Add62b, Add64b, Add69, BG66, Cra68, DW66a, Gup62, Hal60, Hal62, Hal67, Hal68, Hil64, Mar69a, Owe67, Pat63c, PB67, Pha68a, Pha68b, PZB62, QNA65, Rag65, SD69, Sta64, ZT68, Owe65b]. Plant [Mil60]. Plants [DGR61, Dun60]. Plot [Add64b]. Plots [WGH62]. Point [Eck69, HM65b, Man68]. Points [Bra66, Dol60, FC60, Har61a, Har66, Har69, WR69]. Poisson [KR68, Coh61a, Gla62]. Policies [Dur66]. Policy [DW66b]. Pólya [Nob63]. Polynomial [Cla65, Hoe68b, Kus69]. polynomials [Wil65]. Pooling [HB69]. Population [Ano66g, Ano68f, Coh60, Har64, Har67b, Jen62, Koc67b, Pat66, PO61, HM65b]. Populations [Ano73, Bec67, Bec68a, Bec68b, GH66, Gov63, HM65a, JS68b, Sin66]. Positive [ML60]. Posner [GM66]. Possibility [WW61]. Possible [STF68]. Posteriori [Smi66]. Potential [Doo69]. Power [BN64, LB69, Man62, Nau66, PG65, Sho67a]. Practical [Dun60]. Practice [Bur69]. Precedence [EN65, Nel63, Sho67a]. Precision [LJ69b]. Prediction [Che68, Hew68, Ker67, Lie61, Rei62]. Predictions [Lie61, PW65]. Preference [Smi68a]. Present [Wil66]. Presented [Ano64a, Ano65a, Ano68a, Ano69a]. Pricing [Dur66]. Primer [Gaf65]. Principles [Dun60, Ost64, Ros64, Roy67]. Prior [Hal60, Hal68, Pfa63b, Wei65]. Probabilistic [Gra69]. Probabilities [Gla62, HS66, Koo64]. Probability [All63, Bar63c, Ber66, CR66, Cra65, Dra63a, Eri66, Gei66, Gho67, Gru67, Hog66, Hol68, Lau68, Luk64a, Luk64b, NWF63, Pfa64, Put65, Sam69, Smi66, Ste65, Tho67, Wei62a, WGH62, Cry67, Bar68, Cor69, Cry67, Gei62, Gei66, Hag66, Sys63, Wat69b, Blu66, Bor66, Gei64, Kaw65, Lig69, Luk63, Mil66]. Probit [Pra67]. Problem [DeC65, Gla67, HB61, Har62, HHW68, JB61, Law66, LB69, LE64, Rob64, Wei61a, Wes62]. Problems [All63, And60, Bak64, Bec67, BD68a, Dun62, Eck69, FVM69, Gav63, Jen62, Ran68a, VT68]. Procedure [AB64, Bec67, Dol60, Dun60, Koc67b, Mod62, BMW65]. Procedures [AC67, BMPS68, Bec68b, BA63, Dod60, Dod62, EBT62, Eps60d, Eps64a, FVM69, Gru69, Man68, Pfa63b, Ril62, Rob66, SD69, Was64, ZD61, BMW66]. Proceedings [Str69, All63, Bar63c, Cha62b]. Process [AC67, BMPS68, BD68b, HS66, Hin68, Lie65, Sew66, SG61, Tay67, Wou66]. Processes [Ano64g, Bra66, CJ64, Doo69, Hin67, Jen62, McK69b, Par64, Qur64, QNA65, SD68, SV68, TS66, Ber61, Ber67, Fle63, Gov66, Hon66, Koo68, Mil63, Mil66, Ney65, Ode66a, Ode66b, Pyk63, Rou68, Zac68]. Processing [TC68]. Producer [Sin66, Hal67]. Product [DHT69, Dur66, Fle65, FPS65, HB60]. Production [PJ67, Wil66]. Productivity [Cha62a]. Products [Bar62b, Wel62]. Program [BS66, Jae64b]. Programed [Ver66]. Programme [DGR61]. Programmed [Li68, Utt66]. Programming [Ber61, Boc63, PJ67, Rob60, Tra65, Man69b, Boo62, Ran68b, Wol68, Wou63, Wou64, Wou65, Chu66a]. Programs [Ano61d, Hil62, Jos64, Leo61a, Leo61b, Leo61c, Leo61d, Leo62a, Leo62b, Leo62c, Cot69]. Progress [Olm62]. Progressively [Coh63, GK69]. Properties [Dub65, Fin63, Wou66, Owe68]. Proportion [Que64, SG66]. Proportional [Blu67]. Proportions [Goo65, QH64]. Proposal [Ber69]. Proschan [Ams65a]. Protection [Kus69]. Pseudo [MPB68]. Pseudoinverse [HH66a]. Publications [Ano62k]. Pulse [Pas62]. Pulse-Height [Pas62].

Quadratic [DW66a, Tic61, Wes62, Man69b]. Qualitative [Kru63, OF69]. Quality [Ano61e, Dod60, Dun60, Dur66, Eps64b, GW61, HB60, Hsi66, Jae64b, Kra63a, TS64, Fei66, Dod62, Dun60, Zim68]. Quantal [Wei64a]. Quantiles [Eis68, Sär64]. Quantitative [Hus69, Wei64a, BB60]. Query [And65, Ano67f, BF65, Bar67b, Bar67a, Bow68, Bur65, Cha66, Che67, Coc65, Coh66b, CHDF67, Dix64, Gun65, HW66, JL67, Joh64b, JA67, Kao64, LMW64, Mac69, Man64b, MPB68, Mos65, Nad67, Nel66, Ros66, Sat64, Smi68b, Smi68a, Wei64b, You64, Zys65]. Queues [DeC65, Kei63]. Quick [Dol63].

R [All68, Ams65a, Bar63a, Bas68, Ber67, Blo66, BP68, Cry67, Dyk66a, Fer66, Gei62, Gov66, Ham62, Har67a, Ire65, Kip65, Lev66, Lig69, McK69c, Min65b, Nad69b, Olm62, Owe65c, Par64, Pet66, Roo63, Ros65a, Sto61, Tra65, Wat69b, Wol68]. R. [Gor66]. Radar [TC68]. Radhakrishna [Joh66b]. Radial [CB62]. Radioactive [Pet64]. Radionuclide [PL65]. Rahman [Rog69a]. Raiffa [Ran69, Sto61]. Rainfall [GF69]. Rall [Coo66]. Ralston [Mai66]. Random [Blo66, Cro68, Fox63, Gav63, Joh66d, Koc67b, Koo68, MPB68, Mil66, Nau66, Ode66a, Par64, SS62, You64]. Randomisation [DS66]. Randomized [HS66]. Randomness [Man62, Noe63]. Range [Bar62a, Bur65, CD67, Har61a]. Rank [Dun64, Mie67, Mos65, She65a, Ste60, TSRR67, BMW65, BMW66]. Ranked [HW66, Joh64b]. Ranking [Bec67, Bec68b, Gup65]. Ranks [Ode67]. Rao [Joh66b]. Rate [BF65, Eps64a, EP63b, Pro63, Sau68]. Rates [Coc68a, EP63b, LM68, TS66]. Ratio [Ben64, CR66, GL67, MT65]. Ray [Pas62]. Reactor [Jae64b]. Readings [Wou64]. Reappraisal [Jon65]. Received [Ano62k, Ano63e, Ano63f, Ano63g, Ano64e, Ano64f, Ano65e, Ano65f, Ano65g, Ano65h]. Reciprocals [Tip63]. Rectangular [BPM63]. rectified [Rif69]. Reduced [WM61]. Redundant [Hal64, ST68, Wei62b]. Reference [Aig68, TC68]. Regession [LJ69b]. Region [Nád69a]. Regions [Lie61]. Regression [Bai67, Bow68, Box66c, Chu68b, Cla65, DS66, FVM69, GM68, Hin68, HL67, JL67, JS68a, Ker67, Kra63b, Kru67, Kru69, Ler65, Lie61, Man64a, Man64b, OM68, PL65, SK65, Tor63, VT68, Wes62, Wil69, Har61b, Kus69, Man69b, Mee69]. Regressions [STF68]. Regressor [Ker67]. Rejection [AG60, Dix64, HB60]. Related [Ber67, Cla67, Gov63, Lev66, Nob63, Wol77]. Relations [Mal69]. Relationship [EP63b, SZ68]. Relationships [Gov63]. Relative [Eps61]. Relay [PO61]. Release [Rob69]. Reliability [BH60, BS66, Bas64, BES61, Dod62, Eps64a, Eps64b, Gav63, Gra69, Hol67, JL66, LE64, Mad65, MS68, PW65, Pra67, SV69, ST68, Wal69, Wei61b, Ams65a, Dre63, Ost64, Ros64, Zim68, Buc67]. Reliable [Eps64a]. Remark [Cox65]. Remarks [Dan69, Koc68, Kru60, SB67, Wet60]. Renewal [Sch69, SL63, Bar63a]. Repair [Hal64]. Repairable [Fle65]. Repetitive [She65b]. Replacement [Gla67]. Replenishment [DW66b]. Replicates [Dan60b]. Replication [HB69, You61]. Replication-Free [HB69]. Reply [KLTW65a]. Requirement [BM60, HO63]. Research [Gol63, Olm62, Sew66, Str69, Van64, Rei63a, Cha62b, Dyk65b, Lei68, Thr64, Zel62b, Ban69]. Reservoirs [Llo63]. Reset [Moh69]. Resetting [MBM64]. Residual [JL67]. Residuals [AT63, De 64, Ler65, Olm63, PL65, TG67, dB64]. Resolution [Add65, Mar69b, Pre66, Web68a, WM61]. Response [Chu66b, Dra61, Dra63b, DS68b, Dyk60, GKKK61, Mod62, TM68, HH66b, KMH69]. Responses [JA67, Wei64a]. Restricted [Man69b]. Results [Ell63, Mar69b, KLTW65b]. Révész [Rog69b]. Review [All62, All68, All63, Ams65b, Ams65a, And66, Bai67, Bak64, Ban69, Bar63a, Bar63c, Bar68, Bas68, Ber61, Ber66, Ber67, Bla66, Blo66, Blu66, BP68, Boo62, Bor66, Bro68, Buc67, Bur68, Bur69, Cad69, Cap66, Cha62a, Cha62b, Chu68a, Chu66a, Coh66d, Coh66c, Con63, Coo66, Cor69, Cot69, Cow67, Cra65, Cro68, Dan61, Dan63, Dav64, Dav66, Dem60, Dod60, Dod62, Doo69, Dra63a, Dre63, Dun60, Dun61, Dyk65a, Dyk65b, Dyk66a, Dyk67a, Dyk67b, Eps64a, Eri66, Fei66, Fer66, Fis67, Fle63, For66, Gaf65, Gan69, Gei62, Gei63, Gei64, Gho67, Gil63, Gil65, Gil66, Gol63, Gov66, Gra69, Gre68, Gru67, Hag66, Ham62, Ham65b, Ham65a, Har67a, Har62, Har66, Hay68, Hea61, Hes68, Hic68, Hog66, Hol68, Hon66, Hus69, Ire65, Jac66, Jef63, Joh66b]. Review [Joh68, Kan61, KR68, Kaw65, Kei63, Kip65, Koe65, Koo64, Koo68, Kor68, Kra66a, Kru63, Ku63a, Lau68, Lea67, Lei68, Lev66, Li68, Lig69, Llo61, Low65, Luk63, Luk64a, Luk64b, Lut68, Mad68, Mai66, Man61, Man64c, McK68b, McK68a, McK69d, McK69a, McK69b, McK69c, McL69, Mee66, Mee69, Mil63, Mil66, Min65b, Min65a, Mul65, Nad69b, Nau68, Neu68, Ney65, Nob63, Noe63, Ode66a, Ode66b, Ode68, Olm62, Ost64, Owe65a, Owe65b, Owe65c, Owe66a, Par64, Pet66, Pet64, Pfa63a, Pfa64, Pos66, Pot67, Put65, Pyk63, Ran68b, Ran68a, Ran69, Rei62, Rei63b, Rob64, Rob63, Rog68a, Rog68b, Rog69b, Rog69a, Roo63, Ros64, Ros65b, Ros65a, Rou68, Roy67, S.65a, S.65b, Sam69, Sat65, Sch66, Sew66, Sha66, Ste65, Sto61, Str69, Sys63, Tan68, Ter65, Tho63, Tho67]. Review [Thr64, TC68, Tra65, Tuc65, Tuc68, Utt66, Van64, Ver66, Wal68, Wat69a, Wat69b, Wei62a, Wol68, Wou63, Wou64, Wou65, Zac68, Zel62b, Zel62a, Zel63b, Zel63a, Zim68, Zys64, HH66b, HK66]. Reviews [Cry67, Gei66, Jen62, Owe63, Rei63a]. Revisited [DH69, Smi68a]. Richard [Ber66, Cow67]. Ridge [Dra63b]. Right [Coh61a]. Riordan [Gan69]. Risk [Hal67]. Risks [Sin66]. Rival [HR65]. Rivett [Thr64]. Robert [Con63, Lut68, McK69c, Ran69]. Roberts [Gei62]. Robin [DeG68b]. Robinson [Cot69, TC68]. Robust [DW66a, LJE67]. Robustness [Pos63, ZD61]. Roessler [Cra65, Dra63a]. Roger [Cry67]. Rohlf [McK69c]. Romig [Hal62, Dun61]. Ronald [Ber61]. Root [Ano62e, Dwy64, SGR62]. Roots [Har66]. Rosenhead [Owe63]. Rotatable [Dra62, MD68]. Round [Chu68b, DeG68b]. Round-Off [Chu68b]. Routine [TS64]. Roy [All68]. Rozanov [Koo68]. Rubinoff [Ter65]. Ruiz [Neu68]. Ruiz-Pala [Neu68]. Rule [OF69]. Rules [Gup65, GS69, Lut68]. Run [Ano61e, DS68a, GW61]. Runs [Bar67b, PZB62]. Russell [Thr64]. Russian [Ros65b]. Ryser [Man64c].

S [And66, Bak64, Boo62, Bro68, Bur69, Cap66, Chu68a, Cow67, Cro68, Dan63, Dyk67a, Gur65, Hon66, Jen62, Lei68, McK68a, McL69, Pet66, Pfa64, Pos66, Rog68a, Rou68, Sew66, Ver66, Wei62a, Wou64, Zel62a, LJ69b]. S-N [LJ69b]. Safe [Sau68]. Safety [Pet64]. Sample [Ano67f, Bar67a, CR66, Cra68, FW68, Gut62, HB60, Hoe68a, Joh62a, Kru66, LB69, McC65, NG68, Pat66, QH64, Sär64, SB67, Sum65, TB69, Wet60, Zel62a, BMW65, BMW66]. Samples [Ano68f, Ano73, Cha66, Coh61a, Coh61b, Coh63, Coh65, Eps60c, Fox63, GK69, Gov63, Gru69, GS69, Hal66, HM65a, Har67b, HM68, KA67, Lac66, Nad67, Sal67, Smi68b, WS68]. Sampling [Ano67f, BG66, Bar63b, Bar67a, Bec67, Blu67, Bur65, Cox60, Cra68, DG67, Dod60, Dod62, Dun62, ES69, Ell63, Eps64a, Eva63, Gue69, Gup62, Hal60, Hal62, Hal67, Hal68, HB60, Hil64, JB61, Kno67, LN66, LNJE68, Mad68, Owe65b, Owe67, Owe69, Pat63c, Pfa63b, PB67, Pha68a, Pha68b, PZB62, QNA65, Sch69, SD69, She65b, Sum65, TS66, Wei65, ZT68, Dyk67a, Rif69, Dun61, Dun61, Eps64a, Har67a]. Samuel [Ros65b]. Sarhan [Zel63b, Ano62f]. Satellite [DW66b]. Satterthwaite [GH69]. Saturated [Rec67, Web68b]. Savage [Jef63, Ode66b]. Savas [Hon66]. Savings [Kao64]. Scale [HM68, Kao64, Men63, HM65b]. Scarf [Wei62a]. Scheffé [Dem60]. Scheme [Wet60]. Schemes [Ano61e, Cox60, GW61, Kem67, Pag62]. Schiffer [Nob63]. Schlaifer [Ran69, Sto61]. Schleifer [Owe65b]. Science [Dyk65a, Llo61, Sch66, Hog66, Wei62a]. Sciences [Kor68, Coh66d, Gaf65, Pet66, Rei63a]. Scientific [Cam69, Con63, Dun60, McK68b, Wat69a]. Scientists [Min65b, Van64]. Scores [Smi68a]. Screening [Ano62a, Dav63, Pat62, Wat61, Gur65]. Seal [Kra66a]. Search [FVM69, Mod62]. Searle [Pet66]. Second [Cha62b]. Sediments [Wat69a]. Segments [Joh66d]. Seismic [TC68]. Selected [Luk63, Luk64a, Luk64b]. Selecting [Hil64]. Selection [GT66, Gup65, HL67]. Self [Wol68]. Self-Training [Wol68]. Semigraphical [And60]. Seminar [Str69]. Sequences [Add69, HS66]. Sequential [Add64a, Ano65k, AR66, Aro68, AR69, BCM62, BMW65, BMW66, CR66, Cra62, DL67, HB60, Hoe68a, Hun64, JB61, Mod62, SHH62, Wal69, Was64, Web68b, Joh68]. Sequentially [JS68b]. Serial [Cox60, TS64]. Serially [Llo63]. Series [Cot69, Goo61, Hea61, Lev66, Mad65, Mar69a, Mar69b, NH69, Str69]. Service [Eis63]. Session [Ano64a, Ano65a, Ano68a, Ano69a]. Sessions [Ano66a, Ano67a]. Set [Dol60, MBM64, Wil65]. Set-Ups [MBM64]. Sets [BCK60, BCK61, Doe65, Gho67, Pre69, BCK78]. Several [Dub65, FVM69, Lie61, WS68]. Shape [Jae64a, Men63, HM65b]. Shapiro [Bro68]. Shirley [Dan63]. Shooman [Gra69]. Short [KLTW65b, KLTW65c, OF69, WW61]. Short-Cut [KLTW65b, KLTW65c, OF69]. Shot [Lin64]. Shreider [Mul65]. Sided [GH66, Kra63a, OF69]. Sign [RS65]. Signals [Par64, TC68]. Significance [BN63, Har66, Hea69, Owe65c, Owe66a, Wei64a]. Significances [See68]. Silverman [Par64]. Simeon [Cor69]. Simple [Box66b, Cra62, Pra67]. Simplex [GH62, SHH62]. Simplified [Dun60]. Simulation [Bur67, DGR61, EBT62]. Simultaneous [BC66, BC67, BC69, Che68, Goo65, QH64, Wei64a, BP68]. Single [Bec68b, Cra68, Dun61, Hal60, Hal62, Hal67, Hal68, KLTW65b, KLTW65c, Lie61, Ros66, Wet60]. Single-Stage [Bec68b]. Singly [Coh61b, HM68]. Singular [Goo69]. Sittig [Fei66]. Size [Doe65, FW68, Hoe68a, Jae65, Joh62a, McC65, NM69]. Sized [Sch69]. Sizes [Ano67f, Bar67a]. Sjoberg [S.65a]. Skorokhod [Ode66a]. Slippage [San68, Nad69b]. Slow [Eis63]. Slow-Downs [Eis63]. Small [Cha66, GS69, LB69, Rei63a, Sta64]. Small-Group [Rei63a]. Smaller [Chu68b, McC65]. Smallest [GPH67]. Smirnov [Zel62a]. Smith [Bai67, Wol68]. Snell [Rei63a]. Social [Rei63a]. Sokal [McK69c]. Solids [Rob69]. Solomon [Hog66, Nob63, Rei62]. Solution [Wet60]. Solutions [Jen62]. Solving [Rob64, Smi66]. Some [Add64b, Bas64, BD68a, BC62, BB60, Cla67, DG67, Dan69, DH69, Dub67b, Dub67a, Fin63, Fow69, Goo69, Goo61, Gup65, Hal62, Hal64, Koc68, Kru60, Lin64, Mie67, MT65, PO61, PZB62, Rob66, SB67, Wet60, Dyk67a]. Spaces [Lig69]. Spare [Mor61, Wig67]. Spearman [BC69]. Special [TC68]. Specified [Fin63]. Spectra [Hin67, Jen61, Par61, Pas62, Pri62, PG65].

Spectral [Goo61, Jon65, Str69, Won61]. Spectrum [LJ69a, Tuk61]. Speed [Ano61d, Ham65b, Ham65a, Jos64, Leo61a, Leo61b, Leo61c, Leo61d, Leo62a, Leo62b, Leo62c]. SPES [Ano64a, Ano65a, Ano66a, Ano67a, Ano68a, Ano69a]. Spherical [DL67]. Split [Add64b]. Splitting [Tin63]. Spurious [VH69]. Square [Ano62e, Bin68, Che67, Cla67, Dwy64, Pot62b, SGR62, Pot62a]. Squared [Mie67]. Squares [BCK60, BCK61, GH69, Lar66, Nel66, Wel62, BCK78, Har61b, Man69b, Sch66]. Stage [AC67, Bec67, Bec68b, KA67]. Stages [Pat62]. Stahl [Gei62]. Standard [Eis68, Gut62, JS68b, Pag63, Ros66, Tip63]. Start [SG61]. Start-Ups [SG61]. States [Str69]. Stationary [GKKK61, Har62, SD68, TS66, Ber67, Koo68]. Statistic [Har64, LSL68, SW68, Web68c, Zel62a, Sho67b]. Statistical [Ams65b, Ano61d, Ano64a, Ano65a, Ano68a, Ano69a, BMPS68, BP68, BM60, Cad69, Cam69, Dan61, EBT62, Eps60d, Eri66, Eva67, For66, HB61, Hol68, Jac66, Jef63, Joh66b, Jos64, Leo61a, Leo61b, Leo61c, Leo61d, Leo62a, Leo62b, Leo62c, Lie65, McK69d, Min65a, MT65, Mul65, Nel68, Ran69, Rei63b, Ros65b, Sto61, Tan68, Tay67, Tin63, Wou66, Zel63a, Zys64, Bro68, Bur68, Dre63, Dyk65a, Dyk66a, Dyk67b, Gei63, Llo61, McK69c, McL69, Pyk63, Rog68b, Sha66, TC68]. Statistics [All63, And66, Bar63c, Bor66, Chu68a, Coc68b, Con62, Cow67, Dan69, Dav66, GF69, Gaf65, Gei66, GPH67, Gov63, Gup60, GQS67, Hay68, Hic68, Hog66, KR67, Lev66, Luk63, Luk64a, Luk64b, Noe63, Pfa63a, Rog68a, Sal67, Str69, Ver66, Whi69, HM65b, Owe63, All62, All68, Bak64, Ban69, Ber66, Cap66, Coh66c, Cra65, Dav64, Dra63a, Gei64, Gre68, Joh68, Koe65, Ku63a, Lau68, Li68, McK69a, Pfa64, Put65, Rog69a, Roy67, Sat65, Tho63, Tuc68, Wal68, Zel63b, Coh66d, Lea67, Min65b, Sch66]. Status [Wil66]. Stegun [Owe65a]. Stepwise [JS68a]. Sterile [Rob69]. Stig [S.65a]. Stimulation [GF69]. Stochastic [Ber67, CJ64, Eis63, Gov66, Ker67, Mar68a, McK69b, Mil66, Ode66b, Rou68, Zac68, Jen62, Jen62, Mee69, Mil63]. Stochastically [Bas68]. Stoller [Sew66]. Storage [Hil62, Mil60]. Straight [Dyk67a, Joh66d, Mal69]. Straight-Line [Dyk67a, Joh66d, Mal69]. Strategic [Bro61]. Strategy [Sew66]. Strategyst [Gil66]. Stratification [DG67]. Stratonovich [Par64]. Strength [LE64]. Stress [LE64]. Structure [EM64, Lev66]. Structures [BES61, EP63a, Koo64]. Stuart [Hay68, Pfa63a]. Student [Owe65b, GB65]. Studentized [Har61a]. Studies [Blu67, LNJE68, Lev66, BMW65, LN66, Nob63, Ode66a, Rei62, Wei62a]. Study [Ano62a, BH65, LJE67, NG68, NH69, Wat61, BB60, Gur65]. Subject [Box66b, CB67, DG67, Mal69]. Subsampling [Bec67]. Subset [HL67]. Successive [Bin68]. Suitable [GF69]. Sum [Bur69, Har66, Joh63, Joh66c, LMW64, Ode67, Pag61, Pag62, Ste60, Tay68, TSRR67, Wei64b, Ewa63]. Summary [Owe69]. Summers [McL69]. Sums [Dun64, She65a, Wel62]. Supersaturated [BC62]. Supplement [Owe66a]. Surface [Dra61, DS68b, Dyk60, TM68, HH66b]. Surfaces [Dra63b, Man69a, KMH69]. Surveillance [Hil62]. Survey [Eck69, HH66b, Owe68]. Survival [Kra63b]. Symbols [Rog68b]. Symmetric [Lar66]. Symmetrical [Add62b].

Symposium [All63, Bar63c]. System [BH60, Box66b, EP63b, EM64, Hal60, MS68, Pra67, Smi66]. Systematic [BC62, Mar67, You62]. Systems [BES61, Eis63, Fle65, GH62, Hal64, Mad65, Mor61, Roo63, ST68, Tic61, Wei62b, Gol63]. Szego [Nob63].

T [Fle63, Gei62, Hes68, Kor68, McK69d, Ney65]. Table [BN63, BN64, Har66, Har69, TSRR67, Owe65c, Owe66a]. Tables [All62, BN63, BG78, Coh61b, CL67, Dod60, Dod62, Dun61, Eps64a, Har61a, Hea69, Ku63b, KKK62, Man67, Nel63, Owe63, Owe66a, RS65, SD69, Sho67b, WR69, WB60, BK68, Lev66, McK69c, Owe63, Owe65a, Rei63b, S.65a, McK68a, Owe65b, Owe65c, Ste65]. Tablets [Rob69]. Tail [Mar64, Sat64]. Tails [Owe64]. Takacs [Jen62, Kei63, Zac68]. Targets [Eck69]. Taro [Mad68]. Taub [Ham65b, Ham65a]. Tchebycheff [Bar62a]. Technique [DS66, Dyk65b, Pra67]. Techniques [Bur68, Eva67, Kra63b, Hes68]. Technometrics [BCK78, Ram66]. Temple [Kan61]. Tensors [Kan61]. Terminated [ST68]. Terms [Bar62a, Rog68b, Ros65b]. Test [BG66, BN64, BCM62, CG62, Eps60b, Eps60d, Fle65, Gup62, Hew68, Kao64, Ku63b, LSL68, Man62, Nel63, Nel68, OF69, PL65, RS65, Ril62, Ste60, Wei64a, Wol77]. Testing [Ano60d, BS66, Bar63b, Bas64, BN63, Coc65, Coh63, Coh66b, Dod62, DS66, EN65, Eis68, Eps64a, Hoe68a, Hoe68b, JS68b, LJ69b, MH68, ML60, Mul65, Owe65c, Owe66a, Sch66, SG66, Zel59, ZD61, Rif69, Gho67]. Tests [AR66, AR69, BD68a, Bur69, CR66, Cra62, Dav63, Eps60e, Eps60f, Har66, Hea69, JB61, Kra67, KKK62, Mie67, Mos65, Nau66, NG68, NWF63, Sho67a, ST68, TB69, BMW65, BMW66, Bur68, Nad69b]. Text [Bas68]. Textbook [Ver66]. Their [BES61, CW69, JB61, Neu68]. Theorem [Cox60]. Theorems [Hal62, Tuc65]. Theoretic [Rog68a]. Theoretical [Pro63, Rog69a]. Theory [Ams65a, Ber66, Bur69, DG67, Dre63, Dyk65a, Dyk67a, Eps60a, Gei62, Gei64, Gov66, Gru67, GM66, Hag66, Hay68, Kei63, Llo61, Neu68, Ney65, Ode66a, Par64, Pfa63a, Pos65, PZB62, Sam69, Sat65, Sch69, Sys63, Wat69b, Zac68, Rif69, Bar63a, Doo69, Fer66, Gei63, Hol68, Kaw65, Mad68, Ran69, Sto61, Tho67, Dyk65b, Fle63]. Third [Dra62]. Thomas [Dav64, Rog68a]. Thornton [Blu66]. Those [WW61]. Three [BA63, Dan63, Dia67, Doe65, Dra62, DS68b, Mye64, Pot62b, BB60]. Three-Factor [Pot62b]. Ties [Mie67]. Time [Cot69, Goo61, Hal64, Hil60, Kao64, Lin64, NH69, Str69, Hea61]. Times [Wei62b]. Tito [Har66]. Todd [Ham65b, Ham65a]. Tolerance [Boh68, Bow68, FW68, GH66, GM62, HO63, Mac69, WB60]. Tolerance-Limit [WB60]. Tolerancing [Eva67]. Tompkins [Ham65b, Ham65a]. Too [BD68b]. Tool [BG78, SV68]. Topics [Nob63, McK69d, Par64]. Total [HB60]. Tou [Kor68]. Training [Wol68]. Transfer [Jen63, Tic61]. Transformation [BT62, Chu68b, Dol63]. Transformations [Blo66, DH69, MD68, Par64]. Transformed [Pat66]. Transition [HS66]. Translations [Luk63, Luk64a, Luk64b]. Transport [Pet64]. Treatment [Ams65b]. Treatments [RS65, Ste60]. Trend [Bla66, Man62]. Trends [DW66a, DS68a, Hil60]. Tribute [Dun66]. Trimmed [Cha66]. trimming [DT68]. trimming/Winsorization [DT68]. Trucking [Pet64]. Truncated [AR69, Coh61a, Coh61b, Cra68, Hol67, LMW64, SV69, SG61, Wei64b, Rif69]. Tukey [Ans65, Jac65, Cox92, De 64, JP61, Olm63, San68]. Two [Add64b, Add69, AC67, Ano64g, Bec67, BG78, Cla65, Coc65, Coh67, Dan63, DS68a, DS68b, Ela61, Eps60c, GH66, GKKK61, Gov63, HB69, HR65, Jae64a, KA67, LB69, Man68, Mar68b, Mar69a, MS68, Nau66, NG68, Nel68, Pat62, Qur64, QNA65, Ril62, Rob60, TB69, Zel62a, BMW65, BMW66]. Two-Factor [Mar68b, Mar69a]. Two-Level [Add64b, Add69, DS68a, HB69]. Two-Parameter [Man68]. Two-Sample [NG68, Zel62a, BMW65, BMW66]. Two-Sided [GH66]. Two-Stage [AC67, Bec67, KA67]. Two-Way [BG78]. Type [Ano65j, Cla67, Sch64, Boh68, HM67, HM68, Har69]. Types [Dun60, Rei63a].

Ulf [Koo64]. Unbalanced [LNJE68]. Unbiased [Gla62, GQS67, McC65, SV69, SD68, Was64]. Underlying [Eps60e, Eps60f]. Undesigned [Ano66h, Dyk66b]. Unequal [Aig68]. Unfolding [NM69]. Unified [Har66]. Uniform [Gun65, Pos63]. United [Str69]. Univariate [Boc63, Har66]. University [Str69]. Unknown [HH69, Owe65b, SG66]. upon [Dun60, HB60]. Upper [Gut62]. Ups [MBM64, SG61]. Use [Box66c, Bra66, Ewa63, Har61a, Kra67, Kru66, TL68, Wei65]. Used [PW65]. Useful [Bas64, BH62, Bur67, ML60, WW61]. User [Ode68]. Uses [Blu66]. Using [BCK60, BCK61, Dun64, Eis68, GQS67, Pag62, Pat66, She65a, BCK78].

V [Gei62, Gei66, Gor66, Ode66a, Sys63]. Vajda [Boo62, Wou64]. Validating [Ell63]. Validity [Eps60e, Eps60f]. Value [Dow66, GM66, HM68, Man68, Pos65, HM67]. Values [Bin68, Dix64, Dra61, Eps60a, LSL68, LMW64, Nad67, Nel66, Smi68b, Wei64b, Zys65]. Variable [BD68b, Mal69, Mar64, OM68]. Variables [Bar62b, BT62, Cla65, CB67, DS66, GT66, Ker67, Mal69, Mil66, Owe67, Owe69, Sum65, BB60, Blo66]. Variance [And65, Boc63, BA63, CR66, Gla62, HH69, JL69, Koc67a, Koc68, KA67, LNJE68, Man64a, NH69, Owe65b, SV69, TM63, Tuk61, ZT68, LN66, Dem60, Low65, Pot67]. Variances [Bec68a, Bec68b, Coc65]. Variate [Kru66]. Variation [Dyk65b, Rob69, Sum65]. Variations [Enr60]. Various [Dun60, Har66, NG68]. Vectors [SS62]. Version [San68]. versus [RS65]. Vertices [Dia67, Gor66, MA66]. Victor [Coh66a]. Viewpoint [Gei66]. ViewpointPart [Gei66]. VII [Cox92]. Vogl [Hes68]. Vol [Bar63c, Coo66, Hay68, Luk64a, Luk64b, McK69d, Par64, Pfa63a, Sam69]. Volume [All63, Ano60j, Ano61l, Ano62l, Ano63n, Ano64m, Ano65p, Ano66o, Ano67m, Ano68k, Ano69k, Gru67, Lea67, Olm62, Ter65, Zys64, Cox92]. Volumes [Ano66i, Ram66]. vs [LE64, Pet64, Tuk60].

W [Blo66, Buc67, Bur68, Cox92, Dan63, De 64, Dyk66a, Dyk67a, Fis67, Gei62, Gei63, Lau68, Man61, McL69, Neu68, Olm63, Roo63, Tho67, Zel62a]. Waiting [Hal64, Neu68]. Waiting-Line [Neu68]. Wallace [Ans65, Jac65]. Walsh [Dav66, Noe63]. Walter [Sch66]. Wani [Lev66]. Warning [Pag63]. Wasow [Ham62]. Watson [Gur65, Pat63a]. Wave [PG65]. Wave-Analyser [PG65]. Way [Ano65i, BG78, WW64]. Wear [SV68]. Weibull [Ano64g, Ano73, Eps64a, BA67, Coh65, Dub65, Dub67b, Dub67a, HM65a, HM65b, Jae64a, JL66, Joh66c, Kao64, Man67, Man68, Men63, Qur64, QNA65, SL63, TBA69, TB69, Whi69]. Weight [DL67, Rob69]. Weighting [Man64a]. Weights [Jae65]. Weiler [Ano67h]. Weiss [Chu66a]. Weissberg [Ano61g]. Wetherill [BLH^{+}60, Cad69, Joh68]. Where [ZT68]. Which [DHT69, Wei62b]. White [Ano80]. Whitney [Sho67b]. Wiener [Rob63]. Wilcoxon [Dun66, Sho67b]. Wild [Kru60]. Wilks [Cap66]. William [Gru67, Low65, Sam69]. Williams [Gil66]. Williamson [Ste65]. Window [Won61]. Wine [Min65b]. Winsorization [DT68]. Winsorized [DT68]. Wisconsin [Str69]. Wolfe [Ano78]. Wolfgang [Ham62]. Work [Owe69]. Workbook [Hic68]. works [Cox92]. World [Hog66]. wrong [Kus69].

Ya [Gei62]. Yamane [Mad68]. Yaspan [Bar68]. Yates [Coo68, Hun66]. Ylvisaker [Cry67]. Youden [Bur68, Fis67]. Young [Dan63]. Yu [Mul65].

Zayachkowski [Zel62a]. Zelditch [Rei63a]. Zelen [Dre63]. Zero [Ku63b]. Zwet [Blo66].

References

Addelman:1964:PCI

[AB64] Sidney Addelman and Seymour Bush. A procedure for constructing incomplete block designs. Technometrics, 6(4):389–403, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

           tronic). URL http://www.jstor.org/stable/1266094.

Anderson:1967:CDE

[AC67] R. L. Anderson and P. P. Crump. Comparisons of designs and estimation procedures for estimating parameters in a two-stage nested process. Technometrics, 9(4):499–516, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

           http://www.jstor.org/stable/1266190.

Addelman:1961:IFF

[Add61] Sidney Addelman. Irregular fractions of the 2^{n} factorial experiments. Technometrics, 3(4):479–496, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266555.

Addelman:1962:OME

[Add62a] Sidney Addelman. Orthogonal main-effect plans for asymmetrical factorial experiments. Technometrics, 4(1):21–46, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266170.

Addelman:1962:SAF

[Add62b] Sidney Addelman. Symmetrical and asymmetrical fractional factorial plans. Technometrics, 4(1):47–58, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266171. See errata [Ano62f].

Addelman:1964:DSA

[Add64a] Sidney Addelman. Designs for the sequential application of factors. Technometrics, 6(4):365–370, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266091.

Addelman:1964:STL

[Add64b] Sidney Addelman. Some two-level factorial plans with split plot confounding. Technometrics, 6(3):253–258, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266042.

Addelman:1965:CRP

[Add65] Sidney Addelman. The construction of a 2^{17−9} resolution V plan in eight blocks of 32. Technometrics, 7(3):439–443, August 1965. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266601.

Addelman:1969:STL

[Add69] Sidney Addelman. Sequences of two-level fractional factorial plans. Technometrics, 11(3):477–509, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267021.

Anscombe:1960:RO

[AG60] F. J. Anscombe and Irwin Guttman. Rejection of outliers. Technometrics, 2(2):123–147, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1266540. See discussion [KFT^{+}60].

Atkinson:1968:DEP

[AH68] Anthony C. Atkinson and William G. Hunter. The design of experiments for parameter estimation. Technometrics, 10(2):271–289, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267043.

Aigner:1968:LAC

[Aig68] D. J. Aigner. A linear approximator for the class marks of a grouped frequency distribution, with especial reference to the unequal interval case. Technometrics, 10(4):793–809, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267460.

Allen:1962:BRB

[All62] W. R. Allen. Book review: Guide to Tables in Mathematical Statistics by J. Arthur Greenwood; H. O. Hartley. Technometrics, 4(3):438–439, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266586.

Alling:1963:BRB

[All63] David W. Alling. Book review: Proceedings of the Fourth Berkeley Symposium on Mathematical Statistics and Probability. Volume IV. Contributions to Biology and Problems of Medicine. Technometrics, 5(2):281, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266076.

Allen:1968:BRB

[All68] David M. Allen. Book review: Handbook of Methods of Applied Statistics by I. M. Chakravarti; R. C. Laha; J. Roy. Technometrics, 10(4):872–873, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267471.

Amster:1965:BRBb

[Ams65a] S. J. Amster. Book review: Mathematical Theory of Reliability by R. E. Barlow; Frank Proschan; Larry C. Hunter. Technometrics, 7(4):656–657, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266408.

Amster:1965:BRBa

[Ams65b] Sigmund J. Amster. Book review: The Statistical Treatment of Fatigue Experiments by Leonard G. Johnson. Technometrics, 7(3): 455, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1266610.

Anderson:1960:SMA

[And60] Edgar Anderson. A semigraphical method for the analysis of complex problems. Technometrics, 2(3):387–391, August 1960. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266252.

Anderson:1965:QNV

[And65] R. L. Anderson. Query 9: Negative variance estimates. Technometrics, 7(1):75–76, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266133.

Andrews:1966:BRB

[And66] Horace P. Andrews. Book review: Biological Statistics: An Introduction by S. C. Pearce. Technometrics, 8(4):716–717, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266651.

Anonymous:1960:BMa

[Ano60a] Anonymous. Back matter. Technometrics, 2(1):122, February 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266538.

Anonymous:1960:BMb

[Ano60b] Anonymous. Back matter. Technometrics, 2(2):269–274, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266550.

Anonymous:1960:BMc

[Ano60c] Anonymous. Back matter. Technometrics, 2(3):407–409, August 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266255.

Anonymous:1960:EFE

[Ano60d] Anonymous. Errata: “Factorial Experiments in Life Testing”. Technometrics, 2(1):121, February 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266537. See [Zel59].

Anonymous:1960:EFG

[Ano60e] Anonymous. Errata to Fisher and Gupta. Technometrics, 2(4): 523–524, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1266462. See [FC60, Gup60].

Anonymous:1960:FMa

[Ano60f] Anonymous. Front matter. Technometrics, 2(1):117–120, February 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266525.

Anonymous:1960:FMb

[Ano60g] Anonymous. Front matter. Technometrics, 2(2):??, May 1960. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266539.

Anonymous:1960:FMc

[Ano60h] Anonymous. Front matter. Technometrics, 2(3):??, August 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266246.

Anonymous:1960:FMd

[Ano60i] Anonymous. Front matter. Technometrics, 2(4):521, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266450.

Anonymous:1960:VI

[Ano60j] Anonymous. Volume information. Technometrics, 2(4):525–526, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266449.

Anonymous:1961:BMa

[Ano61a] Anonymous. Back matter. Technometrics, 3(2):305–310, May 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266121.

Anonymous:1961:BMb

[Ano61b] Anonymous. Back matter. Technometrics, 3(3):447–448, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266741.

Anonymous:1961:BMc

[Ano61c] Anonymous. Back matter. Technometrics, 3(4):582, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266568.

Anonymous:1961:CSP

[Ano61d] Anonymous. Correction: “Statistical Programs for High Speed Computers”. Technometrics, 3(3):446, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266740. See [Leo61b].

Anonymous:1961:EAR

[Ano61e] Anonymous. Errata: “Average Run Lengths in Cumulative Chart Quality Control Schemes”. Technometrics, 3(3):442, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266738. See

[GW61].

Anonymous:1961:EBH

[Ano61f] Anonymous. Errata to Box and Halperin. Technometrics, 3(1):131, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266486.

See [BB60, HB61].

Anonymous:1961:EBW

[Ano61g] Anonymous. Errata to Box and Weissberg. Technometrics, 3 (4):576–577, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266567. See [BH61b, WB60].

Anonymous:1961:FMa

[Ano61h] Anonymous. Front matter. Technometrics, 3(1):127–130, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266471.

Anonymous:1961:FMb

[Ano61i] Anonymous. Front matter. Technometrics, 3(2):??, May 1961. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266109.

Anonymous:1961:FMc

[Ano61j] Anonymous. Front matter. Technometrics, 3(3):??, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266724.

Anonymous:1961:FMd

[Ano61k] Anonymous. Front matter. Technometrics, 3(4):575, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266552.

Anonymous:1961:VI

[Ano61l] Anonymous. Volume information. Technometrics, 3(4):579–581, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266551.

Anonymous:1962:ASG

[Ano62a] Anonymous. Addendum: A study of the group screening method. Technometrics, 4(2):299, May 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266628.

Anonymous:1962:BMa

[Ano62b] Anonymous. Back matter. Technometrics, 4(1):150, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266187.

Anonymous:1962:BMb

[Ano62c] Anonymous. Back matter. Technometrics, 4(2):300, May 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266629.

Anonymous:1962:BMc

[Ano62d] Anonymous. Back matter. Technometrics, 4(4):628, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266303.

Anonymous:1962:EMS

[Ano62e] Anonymous. Errata: “Modified Square Root Method of Matrix Inversion”. Technometrics, 4(4):622, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266301. See [SGR62].

Anonymous:1962:EAG

[Ano62f] Anonymous. Errata to Addelman, Guterman and Sarhan. Technometrics, 4(3):440, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1266587. See [Add62b, Gut62, SGR62].

Anonymous:1962:FMa

[Ano62g] Anonymous. Front matter. Technometrics, 4(1):??, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266168.

Anonymous:1962:FMb

[Ano62h] Anonymous. Front matter. Technometrics, 4(2):293–298, May 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266615.

Anonymous:1962:FMc

[Ano62i] Anonymous. Front matter. Technometrics, 4(3):??, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266569.

Anonymous:1962:FMd

[Ano62j] Anonymous. Front matter. Technometrics, 4(4):621, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266282.

Anonymous:1962:PR

[Ano62k] Anonymous. Publications received. Technometrics, 4(4):623–624, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266302.

Anonymous:1962:VI

[Ano62l] Anonymous. Volume information. Technometrics, 4(4):625–627, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266281.

Anonymous:1963:BMa

[Ano63a] Anonymous. Back matter. Technometrics, 5(1):135–140, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266503.

Anonymous:1963:BMb

[Ano63b] Anonymous. Back matter. Technometrics, 5(2):294–296, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266086.

Anonymous:1963:BMc

[Ano63c] Anonymous. Back matter. Technometrics, 5(3):418–420, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266355.

Anonymous:1963:BMd

[Ano63d] Anonymous. Back matter. Technometrics, 5(4):539–546, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266039.

Anonymous:1963:BRa

[Ano63e] Anonymous. Books received. Technometrics, 5(1):133, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266501.

Anonymous:1963:BRb

[Ano63f] Anonymous. Books received. Technometrics, 5(2):291, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266084.

Anonymous:1963:BRc

[Ano63g] Anonymous. Books received. Technometrics, 5(4):533–534, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266037.

Anonymous:1963:EFN

[Ano63h] Anonymous. Errata: “Finding New Fractions of Factorial Experimental Designs”. Technometrics, 5(1):134, February 1963. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266502. See [Fry61].

Anonymous:1963:EFB

[Ano63i] Anonymous. Errata to Finney, Box and Guterman. Technometrics, 5(3):417, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1266354. See [Fin63, Box63, Gut62].

Anonymous:1963:FMa

[Ano63j] Anonymous. Front matter. Technometrics, 5(1):??, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266487.

Anonymous:1963:FMb

[Ano63k] Anonymous. Front matter. Technometrics, 5(2):??, May 1963. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266058.

Anonymous:1963:FMc

[Ano63l] Anonymous. Front matter. Technometrics, 5(3):??, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266333.

Anonymous:1963:FMd

[Ano63m] Anonymous. Front matter. Technometrics, 5(4):??, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266018.

Anonymous:1963:VI

[Ano63n] Anonymous. Volume information. Technometrics, 5(4):543–545, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266017.

Anonymous:1964:APP

[Ano64a] Anonymous. Abstracts of papers presented at SPES Session of the 123rd Annual Meeting of the American Statistical Association. Technometrics, 6(1):115–125, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266756.

Anonymous:1964:BMa

[Ano64b] Anonymous. Back matter. Technometrics, 6(1):129–132, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266759.

Anonymous:1964:BMb

[Ano64c] Anonymous. Back matter. Technometrics, 6(2):??, May 1964. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266167.

Anonymous:1964:BMc

[Ano64d] Anonymous. Back matter. Technometrics, 6(3):341–342, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266057.

Anonymous:1964:BRa

[Ano64e] Anonymous. Books received. Technometrics, 6(3):339–340, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266056.

Anonymous:1964:BRb

[Ano64f] Anonymous. Books received. Technometrics, 6(4):476, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266106.

Anonymous:1964:EDB

[Ano64g] Anonymous. Errata: “The Discrimination between Two Weibull Processes”. Technometrics, 6(2):??, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266166. See [Qur64].

Anonymous:1964:EHH

[Ano64h] Anonymous. Errata to Halperin and Harter. Technometrics, 6 (4):482–483, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266108. See [Hal64, Har64].

Anonymous:1964:FMa

[Ano64i] Anonymous. Front matter. Technometrics, 6(1):??, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266742.

Anonymous:1964:FMb

[Ano64j] Anonymous. Front matter. Technometrics, 6(2):239–240, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266147.

Anonymous:1964:FMc

[Ano64k] Anonymous. Front matter. Technometrics, 6(3):??, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266040.

Anonymous:1964:FMd

[Ano64l] Anonymous. Front matter. Technometrics, 6(4):479–481, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266088.

Anonymous:1964:VI

[Ano64m] Anonymous. Volume information. Technometrics, 6(4):484–486, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266087.

Anonymous:1965:APP

[Ano65a] Anonymous. Abstracts of papers presented at SPES Session of the 124th Annual Meeting of the American Statistical Association. Technometrics, 7(2):267–277, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266682.

Anonymous:1965:BMa

[Ano65b] Anonymous. Back matter. Technometrics, 7(1):94, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266146.

Anonymous:1965:BMb

[Ano65c] Anonymous. Back matter. Technometrics, 7(2):282, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266684.

Anonymous:1965:BMc

[Ano65d] Anonymous. Back matter. Technometrics, 7(4):665–672, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266415.

Anonymous:1965:BRa

[Ano65e] Anonymous. Books received. Technometrics, 7(1):89–90, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266144.

Anonymous:1965:BRb

[Ano65f] Anonymous. Books received. Technometrics, 7(2):266, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266681.

Anonymous:1965:BRc

[Ano65g] Anonymous. Books received. Technometrics, 7(3):457–458, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266612.

Anonymous:1965:BRd

[Ano65h] Anonymous. Books received. Technometrics, 7(4):663–664, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266414.

Anonymous:1965:END

[Ano65i] Anonymous. Errata: “A Note on the Determination of Connectedness in an N-Way Cross Classification”. Technometrics, 7(2):281, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266683. See

[WW64].

Anonymous:1965:EEC

[Ano65j] Anonymous. Errata: “Estimation of a Component of a Convolution, When the Other Component Is of Exponential Type”. Technometrics, 7(3):462, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266614. See [Sch64].

Anonymous:1965:ESF

[Ano65k] Anonymous. Errata: “Sequential Factorial Estimation”. Technometrics, 7(1):93, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266145. See [Hun64].

Anonymous:1965:FMa

[Ano65l] Anonymous. Front matter. Technometrics, 7(1):91–92, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266122.

Anonymous:1965:FMb

[Ano65m] Anonymous. Front matter. Technometrics, 7(2):279–280, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266665.

Anonymous:1965:FMc

[Ano65n] Anonymous. Front matter. Technometrics, 7(3):459–460, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266588.

Anonymous:1965:FMd

[Ano65o] Anonymous. Front matter. Technometrics, 7(4):??, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266387.

Anonymous:1965:VI

[Ano65p] Anonymous. Volume information. Technometrics, 7(4):667–670, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266386.

Anonymous:1966:ASS

[Ano66a] Anonymous. Abstracts of SPES Sessions of 125th Annual Meeting of ASA. Technometrics, 8(1):203–211, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266278.

Anonymous:1966:BMa

[Ano66b] Anonymous. Back matter. Technometrics, 8(1):216, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266280.

Anonymous:1966:BMb

[Ano66c] Anonymous. Back matter. Technometrics, 8(2):388, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266385.

Anonymous:1966:BMc

[Ano66d] Anonymous. Back matter. Technometrics, 8(3):??, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266712.

Anonymous:1966:BMd

[Ano66e] Anonymous. Back matter. Technometrics, 8(4):737–738, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266664.

Anonymous:1966:E

[Ano66f] Anonymous. Editorial. Technometrics, 8(1):1–2, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266257.

Anonymous:1966:EMD

[Ano66g] Anonymous. Errata: “Misclassified Data from a Binomial Population”. Technometrics, 8(1):215, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266279. See [Coh60].

Anonymous:1966:EOU

[Ano66h] Anonymous. Errata: “The Orthogonalization of Undesigned Experiments”. Technometrics, 8(4):731, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266663. See [Dyk66b].

Anonymous:1966:EIV

[Ano66i] Anonymous. Errata to index of volumes 1–7. Technometrics, 8(2): 387, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266384.

Anonymous:1966:EOC

[Ano66j] Anonymous. Errata to Owen and Cohen. Technometrics, 8(3):570, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266711.

See [Owe65c, Coh66b].

Anonymous:1966:FMa

[Ano66k] Anonymous. Front matter. Technometrics, 8(1):213, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266256.

Anonymous:1966:FMb

[Ano66l] Anonymous. Front matter. Technometrics, 8(2):385–386, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266356.

Anonymous:1966:FMc

[Ano66m] Anonymous. Front matter. Technometrics, 8(3):569, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266685.

Anonymous:1966:FMd

[Ano66n] Anonymous. Front matter. Technometrics, 8(4):??, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266631.

Anonymous:1966:VI

[Ano66o] Anonymous. Volume information. Technometrics, 8(4):733–736, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266630.

Anonymous:1967:ASS

[Ano67a] Anonymous. Abstracts of SPES Sessions of 126th Annual Meeting of ASA. Technometrics, 9(1):185–194, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266330.

Anonymous:1967:BMa

[Ano67b] Anonymous. Back matter. Technometrics, 9(1):196, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266332.

Anonymous:1967:BMb

[Ano67c] Anonymous. Back matter. Technometrics, 9(2):354, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266440.

Anonymous:1967:BMc

[Ano67d] Anonymous. Back matter. Technometrics, 9(3):??, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266524.

Anonymous:1967:BMd

[Ano67e] Anonymous. Back matter. Technometrics, 9(4):707–710, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266217.

Anonymous:1967:EQC

[Ano67f] Anonymous. Errata: “Query 21: Comparison of Sample Sizes in Inverse Binomial Sampling”. Technometrics, 9(4):695, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266208. See

[Bar67a].

Anonymous:1967:EBM

[Ano67g] Anonymous. Errata to Bartko and Mohan. Technometrics, 9(3): 498, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1266523. See [BGCR67, Moh67].

Anonymous:1967:EHW

[Ano67h] Anonymous. Errata to Harter and Weiler. Technometrics, 9(1):195, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266331.

See [Har67b, Wei66].

Anonymous:1967:FMa

[Ano67i] Anonymous. Front matter. Technometrics, 9(1):??, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266313.

Anonymous:1967:FMb

[Ano67j] Anonymous. Front matter. Technometrics, 9(2):351–352, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266416.

Anonymous:1967:FMc

[Ano67k] Anonymous. Front matter. Technometrics, 9(3):??, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266504.

Anonymous:1967:FMd

[Ano67l] Anonymous. Front matter. Technometrics, 9(4):??, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266189.

Anonymous:1967:VI

[Ano67m] Anonymous. Volume information. Technometrics, 9(4):711–714, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266188.

Anonymous:1968:APP

[Ano68a] Anonymous. Abstracts of papers presented at SPES Session of the 127th Annual Meeting of the American Statistical Association. Technometrics, 10(2):419–427, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267068.

Anonymous:1968:BMa

[Ano68b] Anonymous. Back matter. Technometrics, 10(1):217–218, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266245.

Anonymous:1968:BMb

[Ano68c] Anonymous. Back matter. Technometrics, 10(2):435–438, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267070.

Anonymous:1968:BMc

[Ano68d] Anonymous. Back matter. Technometrics, 10(3):636, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267128.

Anonymous:1968:BMd

[Ano68e] Anonymous. Back matter. Technometrics, 10(4):885–892, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267483.

Anonymous:1968:EML

[Ano68f] Anonymous. Erratum: “Maximum-Likelihood Estimation of the Parameters of a Four-Parameter Generalized Gamma Population from Complete and Censored Samples”. Technometrics, 10(3):635, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267127.

See [Har67b].

Anonymous:1968:FMa

[Ano68g] Anonymous. Front matter. Technometrics, 10(1):??, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266218.

Anonymous:1968:FMb

[Ano68h] Anonymous. Front matter. Technometrics, 10(2):??, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267039.

Anonymous:1968:FMc

[Ano68i] Anonymous. Front matter. Technometrics, 10(3):??, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267099.

Anonymous:1968:FMd

[Ano68j] Anonymous. Front matter. Technometrics, 10(4):??, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267449.

Anonymous:1968:VI

[Ano68k] Anonymous. Volume information. Technometrics, 10(4):887–891, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267448.

Anonymous:1969:APP

[Ano69a] Anonymous. Abstracts of papers presented at SPES session of the 128th annual meeting of the American statistical association. Technometrics, 11(1):211–221, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266782.

Anonymous:1969:BMa

[Ano69b] Anonymous. Back matter. Technometrics, 11(1):223–239, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266783.

Anonymous:1969:BMb

[Ano69c] Anonymous. Back matter. Technometrics, 11(2):??, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267279.

Anonymous:1969:BMc

[Ano69d] Anonymous. Back matter. Technometrics, 11(3):628, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267038.

Anonymous:1969:BMd

[Ano69e] Anonymous. Back matter. Technometrics, 11(4):845–847, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266911.

Anonymous:1969:EDD

[Ano69f] Anonymous. Errata to Draper and DeGray. Technometrics, 11(3): 627, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1267037. See [DHT69, DeG68b].

Anonymous:1969:FMa

[Ano69g] Anonymous. Front matter. Technometrics, 11(1):207–210, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266760.

Anonymous:1969:FMb

[Ano69h] Anonymous. Front matter. Technometrics, 11(2):401–402, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267256.

Anonymous:1969:FMc

[Ano69i] Anonymous. Front matter. Technometrics, 11(3):??, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267016.

Anonymous:1969:FMd

[Ano69j] Anonymous. Front matter. Technometrics, 11(4):835, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266884.

Anonymous:1969:VI

[Ano69k] Anonymous. Volume information. Technometrics, 11(4):848–851, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266883.

Anonymous:1973:EML

[Ano73] Anonymous. Erratum: “Maximum-Likelihood Estimation of the Parameters of Gamma and Weibull Populations from Complete and from Censored Samples”. Technometrics, 15(2):431, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267014. See

[Coh65].

Anonymous:1978:EWB

[Ano78] Anonymous. Errata to Wolfe, Bose and Bradu. Technometrics, 20(2):219, May 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268731. See [Wol77, BCK61, BG78].

Anonymous:1980:CLW

[Ano80] Anonymous. Corrigenda to Leone, White, Narula and Narula. Technometrics, 22(3):452, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268342.

Anscombe:1965:CKL

[Ans65] F. J. Anscombe. Comments on Kurtz–Link–Tukey–Wallace paper. Technometrics, 7(2):167–168, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266668. See [KLTW65b, KLTW65a].

Aroian:1966:SLT

[AR66] L. A. Aroian and D. E. Robison. Sequential life tests for the exponential distribution with changing parameter. Technometrics, 8(2):217–227, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266357.

Aroian:1969:DME

[AR69] L. A. Aroian and D. E. Robison. Direct methods for exact truncated sequential tests of the mean of a normal distribution. Technometrics, 11(4):661–675, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266890.

Aroian:1968:SAD

[Aro68] Leo A. Aroian. Sequential analysis, direct method. Technometrics, 10(1):125–132, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266229.

Anscombe:1963:EAR

[AT63] F. J. Anscombe and John W. Tukey. The examination and analysis of residuals. Technometrics, 5(2):141–160, May 1963. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266059. See corrections

[Olm63, De 64].

Atkinson:1969:CMD

[Atk69] Anthony C. Atkinson. Constrained maximisation and the design of experiments. Technometrics, 11(3):616–618, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267030.

Bush:1963:CTD

[BA63] Norman Bush and R. L. Anderson. A comparison of three different procedures for estimating variance components. Technometrics, 5(4):421–440, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266019.

Bain:1967:EPW

[BA67] Lee J. Bain and Charles E. Antle. Estimation of parameters in the Weibull distribution. Technometrics, 9(4):621–627, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266200.

Bainbridge:1963:LE

[Bai63] J. R. Bainbridge. Letter to the Editor. Technometrics, 5(2):292– 293, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266085.

Bain:1967:BRB

[Bai67] Lee J. Bain. Book review: Applied Regression Analysis by N. H. Draper; H. Smith. Technometrics, 9(1):182–183, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266329.

Baker:1964:BRB

[Bak64] G. A. Baker. Book review: Applications and Problems in Elementary Statistics by Leonard S. Cahen; Duncan N. Hansen. Technometrics, 6(2):231, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266161.

Bancroft:1969:BRB

[Ban69] T. A. Bancroft. Book review: Research Papers in Statistics by F. N. David. Technometrics, 11(4):838–839, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266906.

Barnard:1962:ATI

[Bar62a] G. A. Barnard. An analogue of Tchebycheff’s inequality in terms of range. Technometrics, 4(1):133–134, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266178.

Barnard:1962:DPI

[Bar62b] G. A. Barnard. Distribution of products of independent variables. Technometrics, 4(2):277–278, May 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266624.

Bartko:1962:NNB

[Bar62c] John J. Bartko. A note on the negative binomial distribution. Technometrics, 4(4):609–610, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266292.

Barlow:1963:BRB

[Bar63a] Richard E. Barlow. Book review: Renewal Theory by D. R. Cox. Technometrics, 5(4):529–530, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266033.

Bartholomew:1963:SDE

[Bar63b] D. J. Bartholomew. The sampling distribution of an estimate arising in life testing. Technometrics, 5(3):361–374, August 1963. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266339.

Bartlett:1963:BRB

[Bar63c] M. S. Bartlett. Book review: Proceedings of the Fourth Berkeley Symposium on Mathematical Statistics and Probability. Vol. III: Contributions to Astronomy, Meteorology and Physics by J. Neyman. Technometrics, 5(2):279–280, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266075.

Bartko:1966:ANB

[Bar66] John J. Bartko. Approximating the negative binomial. Technometrics, 8(2):345–350, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266367.

Barton:1967:QCS

[Bar67a] D. E. Barton. Query 21: Comparison of sample sizes in inverse binomial sampling. Technometrics, 9(2):337–339, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266432. See er-

rata [Ano67f].

Barton:1967:QCR

[Bar67b] D. E. Barton. Query 25: Completed runs of length k above and below median. Technometrics, 9(4):682–694, November 1967. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266207.

Bartoo:1968:BRB

[Bar68] James B. Bartoo. Book review: Essentials of Probability by A. Yaspan. Technometrics, 10(4):878, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267478.

Basu:1964:ERS

[Bas64] A. P. Basu. Estimates of reliability for some distributions useful in life testing. Technometrics, 6(2):215–219, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266153.

Basmann:1968:BRB

[Bas68] R. L. Basmann. Book review: Stochastically Dependent Equations: An Introductory Text for Econometricians by P. R. Fisk.

Technometrics, 10(4):874, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267473.

Box:1960:SNT

[BB60] G. E. P. Box and D. W. Behnken. Some new three level designs for the study of quantitative variables. Technometrics, 2(4):455–475, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266454.

See errata [Ano61f].

Booth:1962:SSS

[BC62] Kathleen H. V. Booth and D. R. Cox. Some systematic supersaturated designs. Technometrics, 4(4):489–495, November 1962. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266285.

Barnard:1963:LE

[BC63] G. A. Barnard and D. R. Cox. Letter to the Editor. Technometrics, 5(3):416, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266353.

Beauchamp:1966:SNE

[BC66] John J. Beauchamp and Richard G. Cornell. Simultaneous nonlinear estimation. Technometrics, 8(2):319–326, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266364. See errata [BC67].

Beauchamp:1967:ESN

[BC67] John J. Beauchamp and Richard G. Cornell. Errata to: “Simultaneous nonlinear estimation”. Technometrics, 9(2):353, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266439. See

[BC66].

Beauchamp:1969:SSE

[BC69] John J. Beauchamp and Richard G. Cornell. Spearman simultaneous estimation for a compartmental model. Technometrics, 11(3):551–560, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267024.

Bose:1960:MCS

[BCK60] R. C. Bose, I. M. Chakravarti, and D. E. Knuth. On methods of constructing sets of mutually orthogonal Latin squares using a computer. I. Technometrics, 2(4):507–516, November 1960. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266458.

Bose:1961:MCS

[BCK61] R. C. Bose, I. M. Chakravarti, and D. E. Knuth. On methods of constructing sets of mutually orthogonal Latin squares using a computer. II. Technometrics, 3(1):111–117, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266482. See er-

rata [Ano78, BCK78].

Bose:1978:EMC

[BCK78] R. C. Bose, I. M. Chakravarti, and D. E. Knuth. Errata: “On methods of constructing sets of mutually orthogonal Latin squares using a computer. II” (Technometrics 3 (1961), 111–117). Technometrics, 20(2):219, 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). See [BCK61].

Bessler:1962:OSA

[BCM62] S. Bessler, H. Chernoff, and A. W. Marshall. An optimal sequential accelerated life test. Technometrics, 4(3):367–379, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266574.

Bhapkar:1968:SNT

[BD68a] V. P. Bhapkar and Jayant V. Deshpande. Some nonparametric tests for multisample problems. Technometrics, 10(3):578–585, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267111.

Box:1968:IMP

[BD68b] Geroge E. P. Box and Norman R. Draper. Isn’t my process too variable for EVOP? Technometrics, 10(3):439–444, August 1968. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267100.

Bechhofer:1967:TSS

[Bec67] Robert E. Bechhofer. A two-stage subsampling procedure for ranking means of finite populations with an application to bulk sampling problems. Technometrics, 9(3):355–364, August 1967. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266505.

Bechhofer:1968:MCC

[Bec68a] Robert E. Bechhofer. Multiple comparisons with a control for multiply-classified variances of normal populations. Technometrics, 10(4):715–718, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267454.

Bechhofer:1968:SSP

[Bec68b] Robert E. Bechhofer. Single-stage procedures for ranking multiplyclassified variances of normal populations. Technometrics, 10(4): 693–714, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267453.

Bennett:1964:NCC

[Ben64] B. M. Bennett. A note on combining correlated estimates of a ratio of multivariate means. Technometrics, 6(4):463–467, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266100.

Beranek:1961:BRB

[Ber61] William Beranek. Book review: Dynamic Programming and Markov Processes by Ronald A. Howard. Technometrics, 3(1): 120–121, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266484.

Bergstrom:1966:BRB

[Ber66] Harald Bergström. Book review: Mathematical Theory of Probability and Statistics by Richard von Mises; Hilda Geiringer. Technometrics, 8(3):547–548, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266701.

Berman:1967:BRB

[Ber67] Simeon M. Berman. Book review: Stationary and Related Stochastic Processes by H. Cramér; M. R. Leadbetter. Technometrics, 9(4):701–702, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266213.

Berkson:1969:ELF

[Ber69] Joseph Berkson. Estimation of a linear function for a calibration line; consideration of a recent proposal. Technometrics, 11 (4):649–660, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266889.

Birnbaum:1961:MCS

[BES61] Z. W. Birnbaum, J. D. Esary, and S. C. Saunders. Multicomponent systems and structures and their reliability. Technometrics, 3(1):55–77, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266477.

Balaam:1965:QER

[BF65] L. N. Balaam and W. T. Federer. Query 11: Error rate bases. Technometrics, 7(2):260–262, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266676.

Barlow:1966:DFL

[BG66] Richard E. Barlow and Shanti S. Gupta. Distribution-free life test sampling plans. Technometrics, 8(4):591–613, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266633.

Bradu:1978:BDT

[BG78] Dan Bradu and K. Ruben Gabriel. The biplot as a diagnostic tool for models of two-way tables. Technometrics, 20(1):47–68, February 1978. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268161.

See errata [Ano78].

Bartko:1967:LE

[BGCR67] John J. Bartko, B. K. Ghosh, C. Philip Cox, and Thomas D. Roseberry. Letters to the Editor. Technometrics, 9(2):347–350, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266438. See

[Ano67g].

Barlow:1960:SER

[BH60] R. E. Barlow and L. C. Hunter. System efficiency and reliability. Technometrics, 2(1):43–53, February 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266530.

Box:1961:FFDa

[BH61a] G. E. P. Box and J. S. Hunter. The 2^{k−p} fractional factorial designs. Part I. Technometrics, 3(3):311–351, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266725.

Box:1961:FFDb

[BH61b] G. E. P. Box and J. S. Hunter. The 2^{k−p} fractional factorial designs Part II. Technometrics, 3(4):449–458, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266553. See errata [Ano61g].

Box:1962:UMM

[BH62] G. E. P. Box and William G. Hunter. A useful method for modelbuilding. Technometrics, 4(3):301–318, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266570.

Box:1965:ESP

[BH65] G. E. P. Box and William G. Hunter. The experimental study of physical mechanisms. Technometrics, 7(1):23–42, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266125.

Box:1967:DAM

[BH67] G. E. P. Box and W. J. Hill. Discrimination among mechanistic models. Technometrics, 9(1):57–71, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266318.

Bingham:1968:AMS

[Bin68] Richard S. Bingham, Jr. Approximations for mean square successive difference critical values. Technometrics, 10(2):397–400, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267056.

Bhapker:1968:HNI

[BK68] V. P. Bhapker and Gary G. Koch. Hypotheses of ‘no interaction’ in multi-dimensional contingency tables. Technometrics, 10 (1):107–123, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266228.

Blankenship:1966:BRB

[Bla66] John H. Blankenship. Book review: Nonparametric Trend Analysis by George A. Ferguson. Technometrics, 8(4):724–725, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266659.

Barnard:1960:DPM

[BLH^{+}60] G. A. Barnard, D. V. Lindley, B. Hill, I. J. Good, G. Horsnell, D. R. Cox, F. J. Anscombe, and A. Hald. Discussion of the papers of Messrs. Hald, Wetherill and Cox. Technometrics, 2(3):361–372, August 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266250.

See [Hal60, Wet60, Cox60].

Blom:1966:BRB

[Blo66] Gunnar Blom. Book review: Convex Transformations of Random Variables by W. R. van Zwet. Technometrics, 8(2):380–381, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266380.

Blumenthal:1966:BRB

[Blu66] Saul Blumenthal. Book review: Probability and Its Engineering Uses by Thornton C. Fry. Technometrics, 8(2):372–375, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266376.

Blumenthal:1967:PSL

[Blu67] Saul Blumenthal. Proportional sampling in life length studies. Technometrics, 9(2):205–218, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266418.

Brinegar:1960:SEG

[BM60] Claude S. Brinegar and Ronald R. Miller. Statistical estimation of the gasoline octane number requirement of new model automobiles. Technometrics, 2(1):5–18, February 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266527.

Barlow:1968:SEP

[BMPS68] Richard E. Barlow, Albert Madansky, Frank Proschan, and Ernest M. Scheuer. Statistical estimation procedures for the “burnin” process. Technometrics, 10(1):51–62, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266223.

Bradley:1965:SRT

[BMW65] Ralph A. Bradley, Donald C. Martin, and Frank Wilcoxon. Sequential rank tests. I. Monte Carlo studies of the two-sample procedure. Technometrics, 7(4):463–483, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266388.

Bradley:1966:SRT

[BMW66] Ralph A. Bradley, Sarla D. Merchant, and Frank Wilcoxon. Sequential rank tests. II. Modified two-sample procedures. Technometrics, 8(4):615–623, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266634.

Bennett:1963:TTS

[BN63] B. M. Bennett and E. Nakamura. Tables for testing significance in a 2 × 3 contingency table. Technometrics, 5(4):501–511, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266026.

Bennett:1964:PFE

[BN64] B. M. Bennett and E. Nakamura. The power function of the exact test for the 2 × 3 contingency table. Technometrics, 6(4):439–458, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266098.

Bock:1963:PUM

[Boc63] R. Darrell Bock. Programming univariate and multivariate analysis of variance. Technometrics, 5(1):95–117, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266494.

Bohrer:1968:NTL

[Boh68] Robert Bohrer. A note on tolerance limits with Type I censoring. Technometrics, 10(2):392, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267054.

Boot:1962:BRB

[Boo62] J. C. G. Boot. Book review: Mathematical Programming by S. Vajda. Technometrics, 4(4):618–619, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266299.

Borsting:1966:BRB

[Bor66] J. R. Borsting. Book review: Probability and Statistics for Engineers by I. Miller; J. E. Freund. Technometrics, 8(2):383–384, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266383.

Bowden:1968:QTI

[Bow68] David C. Bowden. Query 26: Tolerance interval in regression. Technometrics, 10(1):207–209, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266238.

Box:1963:EEF

[Box63] G. E. P. Box. The effects of errors in the factor levels and experimental design. Technometrics, 5(2):247–262, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266066. See errata [Ano63i].

Box:1966:NAD

[Box66a] George E. P. Box. A note on augmented designs. Technometrics, 8(1):184–188, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266270.

Box:1966:SSE

[Box66b] George E. P. Box. A simple system of evolutionary operation subject to empirical feedback. Technometrics, 8(1):19–26, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266259.

Box:1966:UAR

[Box66c] George E. P. Box. Use and abuse of regression. Technometrics, 8(4):625–629, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266635.

Bondar:1968:BRB

[BP68] James Bondar and Joseph Putter. Book review: Simultaneous Statistical Inference by R. G. Miller. Technometrics, 10(2):415–416, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267066.

Bhattacharjee:1963:DCI

[BPM63] G. P. Bhattacharjee, S. N. N. Pandit, and R. Mohan. Dimensional chains involving rectangular and normal error-distributions. Technometrics, 5(3):404–406, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266346.

Bracken:1966:PPB

[Bra66] Jerome Bracken. Percentage points of the beta distribution for use in Bayesian analysis of Bernoulli processes. Technometrics, 8(4):687–694, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266640.

Bradley:1968:MCA

[Bra68] Hugh E. Bradley. Multiple classification analysis for arbitrary experimental arrangements. Technometrics, 10(1):13–27, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266220.

Bross:1961:OPE

[Bro61] Irwin D. J. Bross. Outliers in patterned experiments: A strategic appraisal. Technometrics, 3(1):91–102, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266479.

Brownlee:1968:BRB

[Bro68] K. A. Brownlee. Book review: Statistical Models in Engineering by G. J. Hahn; S. S. Shapiro. Technometrics, 10(1):214–215, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266243.

Barlow:1966:RGD

[BS66] Richard E. Barlow and Ernest M. Scheuer. Reliability growth during a development testing program. Technometrics, 8(1):53–60, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266261.

Box:1962:TIV

[BT62] G. E. P. Box and Paul W. Tidwell. Transformation of the independent variables. Technometrics, 4(4):531–550, November 1962. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266288.

Buckland:1967:BRB

[Buc67] Wm. R. Buckland. Book review: Reliability Handbook by W. G. Ireson. Technometrics, 9(2):344–345, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266437.

Burr:1965:QCS

[Bur65] Irving W. Burr. Query 8: Calculation of the sampling distribution of the range. Technometrics, 7(1):73–75, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266132.

Burr:1967:UAN

[Bur67] Irving W. Burr. A useful approximation to the normal distribution function, with application to simulation. Technometrics, 9 (4):647–651, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266202.

Burr:1968:BRB

[Bur68] Irving W. Burr. Book review: Statistical Techniques for Collaborative Tests by W. J. Youden. Technometrics, 10(1):214, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266242.

Burr:1969:BRB

[Bur69] Irving W. Burr. Book review: Cumulative Sum Tests: Theory and Practice by C. S. van Dobben de Bruyn. Technometrics, 11 (3):625–626, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267036.

Cady:1969:BRB

[Cad69] Foster B. Cady. Book review: Elementary Statistical Methods by G. B. Wetherill. Technometrics, 11(4):842–843, November 1969. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266910.

Cameron:1969:SCS

[Cam69] J. M. Cameron. The statistical consultant in a scientific laboratory. Technometrics, 11(2):247–254, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267258.

Capobianco:1966:BRB

[Cap66] Michael F. Capobianco. Book review: Introductory Engineering Statistics by Irwin Guttman; S. S. Wilks. Technometrics, 8 (3):554–555, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266706.

Chew:1962:DRE

[CB62] Victor Chew and Ray Boyce. Distribution of radial error in the bivariate elliptical normal distribution. Technometrics, 4(1):138–140, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266181.

Clutton-Brock:1967:LDE

[CB67] M. Clutton-Brock. Likelihood distributions for estimating functions when both variables are subject to error. Technometrics, 9(2):261–269, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266422.

Cacoullos:1967:DBR

[CD67] T. Cacoullos and H. DeCicco. On the distribution of the bivariate range. Technometrics, 9(3):476–480, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266516.

Csorgo:1962:ECT

[CG62] M. Csörgő and Irwin Guttman. On the empty cell test. Technometrics, 4(2):235–247, May 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266621.

Chanmugan:1962:BRBa

[Cha62a] J. Chanmugan. Book review: Industrial Productivity and Motivation by H. C. Ganguli. Technometrics, 4(3):437, August 1962. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266583.

Chanmugan:1962:BRBb

[Cha62b] J. Chanmugan. Book review: Proceedings of the Second International Conference on Operational Research by J. Banbury; J. Maitland. Technometrics, 4(3):437, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266584.

Chakravarti:1966:QST

[Cha66] I. M. Chakravarti. Query 15: Small trimmed samples. Technometrics, 8(1):193, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266272.

Chalmers:1967:LE

[Cha67] C. P. Chalmers. Letters to the Editor. Technometrics, 9(4):705, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266216.

Cox:1967:QAF

[CHDF67] D. R. Cox, Agnes Herzberg, Cuthbert Daniel, and D. J. Finney. Query 20: Analysis of a factorial experiment (partially confounded 2^{3} ). Technometrics, 9(1):167–179, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266327.

Chernoff:1962:OAL

[Che62] Herman Chernoff. Optimal accelerated life designs for estimation. Technometrics, 4(3):381–408, August 1962. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266575.

Chernoff:1967:QDF

[Che67] Herman Chernoff. Query 23: Degrees of freedom for chi-square. Technometrics, 9(3):489–490, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266518.

Chew:1968:SPI

[Che68] Victor Chew. Simultaneous prediction intervals. Technometrics, 10(2):323–330, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267047.

Chow:1962:NCC

[Cho62] Wen M. Chow. A note on the calculation of certain constrained maxima. Technometrics, 4(1):135–137, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266180.

Church:1966:BRB

[Chu66a] John O. Church. Book review: Programming the IBM 1620 by Eric A. Weiss. Technometrics, 8(3):557–558, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266709.

Church:1966:ADW

[Chu66b] Alonzo Church, Jr. Analysis of data when the response is a curve. Technometrics, 8(2):229–246, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266358.

Chun:1968:BRB

[Chu68a] D. Chun. Book review: Mathematics and Statistics for Chemists by C. J. Brooks; I. G. Betteley; S. M. Loxston. Technometrics, 10(1):212–213, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266240.

Chun:1968:NRT

[Chu68b] Dan Chun. A note on a regression transformation for smaller round-off error. Technometrics, 10(2):393–396, May 1968. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267055.

Cramer:1964:MBA

[CJ64] Harald Cramér and N. L. Johnson. Model building with the aid of stochastic processes. Technometrics, 6(2):133–159, May 1964. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266148.

Cox:1967:NGA

[CL67] D. R. Cox and Elizabeth Lauh. A note on the graphical analysis of multidimensional contingency tables. Technometrics, 9 (3):481–488, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266517.

Clark:1965:CLP

[Cla65] Virginia Clark. Choice of levels in polynomial regression with one or two variables. Technometrics, 7(3):325–333, August 1965. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266592.

Clatworthy:1967:SNF

[Cla67] Willard H. Clatworthy. Some new families of partially balanced designs of the Latin square type and related designs. Technometrics, 9(2):229–244, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266420.

Cochran:1964:PLD

[Coc64] William G. Cochran. On the performance of the linear discriminant function. Technometrics, 6(2):179–190, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266150.

Cochran:1965:QTT

[Coc65] William G. Cochran. Query 12: Testing two correlated variances. Technometrics, 7(3):447–449, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266603.

Cochran:1968:CEE

[Coc68a] W. G. Cochran. Commentary on “Estimation of Error Rates in Discriminant Analysis”. Technometrics, 10(1):204–205, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266237. See

[LM68].

Cochran:1968:EMS

[Coc68b] W. G. Cochran. Errors of measurement in statistics. Technometrics, 10(4):637–666, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267450.

Cohen:1960:MDB

[Coh60] A. Clifford Cohen, Jr. Misclassified data from a binomial population. Technometrics, 2(1):109–113, February 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266535. See errata [Ano66g].

Cohen:1961:EPP

[Coh61a] A. Clifford Cohen, Jr. Estimating the Poisson parameter from samples that are truncated on the right. Technometrics, 3(3):433–438, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266735.

Cohen:1961:TML

[Coh61b] A. Clifford Cohen, Jr. Tables for maximum likelihood estimates: Singly truncated and singly censored samples. Technometrics, 3 (4):535–541, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266559.

Cohen:1963:PCS

[Coh63] A. Clifford Cohen, Jr. Progressively censored samples in life testing. Technometrics, 5(3):327–339, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266337.

Cohen:1965:MLE

[Coh65] A. Clifford Cohen. Maximum likelihood estimation in the Weibull distribution based on complete and on censored samples. Technometrics, 7(4):579–588, November 1965. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266397. See errata [Ano73].

Cohen:1966:DEP

[Coh66a] A. Clifford Cohen. Discussion of “Estimation of Parameters for a Mixture of Normal Distributions” by Victor Hasselblad. Technometrics, 8(3):445–446, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266690. See [Has66].

Cohen:1966:QLT

[Coh66b] A. Clifford Cohen. Query 18: Life testing and early failure. Technometrics, 8(3):539–545, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266699. See errata [Ano66j].

Cohen:1966:BRBb

[Coh66c] Arthur Cohen. Book review: An Introduction to Mathematical Statistics by H. D. Brunk. Technometrics, 8(4):717–718, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266652.

Cohen:1966:BRBa

[Coh66d] A. Clifford Cohen, Jr. Book review: Statistics and Experimental Design in Engineering and the Physical Sciences by Norman Johnson; Fred Leone. Technometrics, 8(2):382–383, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266382.

Cohen:1967:EMT

[Coh67] A. Clifford Cohen. Estimation in mixtures of two normal distributions. Technometrics, 9(1):15–28, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266315.

Conkling:1962:DCE

[Con62] Randall M. Conkling. Decimal correction error: An example in statistics. Technometrics, 4(3):421–425, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266578.

Connor:1963:BRB

[Con63] W. S. Connor. Book review: Introduction to Scientific Inference by Robert Hooke. Technometrics, 5(3):410–411, August 1963. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266348.

Cooper:1966:BRB

[Coo66] Leon Cooper. Book review: Error in Digital Computation, Vol. II by L. B. Rall. Technometrics, 8(4):715–716, November 1966. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266650.

Cooper:1968:EYA

[Coo68] B. E. Cooper. The extension of Yates’ 2^{n} algorithm to any complete factorial experiment. Technometrics, 10(3):575–577, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267110.

Cornell:1969:BRB

[Cor69] John A. Cornell. Book review: The Elements of Probability by Simeon M. Berman. Technometrics, 11(3):619–620, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267031.

Cote:1969:BRB

[Cot69] Louis J. Cote. Book review: Multichannel Time Series Analysis with Digital Computer Programs by Enders A. Robinson. Technometrics, 11(2):408–409, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267278.

Cowden:1967:BRB

[Cow67] Dudley J. Cowden. Book review: Basic Statistics with Business Applications by Richard C. Clelland; John S. de Cani; Francis E. Brown; J. Parker Bursk; Donald S. Murray. Technometrics, 9(4):700–701, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266212.

Cox:1960:SSA

[Cox60] D. R. Cox. Serial sampling acceptance schemes derived from Bayes’s theorem. Technometrics, 2(3):353–360, August 1960. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266249. See discussion

[BLH^{+}60].

Cox:1965:RMC

[Cox65] D. R. Cox. A remark on multiple comparison methods. Technometrics, 7(2):223–224, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266671.

Cox:1992:CWJ

[Cox92] D. R. Cox, editor. The collected works of John W. Tukey. Volume VII. Factorial & ANOVA (1949–1962). Wadsworth & Brooks/Cole Statistics/Probability Series. Wadsworth & Brooks/Cole Advanced Books & Software, Monterey, CA, USA, 1992. ISBN 0-534-05104-9. lxvi + 260 + I6 pp. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Cox:1966:NVD

[CR66] C. Philip Cox and Thomas D. Roseberry. A note on the variance of the distribution of sample number in sequential probability ratio tests. Technometrics, 8(4):700–704, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266642.

Craig:1962:CSS

[Cra62] C. C. Craig. On a class of simple sequential tests on means. Technometrics, 4(3):345–359, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266572.

Craswell:1965:BRB

[Cra65] K. J. Craswell. Book review: Introduction to Probability and Statistics by Henry L. Alder; Edward B. Roessler. Technometrics, 7(2): 264, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266679.

Craig:1968:ASN

[Cra68] C. C. Craig. The average sample number for truncated single and double attributes acceptance sampling plans. Technometrics, 10 (4):685–692, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267452.

Crow:1968:BRB

[Cro68] Edwin L. Crow. Book review: Measurement and Analysis of Random Data by J. S. Bendat; A. G. Piersol. Technometrics, 10 (4):869–871, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267469.

Cryer:1967:BRB

[Cry67] Jon Cryer. Book reviews: Discrete Probability by R. A. Gangolli; Donald Ylvisaker; Introduction to Finite Probability by Roger L. Burford. Technometrics, 9(4):698–700, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266211.

Choi:1969:MLE

[CW69] S. C. Choi and R. Wette. Maximum likelihood estimation of the parameters of the gamma distribution and their bias. Technometrics, 11(4):683–690, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266892.

Daniel:1960:LOF

[Dan60a] Cuthbert Daniel. Locating outliers in factorial experiments. Technometrics, 2(2):149–156, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266541. See discussion [KFT^{+}60].

Daniel:1960:PFR

[Dan60b] Cuthbert Daniel. Parallel fractional replicates. Technometrics, 2(2):263–268, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266549.

Daniel:1961:BRB

[Dan61] Cuthbert Daniel. Book review: A Statistical Manual for Chemists by E. L. Bauer. Technometrics, 3(2):299–300, May 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266119.

Daniel:1963:BRB

[Dan63] Cuthbert Daniel. Book review: Fractional Factorial Designs for Experiments with Factors at Two or Three Levels by W. S. Connor; Shirley Young. Technometrics, 5(2):275–276, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266072.

Daniel:1969:SGR

[Dan69] Cuthbert Daniel. Some general remarks on consulting in statistics. Technometrics, 11(2):241–245, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267257.

Davies:1963:DST

[Dav63] O. L. Davies. The design of screening tests. Technometrics, 5 (4):481–489, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266024.

David:1964:BRB

[Dav64] H. A. David. Book review: Basic Statistics by Thomas E. Kurtz. Technometrics, 6(1):111–112, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266753.

David:1966:BRB

[Dav66] H. A. David. Book review: Handbook of Nonparametric Statistics, II by John E. Walsh. Technometrics, 8(3):553–554, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266705.

deBaun:1964:AR

[dB64] R. M. de Baun. Analysis of residuals. Technometrics, 6(1):127, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266757.

DeBaun:1964:CEA

[De 64] R. M. De Baun. Correction: “The Examination and Analysis of Residuals”, by F. J. Anscombe and John W. Tukey. Technometrics, 6(1):127, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266059. See [AT63].

DeCicco:1965:NAF

[DeC65] Henry DeCicco. Note on an application of four-moment inequalities to a problem in queues. Technometrics, 7(3):435–439, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266600.

DeGray:1968:DI

[DeG68a] R. J. DeGray. Design for interactions. Technometrics, 10(2):389– 391, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267053.

DeGray:1968:RR

[DeG68b] R. J. DeGray. The “round robin”. Technometrics, 10(4):841–842, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267463.

See errata [Ano69f].

Dempster:1960:BRB

[Dem60] A. P. Dempster. Book review: The Analysis of Variance by Henry Scheffé. Technometrics, 2(4):517, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266459.

Dalenius:1967:STS

[DG67] Tore Dalenius and Sakti P. Ghosh. Some theory of sampling when the stratification is subject to error. Technometrics, 9(1):1–13, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266314.

Dyson:1961:GSP

[DGR61] J. Dyson, P. L. Goldsmith, and J. S. M. Robertson. A general simulation programme for material flow in batch chemical plants. Technometrics, 3(4):497–507, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266556.

Draper:1969:TSE

[DH69] Norman R. Draper and William G. Hunter. Transformations: Some examples revisited. Technometrics, 11(1):23–40, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266762.

Draper:1969:WPB

[DHT69] Norman R. Draper, William G. Hunter, and David E. Tierney. Which product is better? Technometrics, 11(2):309–320, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267263. See er-

rata [Ano69f].

Diamond:1967:TDM

[Dia67] W. J. Diamond. Three dimensional models of extreme vertices designs for four component mixtures. Technometrics, 9(3):472–475, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266515.

Dixon:1964:QRO

[Dix64] W. J. Dixon. Query 4: Rejection of outlying values. Technometrics, 6(2):228, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266157.

Draper:1967:SDS

[DL67] Norman R. Draper and Willard E. Lawrence. Sequential designs for spherical weight functions. Technometrics, 9(4):517–529, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266191.

Dodge:1960:BRB

[Dod60] H. F. Dodge. Book review: Inspection and Quality Control Handbook (Interim) H 106, “Multi-Level Continuous Sampling Procedures and Tables for Inspection by Attributes,” 31 October 1958. Technometrics, 2(4):518–520, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266460.

Dodson:1962:BRB

[Dod62] G. A. Dodson. Book review: Quality Control and Reliability Handbook (Interim) H 108, “Sampling Procedures and Tables for Life and Reliability Testing” (Based on Exponential Distribution). Technometrics, 4(1):143, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266183.

Doehlert:1965:BSB

[Doe65] D. H. Doehlert. Balanced sets of balanced incomplete block designs of block size three. Technometrics, 7(4):561–577, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266396.

Dolby:1960:GPF

[Dol60] J. L. Dolby. Graphical procedure for fitting the best line to a set of points. Technometrics, 2(4):477–481, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266455.

Dolby:1963:QMC

[Dol63] James L. Dolby. A quick method for choosing a transformation. Technometrics, 5(3):317–325, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266336.

Doob:1969:BRB

[Doo69] J. L. Doob. Book review: Markov Processes and Potential Theory by Joshua Chover. Technometrics, 11(3):620–621, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267032.

Downton:1966:LEP

[Dow66] Frank Downton. Linear estimates of parameters in the extreme value distribution. Technometrics, 8(1):3–17, February 1966. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266258.

Draper:1961:MVR

[Dra61] Norman R. Draper. Missing values in response surface designs. Technometrics, 3(3):389–398, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266729.

Draper:1962:TOR

[Dra62] Norman R. Draper. Third order rotatable designs in three factors: Analysis. Technometrics, 4(2):219–234, May 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266620.

Draper:1963:BRB

[Dra63a] Norman R. Draper. Book review: Introduction to Probability and Statistics by H. L. Alder; E. B. Roessler. Technometrics, 5(1):132, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266500.

Draper:1963:RAR

[Dra63b] Norman R. Draper. “ridge analysis” of response surfaces. Technometrics, 5(4):469–479, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266023.

Drenick:1963:BRB

[Dre63] R. F. Drenick. Book review: Statistical Theory of Reliability by Marvin Zelen. Technometrics, 5(3):415, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266352.

Draper:1966:TIV

[DS66] Norman R. Draper and David M. Stoneman. Testing for the inclusion of variables in linear regression by a randomisation technique. Technometrics, 8(4):695–699, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266641.

Draper:1968:FCL

[DS68a] Norman R. Draper and David M. Stoneman. Factor changes and linear trends in eight-run two-level factorial designs. Technometrics, 10(2):301–311, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267045.

Draper:1968:RSD

[DS68b] Norman R. Draper and David M. Stoneman. Response surface designs for factors at two and three levels and at two and four levels. Technometrics, 10(1):177–192, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266233.

Dixon:1968:ABD

[DT68] W. J. Dixon and John W. Tukey. Approximate behavior of the distribution of Winsorized t (trimming/Winsorization 2). Technometrics, 10(1):83–98, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266226.

Dubey:1965:APS

[Dub65] Satya D. Dubey. Asymptotic properties of several estimators of Weibull parameters. Technometrics, 7(3):423–434, August 1965.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266599.

Dubey:1967:SPEb

[Dub67a] Satya D. Dubey. On some permissible estimators of the location parameter of the Weibull and certain other distributions. Technometrics, 9(2):293–307, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266425.

Dubey:1967:SPEa

[Dub67b] Satya D. Dubey. Some percentile estimators for Weibull parameters. Technometrics, 9(1):119–129, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266322.

Duncan:1960:BRB

[Dun60] A. J. Duncan. Book review: Quality Control. A Manual of Quality Control Procedure Based upon Scientific Principles and Simplified for Practical Application in Various Types of Manufacturing Plants by Norbert L. Enrick. Technometrics, 2(1):115, February 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266536.

Duncan:1961:BRB

[Dun61] Acheson J. Duncan. Book review: Sampling Inspection Tables, Single and Double Sampling by Harold F. Dodge; Harry G. Romig. Technometrics, 3(1):119–120, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266483.

Duncan:1962:BSP

[Dun62] Acheson J. Duncan. Bulk sampling: Problems and lines of attack. Technometrics, 4(3):319–344, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266571.

Dunn:1964:MCU

[Dun64] Olive Jean Dunn. Multiple comparisons using rank sums. Technometrics, 6(3):241–252, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266041.

Duncan:1965:BAM

[Dun65] David B. Duncan. A Bayesian approach to multiple comparisons. Technometrics, 7(2):171–222, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266670.

Dunnett:1966:TFW

[Dun66] Charles W. Dunnett. A tribute to Frank Wilcoxon 1892–1965. Technometrics, 8(1):195–196, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266273.

Durbin:1966:PPC

[Dur66] Eugene P. Durbin. Pricing policies contingent on observed product quality. Technometrics, 8(1):123–134, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266267.

Daniel:1966:FPR

[DW66a] Cuthbert Daniel and Frank Wilcoxon. Factorial 2^{p−q} plans robust against linear and quadratic trends. Technometrics, 8(2):259–278, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266360.

Dishon:1966:CSR

[DW66b] Menachem Dishon and George H. Weiss. A communications satellite replenishment policy. Technometrics, 8(3):399–409, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266687.

Dwyer:1964:MIS

[Dwy64] P. S. Dwyer. Matrix inversion with the square root method. Technometrics, 6(2):197–213, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266152.

Dykstra:1960:PDR

[Dyk60] O. Dykstra, Jr. Partial duplication of response surface designs. Technometrics, 2(2):185–195, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266544.

Dykstra:1965:BRBa

[Dyk65a] Otto Dykstra, Jr. Book review: Statistical Theory and Methodology in Science and Engineering by K. A. Brownlee. Technometrics, 7(3):451–453, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266606.

Dykstra:1965:BRBb

[Dyk65b] Otto Dykstra, Jr. Book review: Theory and Technique of Variation Research by Leonard G. Johnson. Technometrics, 7(4):654–655, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266407.

Dykstra:1966:BRB

[Dyk66a] Otto Dykstra, Jr. Book review: Statistical Assessment of the Life Characteristic by W. R. Buckland. Technometrics, 8(2):371, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266374.

Dykstra:1966:OUE

[Dyk66b] Otto Dykstra, Jr. The orthogonalization of undesigned experiments. Technometrics, 8(2):279–290, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266361. See errata [Ano66h].

Dykstra:1967:BRBa

[Dyk67a] Otto Dykstra, Jr. Book review: Analysis of Straight-Line Data by Forman S. Acton; Some Theory of Sampling by W. Edwards Deming. Technometrics, 9(3):493, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266521.

Dykstra:1967:BRBb

[Dyk67b] Otto Dykstra, Jr. Book review: Statistical Methods and Formulae by C. G. Lambe. Technometrics, 9(4):702, November 1967. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266214.

Ehrenfeld:1962:ESS

[EBT62] S. Ehrenfeld and S. Ben-Tuvia. The efficiency of statistical simulation procedures. Technometrics, 4(2):257–275, May 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266623.

Eckler:1969:SCP

[Eck69] A. Ross Eckler. A survey of coverage problems associated with point and area targets. Technometrics, 11(3):561–589, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267025.

Eisen:1963:ESD

[Eis63] M. M. Eisen. Effects of slow-downs and failure on stochastic service systems. Technometrics, 5(3):385–392, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266341.

Eisenberger:1964:GBD

[Eis64] Isidore Eisenberger. Genesis of bimodal distributions. Technometrics, 6(4):357–363, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266090.

Eisenberger:1968:TMS

[Eis68] I. Eisenberger. Testing the mean and standard deviation of a normal distribution using quantiles. Technometrics, 10(4):781–792, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267459.

Elandt:1961:FND

[Ela61] Regina C. Elandt. The folded normal distribution: Two methods of estimating parameters from moments. Technometrics, 3 (4):551–562, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266561.

Ellner:1963:VRS

[Ell63] Henry Ellner. Validating results of sampling inspection by attributes. Technometrics, 5(1):23–46, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266489.

Esary:1964:SSE

[EM64] J. D. Esary and A. W. Marshall. System structure and the existence of a system life. Technometrics, 6(4):459–462, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266099.

Eilbott:1965:PLT

[EN65] Joan Eilbott and Jack Nadler. On precedence life testing. Technometrics, 7(3):359–377, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266595.

Enrick:1960:VFA

[Enr60] Norbert L. Enrick. Variations flow analysis. Technometrics, 2 (3):373–386, August 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266251.

Esary:1963:CSN

[EP63a] J. D. Esary and F. Proschan. Coherent structures of non-identical components. Technometrics, 5(2):191–209, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266063.

Esary:1963:RBS

[EP63b] James D. Esary and Frank Proschan. Relationship between system failure rate and component failure rates. Technometrics, 5(2):183– 189, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266062.

Epstein:1960:ETE

[Eps60a] Benjamin Epstein. Elements of the theory of extreme values. Technometrics, 2(1):27–41, February 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266529.

Epstein:1960:ELT

[Eps60b] Benjamin Epstein. Estimation from life test data. Technometrics, 2(4):447–454, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266453.

Epstein:1960:EPT

[Eps60c] Benjamin Epstein. Estimation of the parameters of two parameter exponential distributions from censored samples. Technometrics, 2(3):403–406, August 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266254.

Epstein:1960:SLT

[Eps60d] Benjamin Epstein. Statistical life test acceptance procedures. Technometrics, 2(4):435–446, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266452.

Epstein:1960:TVAa

[Eps60e] Benjamin Epstein. Tests for the validity of the assumption that the underlying distribution of life is exponential. Part I. Technometrics, 2(1):83–101, February 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266533.

Epstein:1960:TVAb

[Eps60f] Benjamin Epstein. Tests for the validity of the assumption that the underlying distribution of life is exponential: Part II. Technometrics, 2(2):167–183, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266543.

Epstein:1961:EBR

[Eps61] Benjamin Epstein. Estimates of bounded relative error for the mean life of an exponential distribution. Technometrics, 3(1): 107–109, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266481.

Epstein:1964:BRB

[Eps64a] Benjamin Epstein. Book review: Sampling Procedures and Tables for Life and Reliability Testing Based on the Weibull Distribution (Hazard Rate Criterion); Sampling Procedures and Tables for Life and Reliability Testing Based on the Weibull Distribution (Reliable Life Criterion). Technometrics, 6(2):229–230, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266159.

Epstein:1964:CQC

[Eps64b] Benjamin Epstein. Comment on “Quality Control and Reliability Handbook: H108”. Technometrics, 6(1):128, February 1964. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266758.

Ericson:1966:BRB

[Eri66] William A. Ericson. Book review: Introductory Probability and Statistical Applications by Paul L. Meyer. Technometrics, 8(4): 720–722, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266656.

El-Sayyad:1969:ISE

[ES69] G. M. El-Sayyad. Information and sampling from the exponential distribution. Technometrics, 11(1):41–45, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266763.

Evans:1963:AMS

[Eva63] David H. Evans. Applied multiplex sampling. Technometrics, 5 (3):341–359, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266338.

Evans:1967:ANI

[Eva67] David H. Evans. An application of numerical integration techniques to statistical tolerancing. Technometrics, 9(3):441–456, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266512.

Ewan:1963:WHU

[Ewa63] W. D. Ewan. When and how to use Cu-Sum charts. Technometrics, 5(1):1–22, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266488.

Fisher:1960:PPD

[FC60] Ronald A. Fisher and E. A. Cornish. The percentile points of distributions having known cumulants. Technometrics, 2(2):209– 225, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266546.

See errata [Ano60e].

Feigenbaum:1966:BRB

[Fei66] A. V. Feigenbaum. Book review: More… through Quality by J. van Ettinger; J. Sittig. Technometrics, 8(4):714, November 1966. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266648.

Ferris:1966:BRB

[Fer66] George E. Ferris. Book review: An Introduction to Mathematical Learning Theory by R. C. Atkinson; G. H. Bower; E. J. Crothers. Technometrics, 8(2):382, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266381.

Finney:1963:SPD

[Fin63] D. J. Finney. Some properties of a distribution specified by its cumulants. Technometrics, 5(1):63–69, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266491. See errata [Ano63i].

Fischer:1967:BRB

[Fis67] Patrick C. Fischer. Book review: Computer Literature Bibliography 1946–1963 by W. W. Youden. Technometrics, 9(3):493, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266520.

Fleming:1963:BRB

[Fle63] W. Fleming. Book review: Theory of Markov Processes by E. B. Dynkin; D. E. Brown; T. Kovary. Technometrics, 5(4):530, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266034.

Flehinger:1965:PTP

[Fle65] B. J. Flehinger. Product test planning for repairable systems. Technometrics, 7(4):485–494, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266389.

Foradori:1966:BRB

[For66] George T. Foradori. Book review: Bibliography of Statistical Literature 1940–1949 by Maurice G. Kendall; Alison G. Doig. Technometrics, 8(3):547, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266700.

Fowlkes:1969:SOA

[Fow69] E. B. Fowlkes. Some operators for ANOVA calculations. Technometrics, 11(3):511–526, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267022.

Fox:1963:GRS

[Fox63] Bennett L. Fox. Generation of random samples from the beta F distributions. Technometrics, 5(2):269–270, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266068.

Folks:1965:EFA

[FPS65] John Leroy Folks, Donald A. Pierce, and Charles Stewart. Estimating the fraction of acceptable product. Technometrics, 7(1):43–50, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266126.

Fraser:1968:BBC

[Fra68] D. A. S. Fraser. A black box or a comprehensive model. Technometrics, 10(2):219–229, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267040.

Fry:1961:FNF

[Fry61] R. E. Fry. Finding new fractions of factorial experimental designs. Technometrics, 3(3):359–370, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266727. See errata [Ano63h].

Flanagan:1969:NIS

[FVM69] P. D. Flanagan, P. A. Vitale, and J. Mendelsohn. A numerical investigation of several one-dimensional search procedures in nonlinear regression problems. Technometrics, 11(2):265–284, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267260.

Faulkenberry:1968:SSD

[FW68] G. David Faulkenberry and David L. Weeks. Sample size determination for tolerance limits. Technometrics, 10(2):343–348, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267049.

Gaffey:1965:BRB

[Gaf65] William R. Gaffey. Book review: Basic Statistics: A Primer for the Biomedical Sciences by Olive Jean Dunn. Technometrics, 7(2):263– 264, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266678.

Gani:1969:BRB

[Gan69] J. Gani. Book review: Combinatorial Identities by J. Riordan. Technometrics, 11(2):406–408, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267277.

Gaver:1963:RHR

[Gav63] D. P. Gaver, Jr. Random hazard in reliability problems. Technometrics, 5(2):211–226, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266064.

Gardiner:1965:AS

[GB65] Donald A. Gardiner and Barbara Flores Bombay. An approximation to Student’s t. Technometrics, 7(1):71–72, February 1965. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266131.

Greenwood:1960:AFG

[GD60] J. Arthur Greenwood and David Durand. Aids for fitting the gamma distribution by maximum likelihood. Technometrics, 2 (1):55–65, February 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266531.

Geisser:1962:BRB

[Gei62] Seymour Geisser. Book review: An Elementary Introduction to the Theory of Probability by B. V. Gnedenko; A. Ya. Khinchin; W. R. Stahl; T. B. Roberts. Technometrics, 4(4):615–616, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266295.

Geisser:1963:BRB

[Gei63] Seymour Geisser. Book review: Statistical Theory by B. W. Lindgren. Technometrics, 5(3):411, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266349.

Geisser:1964:BRB

[Gei64] Seymour Geisser. Book review: Probability Theory and Mathematical Statistics by M. Fisz. Technometrics, 6(4):473, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266103.

Geisser:1966:BRB

[Gei66] Seymour Geisser. Book reviews: Introduction to Probability and Statistics from a Bayesian Viewpoint. Part 1. Probability by D. V. Lindley; Introduction to Probability and Statistics from a Bayesian ViewpointPart 2. Inference by D. V. Lindley. Technometrics, 8 (2):375–376, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266377.

Gabriel:1969:DSS

[GF69] K. R. Gabriel and Paul Feder. On the distribution of statistics suitable for evaluating rainfall stimulation experiments. Technometrics, 11(1):149–160, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266771.

Gorman:1962:SLD

[GH62] J. W. Gorman and J. E. Hinman. Simplex lattice designs for multicomponent systems. Technometrics, 4(4):463–487, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266284.

Gardiner:1966:ATS

[GH66] Donald A. Gardiner and Norma C. Hull. An approximation to two-sided tolerance limits for normal populations. Technometrics, 8(1):115–122, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266266.

Gaylor:1969:EDF

[GH69] D. W. Gaylor and F. N. Hopper. Estimating the degrees of freedom for linear combinations of mean squares by Satterthwaite’s formula. Technometrics, 11(4):691–706, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266893.

Ghosh:1967:BRB

[Gho67] B. K. Ghosh. Book review: An Introduction to Sets, Probability and Hypothesis Testing by H. F. Fehr; L. N. H. Bunt; G. Grossman. Technometrics, 9(1):181–182, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266328.

Gilbert:1963:BRB

[Gil63] E. N. Gilbert. Book review: Combinatorial Chance by F. N. David; D. E. Barton. Technometrics, 5(2):289–290, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266083.

Gilbert:1965:BRB

[Gil65] E. N. Gilbert. Book review: Applied Combinatorial Mathematics by Edwin F. Beckenbach. Technometrics, 7(3):451, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266605.

Gilbert:1966:BRB

[Gil66] E. N. Gilbert. Book review: The Compleat Strategyst by J. D. Williams. Technometrics, 8(3):548–549, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266702.

Gajjar:1969:PCS

[GK69] A. V. Gajjar and C. G. Khatri. Progressively censored samples from log-normal and logistic distributions. Technometrics, 11(4): 793–803, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266899.

Goodman:1961:FRS

[GKKK61] N. R. Goodman, S. Katz, B. H. Kramer, and M. T. Kuo. Frequency response from stationary noise: Two case histories. Technometrics, 3(2):245–268, May 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266116.

Gray:1967:CIA

[GL67] H. L. Gray and T. O. Lewis. A confidence interval for the availability ratio. Technometrics, 9(3):465–471, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266514.

Glasser:1962:MVU

[Gla62] Gerald J. Glasser. Minimum variance unbiased estimators for Poisson probabilities. Technometrics, 4(3):409–418, August 1962. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266576.

Glasser:1967:ARP

[Gla67] Gerald J. Glasser. The age replacement problem. Technometrics, 9(1):83–91, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266320.

Goodman:1962:PFN

[GM62] Leo A. Goodman and Albert Madansky. Parameter-free and nonparametric tolerance limits: The exponential case. Technometrics, 4(1):75–95, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266174.

Guttman:1965:BMN

[GM65] Irwin Guttman and Duane A. Meeter. On Beale’s measures of nonlinearity. Technometrics, 7(4):623–637, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266400.

Gumbel:1966:CEC

[GM66] E. J. Gumbel and C. K. Mustafi. Comments to: Edward C. Posner, “The Application of Extreme Value Theory to Error Free Communication”. Technometrics, 8(2):363–366, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266371. See [Pos65].

Gaylor:1968:AED

[GM68] D. W. Gaylor and J. A. Merrill. Augmenting existing data in multiple regression. Technometrics, 10(1):73–81, February 1968. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266225.

Goldman:1963:BRB

[Gol63] A. J. Goldman. Book review: Systems: Research and Design by Donald P. Eckmann. Technometrics, 5(1):130–131, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266499.

Goodman:1961:SCS

[Goo61] N. R. Goodman. Some comments on spectral analysis of time series. Technometrics, 3(2):221–228, May 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266113. See comments [JP61].

Good:1962:CFA

[Goo62] I. J. Good. A classification of fallacious arguments and interpretations. Technometrics, 4(1):125–132, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266177.

Goodman:1965:SCI

[Goo65] Leo A. Goodman. On simultaneous confidence intervals for multinomial proportions. Technometrics, 7(2):247–254, May 1965. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266673.

Good:1969:SAS

[Goo69] I. J. Good. Some applications of the singular decomposition of a matrix. Technometrics, 11(4):823–831, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266902.

Gorman:1966:DEV

[Gor66] J. W. Gorman. Discussion of “Extreme Vertices Design of Mixture Experiments” by R. A. McLean and V. L. Anderson. Technometrics, 8(3):455–456, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266692. See [MA66].

Govindarajulu:1963:RAM

[Gov63] Z. Govindarajulu. Relationships among moments of order statistics in samples from two related populations. Technometrics, 5 (4):514–518, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266028.

Govindarajulu:1966:BRB

[Gov66] Zakkula Govindarajulu. Book review: The Theory of Stochastic Processes by D. R. Cox; H. D. Miller. Technometrics, 8(4):718, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266653.

Gnanadesikan:1967:MLE

[GPH67] R. Gnanadesikan, R. S. Pinkham, and Laura P. Hughes. Maximum likelihood estimation of the parameters of the beta distribution from smallest order statistics. Technometrics, 9(4):607–620, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266199.

Gupta:1967:BLU

[GQS67] S. S. Gupta, A. S. Qureishi, and B. K. Shah. Best linear unbiased estimators of the parameters of the logistic distribution using order statistics. Technometrics, 9(1):43–56, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266317.

Grace:1969:BRB

[Gra69] K. Grace, Jr. Book review: Probabilistic Reliability: An Engineering Approach by M. L. Shooman. Technometrics, 11(1):206, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266781.

Greene:1968:BRB

[Gre68] Charles H. Greene. Book review: Introduction to Business Statistics by G. Hadley. Technometrics, 10(4):875–876, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267474.

Grubbs:1967:BRB

[Gru67] Frank E. Grubbs. Book review: An Introduction to Probability Theory and Its Applications, Volume II by William Feller. Technometrics, 9(2):342, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266435.

Grubbs:1969:PDO

[Gru69] Frank E. Grubbs. Procedures for detecting outlying observations in samples. Technometrics, 11(1):1–21, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266761.

Guttman:1969:IRD

[GS69] Irwin Guttman and Dennis E. Smith. Investigation of rules for dealing with outliers in small samples from the normal distribution: I: Estimation of the mean. Technometrics, 11(3):527–550, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267023.

Gorman:1966:SVF

[GT66] J. W. Gorman and R. J. Toman. Selection of variables for fitting equations to data. Technometrics, 8(1):27–51, February 1966. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266260.

Guenther:1969:MSB

[Gue69] William C. Guenther. Modified sampling, binomial and hypergeometric cases. Technometrics, 11(4):639–647, November 1969. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266888.

Gunther:1965:QCN

[Gun65] Paul Gunther. Query 13: Combination of a normal and a uniform distribution. Technometrics, 7(3):449–450, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266604.

Gupta:1960:OSG

[Gup60] Shanti S. Gupta. Order statistics from the gamma distribution. Technometrics, 2(2):243–262, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266548. See errata [Ano60e].

Gupta:1962:LTS

[Gup62] Shanti S. Gupta. Life test sampling plans for normal and lognormal distributions. Technometrics, 4(2):151–175, May 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266616.

Gupta:1965:SMD

[Gup65] Shanti S. Gupta. On some multiple decision (selection and ranking) rules. Technometrics, 7(2):225–245, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266672.

Gurnow:1965:NGW

[Gur65] R. N. Gurnow. A note on G. S. Watson’s paper “A study of the group screening method”. Technometrics, 7(3):444–446, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266602. See

[Wat61].

Guterman:1962:UBS

[Gut62] H. E. Guterman. An upper bound for the sample standard deviation. Technometrics, 4(1):134–135, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266179. See errata [Ano62f,

Ano63i].

Goldsmith:1961:ARL

[GW61] P. L. Goldsmith and H. Whitfield. Average run lengths in cumulative chart quality control schemes. Technometrics, 3(1):11–20, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266473.

See errata [Ano61e].

Haggstrom:1966:BRB

[Hag66] Gus W. Haggstrom. Book review: Elements of the Theory of Probability by Émile Borel; John E. Freund. Technometrics, 8(1):197, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266274.

Hald:1960:CHD

[Hal60] A. Hald. The compound hypergeometric distribution and a system of single sampling inspection plans based on prior distributions and costs. Technometrics, 2(3):275–340, August 1960. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266247. See discussion

[BLH^{+}60].

Hald:1962:SLT

[Hal62] A. Hald. Some limit theorems for the Dodge–Romig LTPD single sampling inspection plans. Technometrics, 4(4):497–513, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266286.

Halperin:1963:ANC

[Hal63] Max Halperin. Approximations to the non-central t, with applications. Technometrics, 5(3):295–305, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266334.

Halperin:1964:SWT

[Hal64] Max Halperin. Some waiting time distributions for redundant systems with repair. Technometrics, 6(1):27–40, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266744. See er-

rata [Ano64h].

Halperin:1966:CIC

[Hal66] Max Halperin. Confidence intervals from censored samples, II. Technometrics, 8(2):291–301, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266362.

Hald:1967:DSS

[Hal67] A. Hald. The determination of single sampling attribute plans with given Producer’s and Consumer’s risk. Technometrics, 9 (3):401–415, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266509.

Hald:1968:BSS

[Hal68] A. Hald. Bayesian single sampling attribute plans for continuous prior distributions. Technometrics, 10(4):667–683, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267451.

Hammer:1962:BRB

[Ham62] Preston C. Hammer. Book review: Finite-Difference Methods for Partial Differential Equations by George E. Forsythe; Wolfgang R. Wasow. Technometrics, 4(1):143–144, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266184.

Hammer:1965:BRBb

[Ham65a] Preston C. Hammer. Book review: Experimental Arithmetic, High Speed Computing and Mathematics by N. C. Metropolis, A. H. Taub, John Todd, and C. B. Tompkins. Technometrics, 7(1):82, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://links.jstor.org/sici?sici=
0040-1706%28196502%297%3A1%3C82%3AEAHSCA%3E2.0.CO%3B2-
9.

Hammer:1965:BRBa

[Ham65b] Preston C. Hammer. Book review: Experimental Arithmetic, High Speed Computing and Mathematics by N. C. Metropolis; A. H. Taub; John Todd; C. B. Tompkins. Technometrics, 7(1):82, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266139.

Han:1969:MLE

[Han69] Chien-Pai Han. Maximum likelihood estimate in intraclass correlation model. Technometrics, 11(4):833–834, November 1969. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266903.

Harter:1961:UTP

[Har61a] H. Leon Harter. Use of tables of percentage points of range and Studentized range. Technometrics, 3(3):407–411, August 1961. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266731.

Hartley:1961:MGN

[Har61b] H. O. Hartley. The modified Gauss–Newton method for the fitting of non-linear regression functions by least squares. Technometrics, 3(2):269–280, May 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266117.

Harris:1962:BRB

[Har62] Bernard Harris. Book review: The Passage Problem for a Stationary Markov Chain by J. H. B. Kemperman. Technometrics, 4(4):617–618, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266297.

Harter:1964:ECB

[Har64] H. Leon Harter. Exact confidence bounds, based on one order statistic, for the parameter of an exponential population. Technometrics, 6(3):301–317, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266047. See errata [Ano64h].

Harter:1966:BRB

[Har66] H. Leon Harter. Book review: Percentage Points of the Sum V_{1}^{(s)} of s Roots (s = 1–50) (A Unified Table for Tests of Significance in Various Univariate and Multivariate Hypotheses) by Tito A. Mijares. Technometrics, 8(1):199–200, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266276.

Harkness:1967:BRB

[Har67a] William L. Harkness. Book review: Sampling from a Graph by A. R. Bloemena. Technometrics, 9(2):343, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266436.

Harter:1967:MLE

[Har67b] H. Leon Harter. Maximum-likelihood estimation of the parameters of a four-parameter generalized gamma population from complete and censored samples. Technometrics, 9(1):159–165, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266326. See

errata [Ano67h, Ano68f].

Harter:1969:NTP

[Har69] H. Leon Harter. A new table of percentage points of the Pearson Type III distribution. Technometrics, 11(1):177–187, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266773.

Hasselblad:1966:EPM

[Has66] Victor Hasselblad. Estimation of parameters for a mixture of normal distributions. Technometrics, 8(3):431–444, August 1966. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266689. See discussion

[Coh66a].

Hayton:1968:BRB

[Hay68] Thomas Hayton. Book review: The Advanced Theory of Statistics, Vol. 3 by M. G. Kendall; A. Stuart. Technometrics, 10(1):211–212, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266239.

Halperin:1960:ESB

[HB60] M. Halperin and G. L. Burrows. The effect of sequential batching for acceptance-rejection sampling upon sample assurance of total product quality. Technometrics, 2(1):19–26, February 1960. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266528.

Halperin:1961:ADO

[HB61] M. Halperin and G. L. Burrows. An asymptotic distribution for an occupancy problem with statistical applications. Technometrics, 3(1):79–89, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1266478. See errata [Ano61f].

Holms:1969:CPA

[HB69] A. G. Holms and J. N. Berrettoni. Chain-pooling ANOVA for two-level factorial replication-free experiments. Technometrics, 11 (4):725–746, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266895.

Healy:1961:BRB

[Hea61] M. J. R. Healy. Book review: Time Series Analysis by E. J. Hannan. Technometrics, 3(3):440–441, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266737.

Healy:1969:ETS

[Hea69] M. J. R. Healy. Exact tests of significance in contingency tables. Technometrics, 11(2):393–395, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267270.

Hesse:1968:BRB

[Hes68] Richard M. Hesse. Book review: Recent Advances in Optimization Techniques by A. Lavi; T. P. Vogl. Technometrics, 10(3):622–623, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267118.

Hewett:1968:NPI

[Hew68] John E. Hewett. A note on prediction intervals based on partial observations in certain life test experiments. Technometrics, 10(4):850–853, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267465.

Harris:1966:APM

[HH66a] W. A. Harris, Jr. and T. N. Helvig. Applications of the pseudoinverse to modeling. Technometrics, 8(2):351–357, May 1966. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266368.

Hill:1966:RRS

[HH66b] William J. Hill and William G. Hunter. A review of response surface methodology: A literature survey. Technometrics, 8(4): 571–590, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266632.

Hill:1969:NDM

[HH69] William J. Hill and William G. Hunter. A note on designs for model discrimination: Variance unknown case. Technometrics, 11(2):396– 400, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267271.

Hill:1968:JDC

[HHW68] William J. Hill, William G. Hunter, and Dean W. Wichern. A joint design criterion for the dual problem of model discrimination and parameter estimation. Technometrics, 10(1):145–160, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266231.

Hicks:1968:BRB

[Hic68] Charles R. Hicks. Book review: Elementary Statistics: A Workbook by K. Hope. Technometrics, 10(2):416–417, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267067.

Hill:1960:EDA

[Hil60] Hubert M. Hill. Experimental designs to adjust for time trends. Technometrics, 2(1):67–82, February 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266532.

Hillier:1962:SPL

[Hil62] Frederick S. Hillier. Surveillance programs for lots in storage. Technometrics, 4(4):515–530, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266287.

Hillier:1964:NCS

[Hil64] Frederick S. Hillier. New criteria for selecting continuous sampling plans. Technometrics, 6(2):161–178, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266149.

Hinich:1967:ESA

[Hin67] Melvin Hinich. Estimation of spectra after hard clipping of Gaussian processes. Technometrics, 9(3):391–400, August 1967. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266508.

Hinchen:1968:MRP

[Hin68] J. D. Hinchen. Multiple regression in process development. Technometrics, 10(2):257–269, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267042.

Hunter:1966:EOR

[HK66] William G. Hunter and J. R. Kittrell. Evolutionary operation: A review. Technometrics, 8(3):389–397, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266686.

Hocking:1967:SBS

[HL67] R. R. Hocking and R. N. Leslie. Selection of the best subset in regression analysis. Technometrics, 9(4):531–540, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266192.

Harter:1965:MLE

[HM65a] H. Leon Harter and Albert H. Moore. Maximum-likelihood estimation of the parameters of gamma and Weibull populations from complete and from censored samples. Technometrics, 7(4):639–643, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266401.

Harter:1965:PIE

[HM65b] H. Leon Harter and Albert H. Moore. Point and interval estimators, based on m order statistics, for the scale parameter of a Weibull population with known shape parameter. Technometrics, 7(3):405–422, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266598.

Harter:1967:NET

[HM67] H. Leon Harter and Albert H. Moore. A note on estimation from a Type I extreme-value distribution. Technometrics, 9(2):325–331, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266430.

Harter:1968:CML

[HM68] H. Leon Harter and Albert H. Moore. Conditional maximumlikelihood estimators, from singly censored samples, of the scale parameters of Type II extreme-value distributions. Technometrics, 10(2):349–359, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267050.

Hanson:1963:DFT

[HO63] D. L. Hanson and D. B. Owen. Distribution-free tolerance limits elimination of the requirement that cumulative distribution functions be continuous. Technometrics, 5(4):518–522, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266029.

Hoel:1968:STS

[Hoe68a] David G. Hoel. Sequential testing of sample size. Technometrics, 10(2):331–341, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267048.

Hoel:1968:TDP

[Hoe68b] Paul G. Hoel. On testing for the degree of a polynomial. Technometrics, 10(4):757–767, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267457.

Hogben:1966:BRB

[Hog66] David Hogben. Book review: The World of Probability: Statistics in Science by Solomon Diamond. Technometrics, 8(4):725–727, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266660.

Holla:1967:RET

[Hol67] M. S. Holla. Reliability estimation of the truncated exponential model. Technometrics, 9(2):332–336, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266431.

Holmes:1968:BRB

[Hol68] Paul T. Holmes. Book review: Introduction to Probability and Statistical Decision Theory by G. Hadley. Technometrics, 10(3):625– 627, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1267121.

Honeycutt:1966:BRB

[Hon66] J. M. Honeycutt. Book review: Computer Control of Industrial Processes by E. S. Savas. Technometrics, 8(3):556–557, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266708.

Hunter:1965:DDB

[HR65] William G. Hunter and Albey M. Reiner. Designs for discriminating between two rival models. Technometrics, 7(3):307–323, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266591.

Heller:1966:DRL

[HS66] R. A. Heller and M. Shinozuka. Development of randomized load sequences with transition probabilities based on a Markov process. Technometrics, 8(1):107–114, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266265.

Hsi:1966:OQC

[Hsi66] Bartholomew P. Hsi. Optimization of quality control in the chemical laboratory. Technometrics, 8(3):519–534, August 1966. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266697.

Hunter:1966:LE

[HTJ^{+}66] J. Stuart Hunter, Leo J. Tick, Richard H. Jones, Allan G. Piersol, Maurice B. Priestley, and C. H. Gibson. Letters to the Editor. Technometrics, 8(3):559–568, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266710.

Hunter:1964:SFE

[Hun64] J. S. Hunter. Sequential factorial estimation. Technometrics, 6 (1):41–55, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1266745. See errata [Ano65k].

Hunter:1966:IYA

[Hun66] J. S. Hunter. The inverse Yates algorithm. Technometrics, 8(1): 177–183, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266269.

Hussain:1969:BRB

[Hus69] K. M. Hussain. Book review: Introduction to Quantitative Management by George J. Brabb. Technometrics, 11(2):403, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267272.

Hartley:1966:QJC

[HW66] H. O. Hartley and A. W. Wortham. Query 17: Joint confidence limits for ranked observations. Technometrics, 8(2):368–370, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266373.

Ireland:1965:BRB

[Ire65] Clarence Ireland. Book review: Fundamental Concepts in the Design of Experiments by Charles R. Hicks. Technometrics, 7(4): 652–653, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266405.

Johnson:1967:QAF

[JA67] Arthur F. Johnson and Ram Avrahmi. Query 20: Analysis of a factorial experiment (partially confounded 2[3]) (additional responses). Technometrics, 9(3):490–492, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266519. See comment [Smi68a].

Jackson:1965:CPK

[Jac65] J. Edward Jackson. Comments on paper by Kurtz, Link, Tukey and Wallace. Technometrics, 7(2):163–165, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266667. See [KLTW65b].

Jackson:1966:BRB

[Jac66] J. Edward Jackson. Book review: The Statistical Analysis of Experimental Data by John Mandel. Technometrics, 8(1):197–199, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266275.

Jaech:1964:EWD

[Jae64a] J. L. Jaech. Estimation of Weibull distribution shape parameter when no more than two failures occur per lot. Technometrics, 6(4):415–422, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266096.

Jaech:1964:PEM

[Jae64b] J. L. Jaech. A program to estimate measurement error in nondestructive evaluation of reactor fuel element quality. Technometrics, 6(3):293–300, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266046.

Jaech:1965:EPS

[Jae65] J. L. Jaech. Estimation of particle size distribution based on observed weights of groups of particles. Technometrics, 7(4):505–515, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266391.

Jaech:1967:IMN

[Jae67] J. L. Jaech. Interference in the manufacture of nuclepore filters. Technometrics, 9(2):319–324, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266429.

Jackson:1961:STT

[JB61] J. Edward Jackson and Ralph A. Bradley. Sequential χ^{2}- and T^{2}-tests and their application to an acceptance sampling problem. Technometrics, 3(4):519–534, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266558.

Jorgensen:1964:LE

[JCP64] Erik Jørgensen, Regne Centralen, and G. P. Patil. Letters to the Editor. Technometrics, 6(4):477–478, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266107.

Jeffreys:1963:BRB

[Jef63] Harold Jeffreys. Book review: The Foundations of Statistical Inference by L. J. Savage. Technometrics, 5(3):407–410, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266347.

Jenkins:1961:GCA

[Jen61] G. M. Jenkins. General considerations in the analysis of spectra. Technometrics, 3(2):133–166, May 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266110.

Jenkins:1962:BRB

[Jen62] G. M. Jenkins. Book reviews: Stochastic Processes, Problems and Solutions by Lajos Takacs; Stochastic Population Models by M. S. Bartlett. Technometrics, 4(4):615, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266294.

Jenkins:1963:EEL

[Jen63] G. M. Jenkins. An example of the estimation of a linear open loop transfer function. Technometrics, 5(2):227–245, May 1963. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266065.

Johns:1966:EAE

[JL66] M. V. Johns, Jr. and G. J. Lieberman. An exact asymptotically efficient confidence bound for reliability in the case of the Weibull distribution. Technometrics, 8(1):135–175, February 1966. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266268.

Jackson:1967:QRR

[JL67] J. Edward Jackson and William H. Lawton. Query 22: Regression residual analysis. Technometrics, 9(2):339–340, May 1967. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266433.

Jackson:1969:CAH

[JL69] J. Edward Jackson and William H. Lawton. Comparison of ANOVA and harmonic components of variance. Technometrics, 11(1):75–90, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266766.

John:1961:ABI

[Joh61] Peter W. M. John. An application of a balanced incomplete block design. Technometrics, 3(1):51–54, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266476.

Johnson:1962:ESS

[Joh62a] N. L. Johnson. Estimation of sample size. Technometrics, 4 (1):59–67, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266172.

Johnson:1962:FND

[Joh62b] N. L. Johnson. The folded normal distribution: Accuracy of estimation by maximum likelihood. Technometrics, 4(2):249–256, May 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266622.

Johnson:1963:CSC

[Joh63] N. L. Johnson. Cumulative sum control charts for the folded normal distribution. Technometrics, 5(4):451–458, November 1963. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266021.

John:1964:BD

[Joh64a] Peter W. M. John. Blocking of 3(2^{n−k}) designs. Technometrics, 6(4):371–376, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266092.

Johnson:1964:QDR

[Joh64b] N. L. Johnson. Query 5: The distribution of a ranked observation. Technometrics, 6(3):329–330, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266050.

John:1966:AD

[Joh66a] Peter W. M. John. Augmenting 2^{n−1} designs. Technometrics, 8 (3):469–480, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266694.

Johnson:1966:BRB

[Joh66b] N. L. Johnson. Book review: Linear Statistical Inference and Its Applications by C. Radhakrishna Rao. Technometrics, 8(3):551– 553, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1266704.

Johnson:1966:CSC

[Joh66c] N. L. Johnson. Cumulative sum control charts and the Weibull distribution. Technometrics, 8(3):481–491, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266695.

Johnson:1966:PCR

[Joh66d] N. L. Johnson. Paths and chains of random straight-line segments. Technometrics, 8(2):303–317, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266363.

Johnson:1968:BRB

[Joh68] Richard A. Johnson. Book review: Sequential Methods in Statistics by G. B. Wetherill. Technometrics, 10(2):405–406, May 1968. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267058.

Jones:1965:RPS

[Jon65] Richard H. Jones. A reappraisal of the periodogram in spectral analysis. Technometrics, 7(4):531–542, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266393.

Joshi:1964:SPH

[Jos64] Madhukar Joshi. Statistical programs for high speed computers. Technometrics, 6(2):235–237, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266165.

Jenkins:1961:CDM

[JP61] G. M. Jenkins and Emanuel Parzen. Comments on the discussions of Messrs. Tukey and Goodman. Technometrics, 3(2):229– 234, May 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266114.

See [Tuk61, Goo61].

Johnson:1968:LE

[JRK^{+}68] Norman L. Johnson, Dave Rothman, Richard G. Krutchkoff, Peter A. Lachenbruch, and R. R. Hocking. Letters to the Editor. Technometrics, 10(2):429–433, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267069.

Jennrich:1968:ASR

[JS68a] R. I. Jennrich and P. F. Sampson. Application of stepwise regression to non-linear estimation. Technometrics, 10(1):63–72, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266224.

Joffe:1968:CST

[JS68b] A. D. Joffe and H. S. Sichel. A chart for sequentially testing observed arithmetic means from lognormal populations against a given standard. Technometrics, 10(3):605–612, August 1968. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267115.

Kussmaul:1967:EVC

[KA67] Keith Kussmaul and R. L. Anderson. Estimation of variance components in two-stage nested designs with composite samples. Technometrics, 9(3):373–389, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266507.

Kanwal:1961:BRB

[Kan61] R. P. Kanwal. Book review: Cartesian Tensors: An Introduction by G. Temple. Technometrics, 3(4):570, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266565.

Kao:1964:QST

[Kao64] John H. K. Kao. Query 7: Savings in test time when comparing Weibull scale parameters. Technometrics, 6(4):471, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266102.

Kao:1969:NPA

[Kao69] Edward P. C. Kao. A non-parametric approach to the 2^{3} factorial design. Technometrics, 11(1):193–196, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266775.

Kapadia:1966:CII

[Kap66] C. H. Kapadia. Combining intra and inter block analysis of group divisible designs. Technometrics, 8(1):189–191, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266271.

Kawata:1965:BRB

[Kaw65] Tatsuo Kawata. Book review: Probability Theory by M. Loève. Technometrics, 7(1):85–87, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266142.

Keilson:1963:BRB

[Kei63] J. Keilson. Book review: Introduction to the Theory of Queues by Lajos Takacs. Technometrics, 5(2):286, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266078.

Kemp:1967:EEI

[Kem67] Kenneth W. Kemp. An example of errors incurred by erroneously assuming normality for Cusum schemes. Technometrics, 9(3):457– 464, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1266513.

Kerridge:1967:EPM

[Ker67] D. Kerridge. Errors of prediction in multiple regression with stochastic regressor variables. Technometrics, 9(2):309–311, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266426.

Kruskal:1960:DPM

[KFT^{+}60] William Kruskal, Thomas S. Ferguson, John W. Tukey, E. J. Gumbel, and F. J. Anscombe. Discussion of the papers of Messrs. Anscombe and Daniel. Technometrics, 2(2):157–166, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266542. See

[AG60, Dan60a].

Kipiniak:1965:BRB

[Kip65] Walerian Kipiniak. Book review: Optimal Estimation, Identification, and Control by R. C. K. Lee. Technometrics, 7(4):653–654, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266406.

Kullback:1962:TCT

[KKK62] S. Kullback, M. Kupperman, and H. H. Ku. Tests for contingency tables and Markov chains. Technometrics, 4(4):573–608, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266291.

Kurtz:1965:ARA

[KLTW65a] T. E. Kurtz, R. F. Link, J. W. Tukey, and D. L. Wallace. Authors’ reply to Anscombe’s comments. Technometrics, 7(2):169, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266669. See

[Ans65].

Kurtz:1965:SCMa

[KLTW65b] T. E. Kurtz, R. F. Link, J. W. Tukey, and D. L. Wallace. Short-cut multiple comparisons for balanced single and double classifications: Part 1, Results. Technometrics, 7(2):95–161, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266666.

Reprinted in [Cox92, pp. 353–430]. See comments [Jac65, Ans65], and Part 2 [KLTW65c].

Kurtz:1965:SCMb

[KLTW65c] T. E. Kurtz, R. F. Link, J. W. Tukey, and D. L. Wallace. Shortcut multiple comparisons for balanced single and double classifications: Part 2. Derivations and approximations. Biometrika, 52 (3/4):485–498, December 1965. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2333700. Reprinted in [Cox92, pp. 431–452].

Karson:1969:MBE

[KMH69] M. J. Karson, A. R. Manson, and R. J. Hader. Minimum bias estimation and experimental design for response surfaces. Technometrics, 11(3):461–475, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267020.

Knott:1967:SMP

[Kno67] M. Knott. Sampling mixtures of particles. Technometrics, 9 (3):365–371, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266506.

Koch:1967:GAE

[Koc67a] Gary G. Koch. A general approach to the estimation of variance components. Technometrics, 9(1):93–118, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266321. See remarks [Koc68].

Koch:1967:PEP

[Koc67b] Gary G. Koch. A procedure to estimate the population mean in random effects models. Technometrics, 9(4):577–585, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266196.

Koch:1968:SFR

[Koc68] Gary G. Koch. Some further remarks concerning “A General Approach to the Estimation of Variance Components”. Technometrics, 10(3):551–558, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267108. See [Koc67a].

Koehler:1965:BRB

[Koe65] Truman L. Koehler. Book review: Experimental Statistics by Mary G. Natrella. Technometrics, 7(4):651, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266404.

Koopmans:1964:BRB

[Koo64] L. H. Koopmans. Book review: Probabilities on Algebraic Structures by Ulf Grenander. Technometrics, 6(3):336–337, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266055.

Koopmans:1968:BRB

[Koo68] L. H. Koopmans. Book review: Stationary Random Processes by Y. A. Rozanov. Technometrics, 10(3):628–630, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267124.

Korfhage:1968:BRB

[Kor68] Robert R. Korfhage. Book review: Computer and Information Sciences — II by J. T. Tou. Technometrics, 10(3):631–633, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267126.

Krishnaiah:1967:NMG

[KR67] P. R. Krishnaiah and M. Haseeb Rizvi. A note on moments of gamma order statistics. Technometrics, 9(2):315–318, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266428.

Katti:1968:BRB

[KR68] S. K. Katti and A. Vijaya Rao. Book review: Handbook of the Poisson Distribution by F. A. Haight. Technometrics, 10(2):412, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267063.

Kraemer:1963:OSC

[Kra63a] Helena Chmura Kraemer. One-sided confidence intervals for the quality indices of a complex item. Technometrics, 5(3):400–403, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266345.

Krane:1963:ASD

[Kra63b] Scott A. Krane. Analysis of survival data by regression techniques. Technometrics, 5(2):161–174, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266060.

Kramer:1966:BRB

[Kra66a] Clyde Y. Kramer. Book review: Multivariate Analysis for Biologists by H. L. Seal. Technometrics, 8(3):549–551, August 1966. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266703.

Kramer:1966:ACD

[Kra66b] Clyde Young Kramer. Approximation to the cumulative tdistribution. Technometrics, 8(2):358–359, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266369.

Kramer:1967:UMD

[Kra67] K. H. Kramer. Use of mean deviation in the analysis of interlaboratory tests. Technometrics, 9(1):149–153, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266324.

Kruskal:1960:SRW

[Kru60] William H. Kruskal. Some remarks on wild observations. Technometrics, 2(1):1–3, February 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266526.

Kruskal:1963:BRB

[Kru63] William Kruskal. Book review: Analysing Qualitative Data by A. E. Maxwell. Technometrics, 5(2):271–273, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266069.

Krutchkoff:1966:CUS

[Kru66] Richard G. Krutchkoff. The correct use of the sample mean absolute deviation in confidence intervals for a normal variate. Technometrics, 8(4):663–674, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266638.

Krutchkoff:1967:CIR

[Kru67] R. G. Krutchkoff. Classical and inverse regression methods of calibration. Technometrics, 9(3):425–439, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266511.

Krutchkoff:1969:CIR

[Kru69] Richard G. Krutchkoff. Classical and inverse regression methods of calibration in extrapolation. Technometrics, 11(3):605–608, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267027.

Kennard:1969:CAD

[KS69] R. W. Kennard and L. A. Stone. Computer aided design of experiments. Technometrics, 11(1):137–148, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266770.

Ku:1963:BRB

[Ku63a] H. H. Ku. Book review: An Introduction to Mathematical Statistics by H. D. Brunk. Technometrics, 5(2):278–279, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266074.

Ku:1963:NCT

[Ku63b] H. H. Ku. A note on contingency tables involving zero frequencies and the 2Î test. Technometrics, 5(3):398–400, August 1963. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266344.

Kussmaul:1969:PAA

[Kus69] Keith Kussmaul. Protection against assuming the wrong degree in polynomial regression. Technometrics, 11(4):677–682, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266891.

Lachenbruch:1966:DAW

[Lac66] Peter A. Lachenbruch. Discriminant analysis when the initial samples are misclassified. Technometrics, 8(4):657–662, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266637.

Larson:1966:LSE

[Lar66] Harold J. Larson. Least squares estimation of the components of a symmetric matrix. Technometrics, 8(2):360–362, May 1966. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266370.

Laurent:1968:BRB

[Lau68] Andre G. Laurent. Book review: Introduction to Probability and Statistics by B. W. Lindgren; G. W. McElrath. Technometrics, 10(2):406–408, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267059.

Lawrence:1966:IBP

[Law66] Michael J. Lawrence. An investigation of the burn-in problem. Technometrics, 8(1):61–71, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266262.

Leaverton:1969:SSP

[LB69] Paul Leaverton and John J. Birch. Small sample power curves for the two sample location problem. Technometrics, 11(2):299–307, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267262.

Lipow:1964:ABN

[LE64] M. Lipow and R. L. Eidemiller. Application of the bivariate normal distribution to a stress vs strength problem in reliability analysis. Technometrics, 6(3):325–328, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266049.

Leaverton:1967:BRB

[Lea67] Paul Leaverton. Book review: Statistics in Biology, Volume I by C. I. Bliss. Technometrics, 9(4):702–703, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266215.

Leimkuhler:1968:BRB

[Lei68] Ferdinand F. Leimkuhler. Book review: Introduction to Operations Research by F. S. Hillier; G. J. Lieberman. Technometrics, 10(2):410–411, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267061.

Leone:1961:SPHa

[Leo61a] Fred C. Leone. Statistical programs for high speed computers. Technometrics, 3(1):123–126, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266485.

Leone:1961:SPHb

[Leo61b] Fred C. Leone. Statistical programs for high speed computers. Technometrics, 3(2):301–304, May 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266120. See correction [Ano61d].

Leone:1961:SPHc

[Leo61c] Fred C. Leone. Statistical programs for high speed computers. Technometrics, 3(3):443–446, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266739.

Leone:1961:SPHd

[Leo61d] Fred C. Leone. Statistical programs for high speed computers. Technometrics, 3(4):571–574, November 1961. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266566.

Leone:1962:SPHa

[Leo62a] Fred C. Leone. Statistical programs for high speed computers. Technometrics, 4(1):145–149, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266186.

Leone:1962:SPHb

[Leo62b] Fred C. Leone. Statistical programs for high speed computers. Technometrics, 4(2):289–292, May 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266627.

Leone:1962:SPHc

[Leo62c] Fred C. Leone. Statistical programs for high speed computers. Technometrics, 4(3):433–435, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266581.

Lerwick:1965:MLE

[Ler65] Trygve R. Lerwick. Maximum likelihood estimators of regression coefficients for the case of autocorrelated residuals. Technometrics, 7(1):51–58, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266127.

Levin:1966:BRB

[Lev66] Sheldon G. Levin. Book review: Certain Studies on the Structure and Statistics of the Logarithmic Series Distribution and Related Tables by G. P. Patil; A. R. Kamat; J. K. Wani. Technometrics, 8(4):723–724, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266658.

Li:1968:BRB

[Li68] C. C. Li. Book review: A Programmed Introduction to Statistics by F. F. Elzey. Technometrics, 10(2):411, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267062.

Lieberman:1961:PRS

[Lie61] Gerald J. Lieberman. Prediction regions for several predictions from a single regression line. Technometrics, 3(1):21–27, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266474.

Lieberman:1965:SPC

[Lie65] Gerald J. Lieberman. Statistical process control and the impact of automatic process control. Technometrics, 7(3):283–292, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266589.

Liggett:1969:BRB

[Lig69] Thomas M. Liggett. Book review: Probability Measures on Metric Spaces by K. R. Parthasarathy. Technometrics, 11(4):839–840, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266907.

Linhart:1964:DST

[Lin64] H. Linhart. On the distribution of some time averages of shot noise. Technometrics, 6(3):287–292, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266045.

Lawton:1969:SMI

[LJ69a] William H. Lawton and J. Edward Jackson. The spectrum of a Model II nested ANOVA and its applications. Technometrics, 11(1):91–102, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266767.

Little:1969:NGP

[LJ69b] R. E. Little and Emil H. Jebe. A note on the gain in precision for optimal allocation in regession as applied to extrapolation in S-N fatigue testing. Technometrics, 11(2):389–392, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267269.

Leone:1967:SRE

[LJE67] Fred C. Leone, Toke Jayachandran, and Stanley Eisenstat. A study of robust estimators. Technometrics, 9(4):652–660, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266203.

Lloyd:1961:BRB

[Llo61] Emlyn Lloyd. Book review: Statistical Theory and Methodology in Science and Engineering by K. A. Brownlee. Technometrics, 3(4):569, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266563.

Lloyd:1963:RSC

[Llo63] E. H. Lloyd. Reservoirs with serially correlated inflows. Technometrics, 5(1):85–93, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266493.

Lachenbruch:1968:EER

[LM68] Peter A. Lachenbruch and M. Ray Mickey. Estimation of error rates in discriminant analysis. Technometrics, 10(1):1–11, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266219. See

commentary [Coc68a].

Lipow:1964:QSV

[LMW64] Myron Lipow, Nathan Mantel, and J. W. Wilkinson. Query 2: The sum of values from a normal and a truncated normal distribution (continued). Technometrics, 6(4):469–471, November 1964. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266101.

Leone:1966:SDV

[LN66] Fred C. Leone and Lloyd S. Nelson. Sampling distributions of variance components. I. Empirical studies of balanced nested designs. Technometrics, 8(3):457–468, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266693.

Leone:1968:SDV

[LNJE68] Fred C. Leone, Lloyd S. Nelson, Norman L. Johnson, and Stanley Eisenstat. Sampling distributions of variance components II. Empirical studies of unbalanced nested designs. Technometrics, 10(4):719–737, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267455.

Leone:1961:FND

[LNN61] F. C. Leone, L. S. Nelson, and R. B. Nottingham. The folded normal distribution. Technometrics, 3(4):543–550, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266560. See cor-

rigenda [Ano80].

Low:1965:BRB

[Low65] Leone Y. Low. Book review: Analysis of Variance by William C. Guenther. Technometrics, 7(3):453–454, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266608.

Laubscher:1968:ECV

[LSL68] Nico F. Laubscher, F. E. Steffens, and Elsie M. De Lange. Exact critical values for Mood’s distribution-free test statistic for dispersion and its normal approximation. Technometrics, 10(3):497–507, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267104.

Lukacs:1963:BRB

[Luk63] Eugene Lukacs. Book review: Selected Translations in Mathematical Statistics and Probability. Technometrics, 5(2):287–288, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266080.

Lukacs:1964:BRBa

[Luk64a] Eugene Lukacs. Book review: Selected Translations in Mathematical Statistics and Probability, Vol. 3. Technometrics, 6(2):230–231, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266160.

Lukacs:1964:BRBb

[Luk64b] Eugene Lukacs. Book review: Selected Translations in Mathematical Statistics and Probability Vol. 4. Technometrics, 6(3):334–336, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266054.

Lutz:1968:BRB

[Lut68] Raymond P. Lutz. Book review: Decision Rules for Inventory Management by Robert G. Brown. Technometrics, 10(2):408–409, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267060.

McLean:1966:EVD

[MA66] R. A. McLean and V. L. Anderson. Extreme vertices design of mixture experiments. Technometrics, 8(3):447–454, August 1966. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266691. See discussion

[Gor66].

Mack:1969:QTL

[Mac69] C. Mack. Query 30: Tolerance limits for a binomial distribution. Technometrics, 11(1):201, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266777.

Madansky:1965:ACL

[Mad65] Albert Madansky. Approximate confidence limits for the reliability of series and parallel systems. Technometrics, 7(4):495–503, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266390.

Madow:1968:BRB

[Mad68] William G. Madow. Book review: Elementary Sampling Theory by Taro Yamane. Technometrics, 10(3):621–622, August 1968. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267117.

Maisel:1966:BRB

[Mai66] Herbert Maisel. Book review: A First Course in Numerical Analysis by Anthony Ralston. Technometrics, 8(3):555–556, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266707.

Mallios:1969:GAI

[Mal69] William S. Mallios. A generalized application of instrumental variable estimation to straight-line relations when both variables are subject to error. Technometrics, 11(2):255–263, May 1969. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267259.

Mann:1961:BRB

[Man61] H. B. Mann. Book review: Error Correcting Codes by W. W. Peterson. Technometrics, 3(4):570, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266564.

Mansfield:1962:PFC

[Man62] Edwin Mansfield. Power functions for Cox’s test of randomness against trend. Technometrics, 4(3):430–432, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266580.

Mandel:1964:EWF

[Man64a] John Mandel. Estimation of weighting factors in linear regression and analysis of variance. Technometrics, 6(1):1–25, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266743.

Mandel:1964:QRA

[Man64b] John Mandel. Query 3: Regression analysis of cumulative data. Technometrics, 6(2):225–227, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266156.

Mann:1964:BRB

[Man64c] H. B. Mann. Book review: Combinatorial Mathematics by H. J. Ryser. Technometrics, 6(1):112–113, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266755.

Mann:1967:TOB

[Man67] Nancy R. Mann. Tables for obtaining the best linear invariant estimates of parameters of the Weibull distribution. Technometrics, 9(4):629–645, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266201.

Mann:1968:PIE

[Man68] Nancy R. Mann. Point and interval estimation procedures for the two-parameter Weibull and extreme-value distributions. Technometrics, 10(2):231–256, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267041.

Mandel:1969:MFE

[Man69a] John Mandel. A method for fitting empirical surfaces to physical or chemical data. Technometrics, 11(3):411–429, August 1969. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267017.

Mantel:1969:RLS

[Man69b] Nathan Mantel. Restricted least squares regression and convex quadratic programming. Technometrics, 11(4):763–773, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266897.

Marsaglia:1964:GVT

[Mar64] George Marsaglia. Generating a variable from the tail of the normal distribution. Technometrics, 6(1):101–102, February 1964. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266749.

Margolin:1967:SMA

[Mar67] Barry H. Margolin. Systematic methods for analyzing 2^{n} 3^{m} factorial experiments with applications. Technometrics, 9(2):245–259, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266421.

Marcus:1968:SC

[Mar68a] Allan H. Marcus. Stochastic coalescence. Technometrics, 10(1): 133–143, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266230.

Margolin:1968:OME

[Mar68b] Barry H. Margolin. Orthogonal main-effect 2^{n} 3^{m} designs and two-factor interaction aliasing. Technometrics, 10(3):559–573, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267109.

Margolin:1969:OME

[Mar69a] Barry H. Margolin. Orthogonal main-effect plans permitting estimation of all two-factor interactions for the 2^{n}3^{m} factorial series of designs. Technometrics, 11(4):747–762, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266896.

Margolin:1969:RFD

[Mar69b] Barry H. Margolin. Results on factorial designs of resolution IV for the 2^{n} and 2^{n}3^{m} series. Technometrics, 11(3):431–444, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267018.

Mohan:1964:EME

[MBM64] R. Mohan, A. K. Bhattacharji, and R. Mishra. Estimating machining errors in set-ups with automatic resetting. Technometrics, 6(4):423–438, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266097.

McCool:1965:CGL

[McC65] John I. McCool. The construction of good linear unbiased estimates from the best linear estimates for a smaller sample size. Technometrics, 7(4):543–552, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266394.

McKean:1968:BRBb

[McK68a] H. E. McKean. Book review: Tables of the Incomplete Beta Function by Karl Pearson; E. S. Pearson; N. L. Johnson. Technometrics, 10(4):879–880, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267480.

McKean:1968:BRBa

[McK68b] H. E. McKean. Book review: The Design and the Analysis of Scientific Experiments by K. C. Peng. Technometrics, 10(3):627–628, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267123.

McKean:1969:BRBb

[McK69a] Harlley E. McKean. Book review: Basic Statistics by David Blackwell. Technometrics, 11(4):837, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266904.

McKean:1969:BRBc

[McK69b] Harlley E. McKean. Book review: Introduction to Stochastic Processes in Biostatistics by Chin Long Chiang. Technometrics, 11 (4):837–838, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266905.

McKean:1969:BRBd

[McK69c] Harlley E. McKean. Book review: Statistical Tables by F. James Rohlf; Robert R. Sokal. Technometrics, 11(4):841–842, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266909.

McKean:1969:BRBa

[McK69d] Harlley E. McKean. Book review: Topics in Intermediate Statistical Materials, Vol. 1 by T. A. Bancroft. Technometrics, 11(2):403– 404, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267273.

McLean:1969:BRB

[McL69] Robert A. McLean. Book review: Statistical Analysis for Business Decisions by W. S. Peters; G. W. Summers. Technometrics, 11(1):203–205, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266779.

Mehta:1968:ARD

[MD68] J. S. Mehta and M. N. Das. Asymmetric rotatable designs and orthogonal transformations. Technometrics, 10(2):313–322, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267046.

Meeter:1966:BRB

[Mee66] Duane Meeter. Book review: Introduction to Numerical Analysis by C. E. Fröberg. Technometrics, 8(4):719–720, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266655.

Meeter:1969:BRB

[Mee69] Duane Meeter. Book review: Stochastic Approximation and Nonlinear Regression by A. E. Albert; L. A. Gardner. Technometrics, 11(1):203, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266778.

Menon:1963:ESS

[Men63] M. V. Menon. Estimation of the shape and scale parameters of the Weibull distribution. Technometrics, 5(2):175–182, May 1963. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266061.

McGilchrist:1968:TMD

[MH68] C. A. McGilchrist and J. Y. Harrison. Testing of means with different alternatives. Technometrics, 10(1):195–198, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266235.

Milton:1969:CEN

[MH69] Roy C. Milton and Richard Hotchkiss. Computer evaluation of the normal and inverse normal distribution functions. Technometrics, 11(4):817–822, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266901.

Mielke:1967:NSS

[Mie67] Paul W. Mielke, Jr. Note on some squared rank tests with existing ties. Technometrics, 9(2):312–314, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266427.

Miller:1960:IPS

[Mil60] H. D. Miller. Inter-plant storage in continuous manufacturing. Technometrics, 2(3):393–401, August 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266253.

Miller:1963:BRB

[Mil63] H. D. Miller. Book review: Stochastic Processes by Emanuel Parzen. Technometrics, 5(4):528–529, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266032.

Miller:1966:BRB

[Mil66] Irwin Miller. Book review: Probability, Random Variables, and Stochastic Processes by Anthanasios Papoulis. Technometrics, 8 (2):378–380, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266379.

Minton:1965:BRBb

[Min65a] Paul D. Minton. Book review: Introduction to Statistical Inference by Harold Freeman. Technometrics, 7(4):658, November 1965. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266410.

Minton:1965:BRBa

[Min65b] Paul D. Minton. Book review: Statistics for Scientists and Engineers by R. Lowell Wine. Technometrics, 7(4):657–658, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266409.

Mendenhall:1960:ANM

[ML60] W. Mendenhall and E. H. Lehman, Jr. An approximation to the negative moments of the positive binomial useful in life testing. Technometrics, 2(2):227–242, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266547.

Moder:1962:SSP

[Mod62] Joseph J. Moder. A sequential search procedure for locating a response jump. Technometrics, 4(4):610–614, November 1962. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266293.

Mohan:1967:DPL

[Moh67] R. Mohan. Diamond-pin location. Technometrics, 9(1):131–148, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266323.

See errata [Ano67g].

Mohan:1969:CDC

[Moh69] R. Mohan. Controlling dimension in centerless-grinding with automatic reset device. Technometrics, 11(1):115–136, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266769.

Morrison:1961:OAS

[Mor61] Donald F. Morrison. The optimum allocation of spare components in systems. Technometrics, 3(3):399–406, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266730.

Moses:1965:QCL

[Mos65] Lincoln E. Moses. Query 10: Confidence limits from rank tests. Technometrics, 7(2):257–260, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266675.

Marsaglia:1968:QPR

[MPB68] George Marsaglia, Minu K. Patel, and Duncan Badenius. Query 27: Pseudo random normal numbers. Technometrics, 10(2):401–404, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267057.

Myhre:1968:CTM

[MS68] J. M. Myhre and Sam C. Saunders. Comparison of two methods of obtaining approximate confidence intervals for system reliability. Technometrics, 10(1):37–49, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266222.

Morrison:1965:SSC

[MT65] Milton Morrison and Filbert Tobias. Some statistical characteristics of a peak to average ratio. Technometrics, 7(3):379–385, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266596.

Muller:1965:BRB

[Mul65] Mervin E. Muller. Book review: Methods of Statistical Testing: Monte Carlo Method by Yu. A. Shreider. Technometrics, 7(4): 659–661, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266412.

Myers:1964:MEC

[Mye64] R. H. Myers. Methods for estimating the composition of a three component liquid mixture. Technometrics, 6(4):343–356, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266089.

Nadler:1967:QBS

[Nad67] Jack Nadler. Query 24: Bivariate samples with missing values. Technometrics, 9(4):679–682, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266206.

Nadas:1969:CRL

[Nád69a] Arthur Nádas. A confidence region for the log-normal hazard function. Technometrics, 11(2):387–388, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267268.

Nadler:1969:BRB

[Nad69b] Jack Nadler. Book review: Slippage Tests by R. Doornbos. Technometrics, 11(2):404, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267274.

Naus:1966:PCT

[Nau66] Joseph I. Naus. A power comparison of two tests of non-random clustering. Technometrics, 8(3):493–517, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266696.

Naus:1968:BRB

[Nau68] Joseph I. Naus. Book review: Classical and Contagious Discrete Distributions by G. P. Patil. Technometrics, 10(1):213–214, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266241.

Nelson:1960:LE

[Nel60] Lloyd S. Nelson. Letters to the Editor. Technometrics, 2(4):522, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266461.

Nelson:1963:TPL

[Nel63] Lloyd S. Nelson. Tables for a precedence life test. Technometrics, 5(4):491–499, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266025.

Nelson:1966:QCV

[Nel66] Lloyd S. Nelson. Query 19: Combining values of observed chi squares. Technometrics, 8(4):709, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266644.

Nelson:1968:STE

[Nel68] Wayne Nelson. A statistical test for equality of two availabilities. Technometrics, 10(3):594–596, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267113.

Neuts:1968:BRB

[Neu68] Marcel F. Neuts. Book review: Waiting-Line Models: An Introduction to Their Theory and Application by E. Ruiz-Pala; C. Avila-Beloso; W. W. Hines. Technometrics, 10(2):413–414, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267065.

Ney:1965:BRB

[Ney65] P. E. Ney. Book review: The Theory of Branching Processes by T. E. Harris. Technometrics, 7(1):79–80, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266137.

Nadler:1968:LE

[NF68] Jack Nadler and Bennett Fox. Letters to the Editor. Technometrics, 10(4):883–884, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267482.

Neave:1968:MCS

[NG68] H. R. Neave and C. W. J. Granger. A Monte Carlo study comparing various two-sample tests for differences in mean. Technometrics, 10(3):509–522, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267105.

Nuri:1969:FMS

[NH69] W. A. Nuri and L. J. Herbst. Fourier methods in the study of variance fluctuations in time series analysis. Technometrics, 11 (1):103–113, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266768.

Nicholson:1969:UPS

[NM69] W. L. Nicholson and K. R. Merckx. Unfolding particle size distributions. Technometrics, 11(4):707–723, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266894.

Noble:1963:BRB

[Nob63] Ben Noble. Book review: Studies in Mathematical Analysis and Related Topics: Essays in Honor of George Pólya by Szegő; Loewner; Bergman; Schiffer; Neyman; Gilbarg; Solomon. Technometrics, 5(4):531, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266036.

Noether:1963:BRB

[Noe63] Gottfried E. Noether. Book review: Handbook of Nonparametric Statistics. Investigation of Randomness, Moments, Percentiles, and Distributions by John E. Walsh. Technometrics, 5(2):273–274, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266070.

Nelson:1963:EPD

[NWF63] A. C. Nelson, Jr., J. S. Williams, and N. T. Fletcher. Estimation of the probability of defective failure from destructive tests. Technometrics, 5(4):459–468, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266022.

Odeh:1966:BRB

[Ode66a] R. E. Odeh. Book review: Studies in the Theory of Random Processes by A. V. Skorokhod. Technometrics, 8(4):718–719, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266654.

Odell:1966:BRB

[Ode66b] Patrick L. Odell. Book review: How to Gamble If You Must: Inequalities for Stochastic Processes by L. E. Dubins; L. J. Savage. Technometrics, 8(4):713, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266647.

Odeh:1967:DMS

[Ode67] Robert E. Odeh. The distribution of the maximum sum of ranks. Technometrics, 9(2):271–278, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266423.

Odell:1968:BRB

[Ode68] Patrick L. Odell. Book review: Digital Computer User’s Handbook by M. Klerer; G. A. Korn. Technometrics, 10(4):876–877, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267475.

Ott:1969:SCR

[OF69] R. Lyman Ott and Spencer M. Free. A short-cut rule for a onesided test of hypothesis for qualitative data. Technometrics, 11 (1):197–200, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266776.

Olmstead:1962:BRB

[Olm62] Paul S. Olmstead. Book review: Progress in Operations Research, Volume I by R. L. Ackoff. Technometrics, 4(1):144, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266185.

Olmstead:1963:CEA

[Olm63] Paul Olmstead. Correction: “The Examination and Analysis of Residuals”, by F. J. Anscombe and John W. Tukey. Technometrics, 5(4):536–537, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). See [AT63].

Ott:1968:OED

[OM68] R. Lyman Ott and Raymond H. Myers. Optimal experimental designs for estimating the independent variable in regression. Technometrics, 10(4):811–823, November 1968. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267461.

Ostle:1964:BRB

[Ost64] Bernard Ostle. Book review: Principles of Reliability by Erich Pieruschka. Technometrics, 6(2):232, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266162.

Owen:1963:BRB

[Owe63] D. B. Owen. Book reviews: An Index of Mathematical Tables by A. Fletcher; J. C. P. Miller; L. Rosenhead; L. J. Comrie; Guide to Tables in Mathematical Statistics by J. Arthur Greenwood; H. O. Hartley. Technometrics, 5(2):276–278, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266073.

Owen:1964:CPB

[Owe64] D. B. Owen. Control of percentages in both tails of the normal distribution. Technometrics, 6(4):377–387, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266093.

Owen:1965:BRBa

[Owe65a] D. B. Owen. Book review: Handbook of Mathematical Functions with Formulas, Graphs, and Mathematical Tables by Milton Abramowitz; Irene A. Stegun. Technometrics, 7(1):78–79, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266136.

Owen:1965:BRBb

[Owe65b] D. B. Owen. Book review: Tables for Normal Sampling with Unknown Variance. The Student Distribution and Economically Optimal Sampling Plans by Jerome Bracken; Arthur Schleifer. Technometrics, 7(1):82–83, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266140.

Owen:1965:BRBc

[Owe65c] D. B. Owen. Book review: Tables for Testing Significance in a 2×2 Contingency Table by D. J. Finney; R. Latscha; B. M. Bennett; P. Hsu. Technometrics, 7(2):264–265, May 1965. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266680. See errata [Ano66j].

Owen:1966:BRB

[Owe66a] D. B. Owen. Book review: Supplement to Tables for Testing Significance in a 2×2 Contingency Table by B. M. Bennett; C. Horst. Technometrics, 8(4):715, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266649.

Owen:1966:LE

[Owe66b] D. B. Owen. Letter to the Editor. Technometrics, 8(4):729–730, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266662.

Owen:1967:VSP

[Owe67] D. B. Owen. Variables sampling plans based on the normal distribution. Technometrics, 9(3):417–423, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266510.

Owen:1968:SPA

[Owe68] D. B. Owen. A survey of properties and applications of the noncentral t-distribution. Technometrics, 10(3):445–478, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267101.

Owen:1969:SRW

[Owe69] D. B. Owen. Summary of recent work on variables acceptance sampling with emphasis on non-normality. Technometrics, 11(4): 631–637, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266887.

Page:1961:CSC

[Pag61] E. S. Page. Cumulative sum charts. Technometrics, 3(1):1–9, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266472.

Page:1962:CSS

[Pag62] E. S. Page. Cumulative sum schemes using gauging. Technometrics, 4(1):97–109, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266175.

Page:1963:CSD

[Pag63] E. S. Page. Controlling the standard deviation by cusums and warning lines. Technometrics, 5(3):307–315, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266335.

Parzen:1961:MCE

[Par61] Emanuel Parzen. Mathematical considerations in the estimation of spectra. Technometrics, 3(2):167–190, May 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266111.

Parzen:1964:BRB

[Par64] Emanuel Parzen. Book review: Topics in the Theory of Random Noise. Vol. 1. General Theory of Random Processes. Non-Linear Transformations of Signals and Noise by R. L. Stratonovich; R. A. Silverman. Technometrics, 6(4):473–474, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266104.

Pasternack:1962:LEA

[Pas62] Bernard S. Pasternack. Linear estimation and the analysis of gamma ray pulse-height spectra. Technometrics, 4(4):565–571, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266290.

Patil:1960:ENB

[Pat60] G. P. Patil. On the evaluation of the negative binomial distribution with examples. Technometrics, 2(4):501–505, November 1960. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266457.

Patel:1962:GSM

[Pat62] M. S. Patel. Group-screening with more than two stages. Technometrics, 4(2):209–217, May 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266619.

Patel:1963:NWP

[Pat63a] M. S. Patel. A note on Watson’s paper. Technometrics, 5(3):397– 398, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1266343. See [Wat61].

Patel:1963:PDF

[Pat63b] M. S. Patel. Partially duplicated fractional factorial designs. Technometrics, 5(1):71–83, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266492.

Patil:1963:EBI

[Pat63c] G. P. Patil. On the equivalence of the binomial and inverse binomial acceptance sampling plans and an acknowledgement. Technometrics, 5(1):119–121, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266495.

Patterson:1966:DIE

[Pat66] R. L. Patterson. Difficulties involved in the estimation of a population mean using transformed sample data. Technometrics, 8 (3):535–537, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266698.

Phatak:1967:EFD

[PB67] A. G. Phatak and N. M. Bhatt. Estimation of the fraction defective in curtailed sampling plans by attributes. Technometrics, 9(2):219– 228, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266419.

Pelto:1969:ANC

[Pel69] Chester R. Pelto. Adaptive nonparametric classification. Technometrics, 11(4):775–792, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266898.

Peterson:1964:BRB

[Pet64] Donald P. Peterson. Book review: Trucking of Radioactive Materials. Safety vs Economy in Highway Transport by Ferdinard F.

Leimkuhler. Technometrics, 6(4):474–475, November 1964. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266105.

Petersen:1966:BRB

[Pet66] Roger G. Petersen. Book review: Matrix Algebra for the Biological Sciences by S. R. Searle. Technometrics, 8(4):722–723, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266657.

Pfanzagl:1963:BRB

[Pfa63a] J. Pfanzagl. Book review: The Advanced Theory of Statistics, Vol. 2 by M. G. Kendall; A. Stuart. Technometrics, 5(4):525–528, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266031.

Pfanzagl:1963:SPB

[Pfa63b] J. Pfanzagl. Sampling procedures based on prior distributions and costs. Technometrics, 5(1):47–61, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266490.

Pfanzagl:1964:BRB

[Pfa64] J. Pfanzagl. Book review: Essays on Probability and Statistics by M. S. Bartlett. Technometrics, 6(1):112, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266754.

Priestley:1965:EPS

[PG65] M. B. Priestley and C. H. Gibson. Estimation of power spectra by a wave-analyser. Technometrics, 7(4):553–559, November 1965. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266395.

Phatak:1968:CSC

[Pha68a] A. G. Phatak. Censored sampling in curtailed sampling plans by attributes. Technometrics, 10(4):854–860, November 1968. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267466.

Phatak:1968:MDC

[Pha68b] A. G. Phatak. Misclassified data from curtailed sampling plans. Technometrics, 10(3):489–496, August 1968. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267103.

Pruzan:1967:DPA

[PJ67] Peter Mark Pruzan and J. T. Ross Jackson. A dynamic programming application in production line inspection. Technometrics, 9 (1):73–81, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266319.

Pasternack:1965:PRT

[PL65] Bernard Pasternack and Anthony Liuzzi. Patterns in residuals: A test for regression model adequacy in radionuclide assay. Technometrics, 7(4):603–621, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266399.

Prairie:1961:ASR

[PO61] R. R. Prairie and B. Ostle. An analysis of some relay failure data from a composite exponential population. Technometrics, 3 (3):423–428, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266733.

Page:1963:LE

[PO63] E. S. Page and Paul S. Olmstead. Letters to the Editor. Technometrics, 5(4):535–537, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266038.

Posten:1963:RUB

[Pos63] H. O. Posten. Robustness of uniform Bayesian encoding. Technometrics, 5(1):122–125, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266496.

Posner:1965:AEV

[Pos65] Edward C. Posner. The application of extreme-value theory to error-free communication. Technometrics, 7(4):517–529, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266392. See

comments [GM66].

Posten:1966:BRB

[Pos66] H. O. Posten. Book review: Multidimensional Gaussian Distributions by Kenneth S. Miller. Technometrics, 8(2):377–378, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266378.

Potthoff:1962:FFA

[Pot62a] Richard F. Potthoff. Four-factor additive designs more general than the Greco–Latin square. Technometrics, 4(3):361–366, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266573.

Potthoff:1962:TFA

[Pot62b] Richard F. Potthoff. Three-factor additive designs more general than the Latin square. Technometrics, 4(2):187–208, May 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266618.

Potthoff:1967:BRB

[Pot67] Richard F. Potthoff. Book review: The Analysis of Variance by A. Huitson. Technometrics, 9(2):341–342, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266434.

Prairie:1967:PAT

[Pra67] R. R. Prairie. Probit analysis as a technique for estimating the reliability of a simple system. Technometrics, 9(2):197–203, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266417.

Preece:1966:ARP

[Pre66] D. A. Preece. On Addelman’s 2^{17−9} resolution V plan. Technometrics, 8(4):705–707, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266643.

Preece:1969:BIB

[Pre69] D. A. Preece. Balanced incomplete block designs with sets of identical blocks. Technometrics, 11(3):613–615, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267029.

Priestley:1962:BCE

[Pri62] M. B. Priestley. Basic considerations in the estimation of spectra. Technometrics, 4(4):551–564, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266289.

Proschan:1963:TEO

[Pro63] Frank Proschan. Theoretical explanation of observed decreasing failure rate. Technometrics, 5(3):375–383, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266340.

Putter:1965:BRB

[Put65] Joseph Putter. Book review: Basic Concepts of Probability and Statistics by J. L. Hodges,; E. L. Lehmann. Technometrics, 7 (1):80–81, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266138.

Parr:1965:MDB

[PW65] Van B. Parr and J. T. Webster. A method for discriminating between failure density functions used in reliability predictions. Technometrics, 7(1):1–10, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266123.

Pyke:1963:BRB

[Pyk63] Ronald Pyke. Book review: Statistical Inference for Markov Processes by Patrick Billingsley. Technometrics, 5(3):413–415, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266351.

Prairie:1962:SAS

[PZB62] R. R. Prairie, W. J. Zimmer, and J. K. Brookhouse. Some acceptance sampling plans based on the theory of runs. Technometrics, 4(2):177–185, May 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266617.

Quesenberry:1964:LSS

[QH64] C. P. Quesenberry and D. C. Hurst. Large sample simultaneous confidence intervals for multinomial proportions. Technometrics, 6(2):191–195, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266151.

Qureishi:1965:SIP

[QNA65] A. S. Qureishi, K. J. Nabavian, and J. D. Alanen. Sampling inspection plans for discriminating between two Weibull processes. Technometrics, 7(4):589–601, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266398.

Quesenberry:1964:CPD

[Que64] C. P. Quesenberry. Controlling the proportion defective from classification data. Technometrics, 6(1):99–100, February 1964. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266748.

Qureishi:1964:DBT

[Qur64] A. S. Qureishi. The discrimination between two Weibull processes. Technometrics, 6(1):57–75, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266746. See errata [Ano64g].

Raghavarao:1965:NFP

[Rag65] Damaraju Raghavarao. A note on fractions of 3^{4n+1} plans. Technometrics, 7(1):69–71, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266130.

Ramachandramurty:1966:IBV

[Ram66] P. V. Ramachandramurty. Index to Technometrics: Volumes 1–7: 1959–1965. Technometrics, 8(4):ii + 28, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266630. Separately-numbered

section following page 738.

Randolph:1968:BRBb

[Ran68a] Paul H. Randolph. Book review: Bayesian Decision Problems and Markov Chains by J. J. Martin. Technometrics, 10(4):878, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267477.

Randolph:1968:BRBa

[Ran68b] Paul H. Randolph. Book review: Nolinear Programming by J. Abadie. Technometrics, 10(3):627, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267122.

Randolph:1969:BRB

[Ran69] Paul H. Randolph. Book review: Applied Statistical Decision Theory by Howard Raiffa; Robert Schlaifer. Technometrics, 11(1):205, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266780.

Rechtschaffner:1967:SFF

[Rec67] R. L. Rechtschaffner. Saturated fractions of 2^{n} and 3^{n} factorial designs. Technometrics, 9(4):569–575, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266195.

Reinhardt:1962:BRB

[Rei62] H. E. Reinhardt. Book review: Studies in Item Analysis and Prediction by Herbert Solomon. Technometrics, 4(4):618, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266298.

Reinhardt:1963:BRBb

[Rei63a] H. E. Reinhardt. Book reviews: Mathematical Models in the Social Sciences by John G. Kemeny; J. Laurie Snell; Types of Formalization in Small-Group Research by Joseph Berger; Bernard P. Cohen; J. Laurie Snell; Morris Zelditch. Technometrics, 5(2):288, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266081.

Reinhardt:1963:BRBa

[Rei63b] Howard Reinhardt. Book review: Handbook of Statistical Tables by Donald B. Owen. Technometrics, 5(1):127–128, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266497.

Renner:1969:GMM

[Ren69] Maynard S. Renner. A graphical method for making multiple comparisons of frequencies. Technometrics, 11(2):321–329, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267264.

Riffenburgh:1969:HRT

[Rif69] Robert H. Riffenburgh. Half-rectified truncated distributions: Sampling theory and hypothesis testing. Technometrics, 11(1):47– 59, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1266764.

Riley:1962:CCT

[Ril62] James D. Riley. Comparative cost of two life test procedures. Technometrics, 4(1):140–142, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266182.

Robertson:1960:PFE

[Rob60] W. H. Robertson. Programming Fisher’s exact method of comparing two percentages. Technometrics, 2(1):103–107, February 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266534.

Robson:1961:MCC

[Rob61] D. S. Robson. Multiple comparisons with a control in balanced incomplete block designs. Technometrics, 3(1):103–105, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266480.

Robinson:1963:BRB

[Rob63] Enders A. Robinson. Book review: Cybernetics, or Control and Communication in the Animal and the Machine by Norbert Wiener. Technometrics, 5(1):128–130, February 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266498.

Robbins:1964:BRB

[Rob64] D. K. Robbins. Book review: Introduction to Electronic Computers Problem Solving with the IBM 1620 by F. J. Gruenberger; D. D. McCracken. Technometrics, 6(3):333, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266052.

Robertson:1965:ADD

[Rob65] H. H. Robertson. Approximate design of digital filters. Technometrics, 7(3):387–403, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266597.

Roberts:1966:CSC

[Rob66] S. W. Roberts. A comparison of some control chart procedures. Technometrics, 8(3):411–430, August 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266688.

Roberts:1969:FWV

[Rob69] Charles Roberts. Fill weight variation release and control of capsules, tablets, and sterile solids. Technometrics, 11(1):161–175, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266772.

Rogers:1968:BRBa

[Rog68a] Gerald S. Rogers. Book review: Mathematical Statistics A Decision Theoretic Approach by Thomas S. Ferguson. Technometrics, 10(1):215–216, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266244.

Rogers:1968:BRBb

[Rog68b] Gerald S. Rogers. Book review: Statistical Dictionary of Terms and Symbols by A. K. Kurty; H. A. Edgerton. Technometrics, 10(4):877–878, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267476.

Rogers:1969:BRBb

[Rog69a] Gerald S. Rogers. Book review: A Course in Theoretical Statistics by N. A. Rahman. Technometrics, 11(4):840–841, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266908.

Rogers:1969:BRBa

[Rog69b] Gerald S. Rogers. Book review: The Laws of Large Numbers by Pál Révész. Technometrics, 11(3):625, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267035.

Root:1963:BRB

[Roo63] William L. Root. Book review: Introduction to Linear Systems Analysis by R. G. Brown; J. W. Nilsson. Technometrics, 5(2):286, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266079.

Rosenblatt:1964:BRB

[Ros64] Judah Rosenblatt. Book review: Principles of Reliability by Erich Pieruschka. Technometrics, 6(2):232–234, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266163.

Rosenblatt:1965:BRBb

[Ros65a] Judah Rosenblatt. Book review: Applications of Characteristic Functions by E. Lukacs; R. G. Laha. Technometrics, 7(4):658–659, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266411.

Rosenblatt:1965:BRBa

[Ros65b] Lisa Rosenblatt. Book review: Russian English Dictionary of Statistical Terms and Expressions by Samuel Kotz. Technometrics, 7(1):77, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266135.

Rosenblatt:1966:QCI

[Ros66] Judah Rosenblatt. Query 16: Confidence interval for standard deviation from a single observation. Technometrics, 8(2):367–368, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266372.

Roussas:1968:BRB

[Rou68] George Roussas. Book review: An Introduction to Stochastic Processes by M. S. Bartlett. Technometrics, 10(3):624–625, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267120.

Roy:1967:BRB

[Roy67] Russell Roy, Jr. Book review: Principles of Statistics by M. G. Bulmer. Technometrics, 9(4):698, November 1967. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266210.

Rhyne:1965:TTV

[RS65] A. L. Rhyne, Jr. and R. G. D. Steel. Tables for a treatments versus control multiple comparisons sign test. Technometrics, 7 (3):293–306, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266590.

S:1965:BRBa

[S.65a] G. P. S. Book review: Elphyma Tables by Erik Ingelstam; Stig Sjoberg. Technometrics, 7(1):77, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266134.

S:1965:BRBb

[S.65b] G. P. S. Book review: International Journal of Computer Mathematics by Peter H. Friedlander. Technometrics, 7(1):87–88, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266143.

Saleh:1967:DEO

[Sal67] A. K. Md. Ehsanes Saleh. Determination of the exact optimum order statistics for estimating the parameters of the exponential distribution from censored samples. Technometrics, 9(2):279–292, May 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266424.

Samuels:1969:BRB

[Sam69] S. M. Samuels. Book review: An Introduction to Probability Theory and Its Applications, Vol. 1 by William Feller. Technometrics, 11(2):405, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267275.

Sandelius:1968:GVT

[San68] Martin Sandelius. A graphical version of Tukey’s confidence interval for slippage. Technometrics, 10(1):193–194, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266234.

Sarndal:1964:EPG

[Sär64] Carl-Erik Särndal. Estimation of the parameters of the gamma distribution by sample quantiles. Technometrics, 6(4):405–414, November 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266095.

Satterthwaite:1964:QMT

[Sat64] F. E. Satterthwaite. Query 6: The mean of the tail of a distribution. Technometrics, 6(3):331, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266051.

Sathe:1965:BRB

[Sat65] Y. S. Sathe. Book review: Introduction to the Theory of Statistics by A. M. Mood; F. A. Graybill. Technometrics, 7(3):456, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266611.

Saunders:1968:DSL

[Sau68] Sam C. Saunders. On the determination of a safe life for distributions classified by failure rate. Technometrics, 10(2):361–377, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267051.

Shenton:1967:RLS

[SB67] L. R. Shenton and K. O. Bowman. Remarks on large sample estimators for some discrete distributions. Technometrics, 9(4): 587–598, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266197.

Scheinok:1964:ECC

[Sch64] P. Scheinok. Estimation of a component of a convolution, when the other component is of exponential type. Technometrics, 6(2):222– 224, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266155.

See errata [Ano65j].

Schneider:1966:BRB

[Sch66] A. M. Schneider. Book review: Statistics in Physical Science: Estimation, Hypothesis Testing, and Least Squares by Walter Clark Hamilton. Technometrics, 8(4):712–713, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266646.

Scheaffer:1969:SMM

[Sch69] Richard L. Scheaffer. Sampling mixtures of multi-sized particles: An application of renewal theory. Technometrics, 11(2):285–298, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267261.

Shumway:1968:BLU

[SD68] Robert H. Shumway and William C. Dean. Best linear unbiased estimation for multivariate stationary processes. Technometrics, 10(3):523–534, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267106.

Schilling:1969:PTE

[SD69] E. G. Schilling and H. F. Dodge. Procedures and tables for evaluating dependent mixed acceptance sampling plans. Technometrics, 11(2):341–372, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267266.

Seeger:1968:NMA

[See68] Paul Seeger. A note on a method for the analysis of significances en masse. Technometrics, 10(3):586–593, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267112.

Seward:1966:BRB

[Sew66] J. S. Seward. Book review: Operations Research: Process and Strategy by David S. Stoller. Technometrics, 8(1):200–201, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266277.

Smith:1961:ATD

[SG61] H. Smith and D. W. Grace. Applications of truncated distributions in process start-ups and inventory control. Technometrics, 3 (3):429–432, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266734.

Sobel:1966:BGT

[SG66] Milton Sobel and Phyllis A. Groll. Binomial group-testing with an unknown proportion of defectives. Technometrics, 8(4):631–656, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266636.

Sarhan:1962:MSR

[SGR62] A. E. Sarhan, B. G. Greenberg, and Eleanor Roberts. Modified square root method of matrix inversion. Technometrics, 4(2):282– 287, May 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266626.

See errata [Ano62e].

Spurrell:1967:LE

[SH67] D. J. Spurrell and Charles Dean Hendrix. Letters to the Editor. Technometrics, 9(3):495–497, August 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266522.

Shanbhag:1965:LE

[Sha65] D. N. Shanbhag. Letter to the Editor. Technometrics, 7(3):461, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266613.

Shapiro:1966:BRB

[Sha66] Samuel S. Shapiro. Book review: Statistical Analysis by E. C. Bryant. Technometrics, 8(4):711–712, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266645.

Sherman:1965:NMC

[She65a] Ellen Sherman. A note on multiple comparisons using rank sums. Technometrics, 7(2):255–256, May 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266674.

Sherman:1965:DER

[She65b] Robert E. Sherman. Design and evaluation of a repetitive group sampling plan. Technometrics, 7(1):11–21, February 1965. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266124.

Spendley:1962:SAS

[SHH62] W. Spendley, G. R. Hext, and F. R. Himsworth. Sequential application of simplex designs in optimisation and evolutionary operation. Technometrics, 4(4):441–461, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266283.

Shorack:1967:PPL

[Sho67a] Roger A. Shorack. On the power of precedence life tests. Technometrics, 9(1):154–158, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266325.

Shorack:1967:TDM

[Sho67b] Roger A. Shorack. Tables of the distribution of the Mann– Whitney–Wilcoxon U-statistic under Lehmann alternatives. Technometrics, 9(4):666–677, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266205.

Singh:1966:PCR

[Sin66] H. R. Singh. Producer and consumer risks in non-normal populations. Technometrics, 8(2):335–343, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266366.

Shah:1965:MFR

[SK65] B. K. Shah and C. G. Khatri. A method of fitting the regression curve E(y) = α + δx + βρ^{x}. Technometrics, 7(1):59–65, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266128.

Smith:1963:RFW

[SL63] W. L. Smith and M. R. Leadbetter. On the renewal function for the Weibull distribution. Technometrics, 5(3):393–396, August 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266342.

Stacy:1965:PEG

[SM65] E. W. Stacy and G. A. Mihram. Parameter estimation for a generalized gamma distribution. Technometrics, 7(3):349–358, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266594.

Smith:1966:PPM

[Smi66] W. E. Smith. An a posteriori probability method for solving an overdetermined system of equations. Technometrics, 8(4):675–686, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266639.

Smith:1968:QCQ

[Smi68a] H. Fairfield Smith. Query 28: Comment on query 20: Preference scores (revisited). Technometrics, 10(3):613–619, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267116. See

[JA67].

Smith:1968:QBS

[Smi68b] William B. Smith. Query 29: Bivariate samples with missing values, II. Technometrics, 10(4):867–868, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267468.

Smith:1969:E

[Smi69a] H. Smith, Jr. Editorial. Technometrics, 11(4):629–630, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266885.

Smith:1969:MJP

[Smi69b] Harry Smith, Jr. In memoriam: John P. Comer. Technometrics, 11(4):630, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266886.

Scheuer:1962:GNR

[SS62] Ernest M. Scheuer and David S. Stoller. On the generation of normal random vectors. Technometrics, 4(2):278–281, May 1962. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266625.

Springer:1968:BCL

[ST68] M. D. Springer and W. E. Thompson. Bayesian confidence limits for reliability of redundant systems when tests are terminated at first failure. Technometrics, 10(1):29–36, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266221.

Starks:1964:NSO

[Sta64] T. H. Starks. A note on small orthogonal main-effect plans for factorial experiments. Technometrics, 6(2):220–222, May 1964. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266154.

Steel:1960:RST

[Ste60] Robert G. D. Steel. A rank sum test for comparing all pairs of treatments. Technometrics, 2(2):197–207, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266545.

Steck:1965:BRB

[Ste65] G. P. Steck. Book review: Tables of the Negative Binomial Probability Distribution by Eric Williamson; Michael H. Bretherton. Technometrics, 7(1):83–85, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266141.

Schatzoff:1968:ECA

[STF68] M. Schatzoff, R. Tsao, and S. Fienberg. Efficient calculation of all possible regressions. Technometrics, 10(4):769–779, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267458.

Stone:1961:BRB

[Sto61] M. Stone. Book review: Applied Statistical Decision Theory by H. Raiffa; R. Schlaifer. Technometrics, 3(3):439–440, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266736.

Striebel:1969:BRB

[Str69] Charlotte Striebel. Book review: Advanced Seminar on Spectral Analysis of Time Series, Proceedings of an Advanced Seminar Conducted by the Mathematics Research Center, United States Army, and the Statistics Department at the University of Wisconsin, Madison, October 3–5, 1966 by Bernard Harris. Technometrics, 11(3):622–624, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267034.

Summers:1965:ISC

[Sum65] Robert D. Summers. An inequality for the sample coefficient of variation and an application to variables sampling. Technometrics, 7(1):67–68, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266129.

Smith:1968:LMP

[SV68] Barnard E. Smith and Ramakrishna R. Vemuganti. A learning model for processes with tool wear. Technometrics, 10(2):379–387, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267052.

Sathe:1969:MVU

[SV69] Y. S. Sathe and S. D. Varde. Minimum variance unbiased estimation of reliability for the truncated exponential distribution. Technometrics, 11(3):609–612, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267028.

Shapiro:1968:AND

[SW68] S. S. Shapiro and M. B. Wilk. Approximations for the null distribution of the W statistic. Technometrics, 10(4):861–866, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267467.

Syski:1963:BRB

[Sys63] R. Syski. Book review: The Theory of Probability by B. V. Gnedenko. Technometrics, 5(2):281–285, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266077.

Steck:1968:RBN

[SZ68] G. P. Steck and W. J. Zimmer. The relationship between Neyman and Bayes confidence intervals for the hypergeometric parameter. Technometrics, 10(1):199–203, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266236.

Tallis:1967:AML

[Tal67] G. M. Tallis. Approximate maximum likelihood estimates from grouped data. Technometrics, 9(4):599–606, November 1967. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266198.

Taneja:1968:BRB

[Tan68] V. S. Taneja. Book review: Introduction to Statistical Mathematics by A. M. Mathai. Technometrics, 10(3):623–624, August 1968. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267119.

Taylor:1967:SCG

[Tay67] Howard M. Taylor. Statistical control of a Gaussian process. Technometrics, 9(1):29–41, February 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266316.

Taylor:1968:EDC

[Tay68] Howard M. Taylor. The economic design of cumulative sum control charts. Technometrics, 10(3):479–488, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267102.

Thoman:1969:TST

[TB69] Darrel R. Thoman and Lee J. Bain. Two sample tests in the Weibull distribution. Technometrics, 11(4):805–815, November 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266900.

Thoman:1969:IPW

[TBA69] Darrel R. Thoman, Lee J. Bain, and Charles E. Antle. Inferences on the parameters of the Weibull distribution. Technometrics, 11(3):445–460, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267019.

Tipton:1968:BRB

[TC68] Ronald B. Tipton and Frank F. Carden. Book review: Statistical Communication and Detection with Special Reference to Digital Data Processing of Radar and Seismic Signals by Enders A.

Robinson. Technometrics, 10(4):880–881, November 1968. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267481.

Terry:1965:BRB

[Ter65] M. E. Terry. Book review: Advances in Computers, Volume 5 by Franz L. Alt; Morris Rubinoff. Technometrics, 7(4):661–662, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266413.

Tiao:1967:AOA

[TG67] G. C. Tiao and Irwin Guttman. Analysis of outliers with adjusted residuals. Technometrics, 9(4):541–559, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266193.

Thompson:1963:BRB

[Tho63] W. A. Thompson, Jr. Book review: Mathematical Statistics by John E. Freund. Technometrics, 5(3):412–413, August 1963. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266350.

Thompson:1967:BRB

[Tho67] W. A. Thompson, Jr. Book review: Elements of Probability Theory by J. Bass; W. McKay. Technometrics, 9(4):697–698, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266209.

Thrall:1964:BRB

[Thr64] R. M. Thrall. Book review: A Manager’s Guide to Operations Research by Russell L. Ackoff; Patrick Rivett. Technometrics, 6 (3):333–334, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266053.

Tick:1961:ETF

[Tic61] Leo J. Tick. The estimation of “transfer functions” of quadratic systems. Technometrics, 3(4):563–567, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266562.

Tick:1962:LE

[Tic62] Leo J. Tick. Letter to the Editor. Technometrics, 4(3):436, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266582.

Tingey:1963:SES

[Tin63] Fred H. Tingey. Statistical evaluation of splitting limit criteria in measurement disputes. Technometrics, 5(2):263–269, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266067.

Tiplitz:1963:AMS

[Tip63] C. Tiplitz. Approximations to the mean and standard deviation of reciprocals of observations. Technometrics, 5(4):522–524, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266030.

Tallis:1968:UFM

[TL68] G. M. Tallis and R. Light. The use of fractional moments for estimating the parameters of a mixed exponential distribution. Technometrics, 10(1):161–175, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266232.

Thompson:1963:NNE

[TM63] W. A. Thompson, Jr. and James R. Moore. Non-negative estimates of variance components. Technometrics, 5(4):441–449, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266020.

Thompson:1968:RSD

[TM68] William O. Thompson and Raymond H. Myers. Response surface designs for experiments with mixtures. Technometrics, 10 (4):739–756, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267456.

Tornheim:1963:CNR

[Tor63] Leonard Tornheim. Convergence in nonlinear regression. Technometrics, 5(4):513–514, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266027.

Trauth:1965:BRB

[Tra65] Charles A. Trauth. Book review: Elements of Linear Programming with Economic Applications by R. C. Geary; M. D. Mc-Carthy. Technometrics, 7(3):453, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266607.

Thompson:1964:SDR

[TS64] H. R. Thompson and K. E. Seal. Serial designs for routine quality control and experimentation. Technometrics, 6(1):77–98, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266747.

Tick:1966:SRA

[TS66] Leo Joseph Tick and Paul Shaman. Sampling rates and appearance of stationary Gaussian processes. Technometrics, 8(1):91–106, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266264.

Tobach:1967:TMR

[TSRR67] Ethel Tobach, Mark Smith, George Rose, and Donald Richter. A table for making rank sum multiple paired comparisons. Technometrics, 9(4):561–567, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266194.

Tucker:1965:BRB

[Tuc65] Howard G. Tucker. Book review: Limit Theorems for Convolutions by Harald Bergström. Technometrics, 7(2):263, May 1965. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266677.

Tucker:1968:BRB

[Tuc68] Henry Tucker. Book review: Introduction to Statistics by J. M. Bevan. Technometrics, 10(4):871–872, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267470.

Tukey:1960:CVD

[Tuk60] John W. Tukey. Conclusions vs. decisions. Technometrics, 2(4): 423–433, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266451.

Tukey:1961:DECb

[Tuk61] John W. Tukey. Discussion, emphasizing the connection between analysis of variance and spectrum analysis. Technometrics, 3(2): 191–219, May 1961. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1266112. See comments [JP61].

Utter:1966:BRB

[Utt66] Robert F. Utter. Book review: A Programmed Introduction to PERT by Federal Electric Corporation. Technometrics, 8(2):371– 372, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266375.

Vandenberg:1964:BRB

[Van64] Steven G. Vandenberg. Book review: Digital Computers in Research, an Introduction for Behavioral Scientists by Bert F. Green. Technometrics, 6(2):234, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266164.

Versace:1966:BRB

[Ver66] John Versace. Book review: Descriptive Statistics. A Programed Textbook by Lassar G. Gotkin; Leo S. Goldstein. Technometrics, 8(4):727, November 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266661.

Veale:1969:EMW

[VH69] James R. Veale and D. V. Huntsberger. Estimation of a mean when one observation May be spurious. Technometrics, 11(2):331–339, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267265.

Vitale:1968:NAD

[VT68] P. Vitale and G. Taylor. A note on the application of Davidon’s method to nonlinear regression problems. Technometrics, 10(4):843–849, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267464.

Walsh:1968:BRB

[Wal68] John E. Walsh. Book review: Elements of Nonparametric Statistics by G. E. Noether. Technometrics, 10(4):878–879, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267479.

Wallenius:1969:SRA

[Wal69] K. T. Wallenius. Sequential reliability assurance in finite lots. Technometrics, 11(1):61–74, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266765.

Wasan:1964:SOP

[Was64] M. T. Wasan. Sequential optimum procedures for unbiased estimation of a binomial parameter. Technometrics, 6(3):259–271, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266043.

Watson:1961:SGS

[Wat61] G. S. Watson. A study of the group screening method. Technometrics, 3(3):371–388, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1266728. See notes [Pat63a, Gur65].

Watson:1969:BRBa

[Wat69a] Geoffrey Watson. Book review: Scientific Method in Analysis of Sediments by John C. Griffiths. Technometrics, 11(2):406, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267276.

Watson:1969:BRBb

[Wat69b] Geoffrey S. Watson. Book review: The Theory of Applied Probability by R. C. Dubes. Technometrics, 11(3):621–622, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267033.

Weissberg:1960:TTL

[WB60] Alfred Weissberg and Glenn H. Beatty. Tables of tolerance-limit factors for normal distributions. Technometrics, 2(4):483–500, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266456.

See errata [Ano61g].

Webb:1968:NOD

[Web68a] Steve Webb. Non-orthogonal designs of even resolution. Technometrics, 10(2):291–299, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267044.

Webb:1968:SSF

[Web68b] Steve R. Webb. Saturated sequential factorial designs. Technometrics, 10(3):535–550, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267107.

Webster:1968:AS

[Web68c] J. T. Webster. An approximate F -statistic. Technometrics, 10 (3):597–604, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267114.

Weiler:1961:POA

[Wei61a] H. Weiler. A problem of optimum allocation arising in chemical analyses by multiple isotope dilution. Technometrics, 3(4):509–518, November 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266557.

Weiss:1961:RCE

[Wei61b] George H. Weiss. The reliability of components exhibiting cumulative damage effects. Technometrics, 3(3):413–422, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266732.

Weiss:1962:BRB

[Wei62a] George Weiss. Book review: Studies in Applied Probability and Management Science by K. J. Arrow; S. Karlin; H. Scarf. Technometrics, 4(4):616–617, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266296.

Weiss:1962:CRS

[Wei62b] George Weiss. On certain redundant systems which operate at discrete times. Technometrics, 4(1):69–74, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266173.

Weiler:1964:STS

[Wei64a] H. Weiler. A significance test for simultaneous quantal and quantitative responses. Technometrics, 6(3):273–285, August 1964. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266044.

Weinstein:1964:QSV

[Wei64b] M. A. Weinstein. Query 2: The sum of values from a normal and a truncated normal distribution. Technometrics, 6(1):104–105, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266751.

Weiler:1965:UIB

[Wei65] H. Weiler. The use of incomplete beta functions for prior distributions in binomial sampling. Technometrics, 7(3):335–347, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266593.

Weiler:1966:CMG

[Wei66] H. Weiler. A coefficient measuring the goodness of fit. Technometrics, 8(2):327–334, May 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1266365. See errata [Ano67h].

Welford:1962:NMC

[Wel62] B. P. Welford. Note on a method for calculating corrected sums of squares and products. Technometrics, 4(3):419–420, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266577.

Westlake:1962:NAP

[Wes62] W. J. Westlake. A numerical analysis problem in constrained quadratic regression analysis. Technometrics, 4(3):426–430, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266579.

Wetherill:1960:SRB

[Wet60] G. B. Wetherill. Some remarks on the Bayesian solution of the single sample inspection scheme. Technometrics, 2(3):341–352, August 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266248. See

discussion [BLH^{+}60].

Wilk:1962:PPG

[WGH62] M. B. Wilk, R. Gnanadesikan, and M. J. Huyett. Probability plots for the gamma distribution. Technometrics, 4(1):1–20, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266169.

White:1969:MLW

[Whi69] John S. White. The moments of log-Weibull order statistics. Technometrics, 11(2):373–386, May 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267267.

Wiggins:1967:MCM

[Wig67] Alvin D. Wiggins. A minimum cost model of spare parts inventory control. Technometrics, 9(4):661–665, November 1967. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266204.

Wilkie:1965:CSL

[Wil65] D. Wilkie. Complete set of leading coefficients, λ(r, n), for orthogonal polynomials up to n = 26. Technometrics, 7(4):644–648, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266402.

Williams:1966:PSA

[Wil66] Theodore J. Williams. The present status of automatic production and control devices and expected future developments. Technometrics, 8(1):73–89, February 1966. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266263.

Williams:1969:NRM

[Wil69] E. J. Williams. A note on regression methods in calibration. Technometrics, 11(1):189–192, February 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266774.

Whitwell:1961:RDR

[WM61] John C. Whitwell and Graham K. Morbey. Reduced designs of resolution five. Technometrics, 3(4):459–477, November 1961. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266554.

Wolfe:1968:BRB

[Wol68] Shirley Wolfe. Book review: The Bases of Fortran: A Self-Training Approach to Computer Programming by R. E. Smith. Technometrics, 10(2):412–413, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1267064.

Wolfe:1977:DFT

[Wol77] Douglas A. Wolfe. A distribution-free test for related correlation coefficients. Technometrics, 19(4):507–509, November 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267893. See er-

rata [Ano78].

Wonnacott:1961:SAC

[Won61] Thomas H. Wonnacott. Spectral analysis combining a Bartlett window with an associated inner window. Technometrics, 3(2):235– 243, May 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266115.

Wouk:1963:BRB

[Wou63] Arthur Wouk. Book review: Linear Programming by G. Hadley. Technometrics, 5(2):288–289, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266082.

Wouk:1964:BRB

[Wou64] Arthur Wouk. Book review: Readings in Mathematical Programming by S. Vajda. Technometrics, 6(2):229, May 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266158.

Wouk:1965:BRB

[Wou65] A. Wouk. Book review: Nonlinear and Dynamic Programming by G. Hadley. Technometrics, 7(3):454, August 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266609.

Wouk:1966:SPC

[Wou66] A. Wouk. Statistical properties of a certain periodic binary process. Technometrics, 8(2):247–258, May 1966. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266359.

Wasan:1969:TIG

[WR69] M. T. Wasan and L. K. Roy. Tables of inverse Gaussian percentage points. Technometrics, 11(3):591–604, August 1969. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267026.

Wilk:1968:JAN

[WS68] M. B. Wilk and S. S. Shapiro. The joint assessment of normality of several independent samples. Technometrics, 10(4):825–839, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267462.

Watson:1961:PIM

[WW61] G. S. Watson and W. T. Wells. On the possibility of improving the mean useful life of items by eliminating those with short lives. Technometrics, 3(2):281–298, May 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266118.

Weeks:1964:NDC

[WW64] David L. Weeks and Donald R. Williams. A note on the determination of connectedness in an N-way cross classification. Technometrics, 6(3):319–324, August 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1266048. See errata [Ano65i].

Youden:1961:PCF

[You61] W. J. Youden. Partial confounding in fractional replication. Technometrics, 3(3):353–358, August 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266726.

Youden:1962:SEP

[You62] W. J. Youden. Systematic errors in physical constants. Technometrics, 4(1):111–123, February 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266176.

Youden:1964:QIR

[You64] W. J. Youden. Query 1: Inadmissible random assignments. Technometrics, 6(1):103–104, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266750.

Zacks:1968:BRB

[Zac68] S. Zacks. Book review: Combinatorial Methods in the Theory of Stochastic Processes by L. Takács. Technometrics, 10(3):630–631, August 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267125.

Zelen:1961:RLT

[ZD61] M. Zelen and Mary C. Dannemiller. The robustness of life testing procedures derived from the exponential distribution. Technometrics, 3(1):29–49, February 1961. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266475.

Zelen:1959:FEL

[Zel59] Marvin Zelen. Factorial experiments in life testing. Technometrics, 1(3):269–288, August 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1266445. See errata [Ano60d].

Zelen:1962:BRBb

[Zel62a] M. Zelen. Book review: The Smirnov Two-Sample Statistic D_{mn}(n ≤ m ≤ 20) by G. S. Marliss; W. Zayachkowski. Technometrics, 4(4):619, November 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266300.

Zelen:1962:BRBa

[Zel62b] Marvin Zelen. Book review: A Guide to Operational Research by Eric Duckworth. Technometrics, 4(3):438, August 1962. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266585.

Zelen:1963:BRBb

[Zel63a] M. Zelen. Book review: Bibliography of Statistical Literature 1950– 1958 by Maurice G. Kendall; Alison G. Doig. Technometrics, 5 (4):530–531, November 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266035.

Zelen:1963:BRBa

[Zel63b] Marvin Zelen. Book review: Contributions to Order Statistics by Ahmed E. Sarhan; Bernard G. Greenberg. Technometrics, 5(2):274–275, May 1963. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266071.

Zimmer:1968:BRB

[Zim68] William J. Zimmer. Book review: Quality Control and Reliability by N. L. Enrick. Technometrics, 10(4):873–874, November 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267472.

Zeigler:1968:DSP

[ZT68] R. K. Zeigler and G. L. Tietjen. Double sampling plans where the acceptance criterion is the variance. Technometrics, 10(1):99–106, February 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266227.

Zyskind:1964:BRB

[Zys64] George Zyskind. Book review: An Introduction to Linear Statistical Models, Volume 1 by Franklin A. Graybill. Technometrics, 6(1):107–111, February 1964. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266752.

Zyskind:1965:QMV

[Zys65] George Zyskind. Query 14: Missing values in factorial experiments. Technometrics, 7(4):649–650, November 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266403.